### v1

# Probing the accretion geometry of black hole X-ray binaries: A multi-mission spectro-polarimetric and timing study

Abstract
We present a comprehensive spectro-polarimetric and timing analysis of eleven black hole X-ray binaries, namely Cyg X11-1- 1, 4U $1630-47$, Cyg X33-3- 3, LMC X11-1- 1, 4U $1957+115$, LMC X33-3- 3, Swift J1727.816131727.816131727.8-16131727.8 - 1613, GX $339-4$, Swift J151857.0572147151857.0572147151857.0-572147151857.0 - 572147, IGR J17091362417091362417091-362417091 - 3624, and MAXI J174429417442941744-2941744 - 294, using quasi-simultaneous observations from IXPE, NICER, NuSTAR, and AstroSat. Timing analyses reveal characteristic type-B and type-C Quasi-periodic Oscillations (QPOs) across different spectral states, often associated with episodic radio ejections. Broadband ($0.7-60$ keV) spectral modelling, employing disc, Comptonization, and reflection components, reveals degeneracies in constraining disc-corona geometries. Polarimetric measurements in $2-8$ keV band detect significant polarization degrees (PDs) ranging from $3-20.6\%$ ($1.2-21.4\%$) in harder (softer) states, with moderate to strong energy dependence, except for LMC X11-1- 1, Swift J151857.0572147151857.0572147151857.0-572147151857.0 - 572147, and MAXI J174429417442941744-2941744 - 294, where no significant polarization is detected. We report the polarization detections of Cyg X33-3- 3 (PD $\sim 21.4\%$, SIMS), LMC X33-3- 3 (PD $\sim 2.4\%$, HSS) and IGR J17091362417091362417091-362417091 - 3624 (PD $\sim 9\%$, LHS) using the recent IXPE observations. A positive correlation is found between PD and the Comptonized photon fraction ($cov_{\rm frac}$), while an anti-correlation is observed with the disc-to-Comptonized flux ratio ($F_{\rm ratio}$) across spectral states. The combined timing, spectral, and polarimetric results, together with constraints from radio jet observations, suggest a radially extended corona within a truncated disc for Cyg X11-1- 1, Swift J1727.816131727.816131727.8-16131727.8 - 1613, IGR J17091362417091362417091-362417091 - 3624, and GX $339-4$, whereas the disc-corona geometry remains unconstrained for 4U $1957+115$, LMC X33-3- 3, and 4U $1630-47$. We discuss the implications of these findings for understanding accretion geometries and highlight prospects for future X-ray polarimetric studies.
keywords:
accretion, accretion disc – black hole physics – polarization – techniques: polarimetric – radiation mechanisms: general – X-rays: binaries – stars: individual
pubyear: 2025
1 Introduction
Black hole X-ray binaries (BH-XRBs) are believed to be the ideal candidates for understanding the physical processes that govern the radiation mechanism around the compact objects. BH-XRBs often exhibit distinct spectral states over timescales ranging from days to months, strongly associated with the underlying accretion dynamics of the system (Morgan et al., 1997; Zhang et al., 1997; Paul et al., 1998; Sobczak et al., 1999; Done & Zycki, 1999; Chakrabarti & Manickam, 2000; Homan et al., 2001; Corbel et al., 2001; Titarchuk & Shrader, 2002; Pottschmidt et al., 2003; Rodriguez et al., 2003; Corbel et al., 2004; Belloni et al., 2005; Kalemci et al., 2006; Remillard & McClintock, 2006; Yu & Yan, 2009; Shaposhnikov et al., 2010; Tarana et al., 2011; Nandi et al., 2012; Sriram et al., 2013; Seifina et al., 2014; Iyer et al., 2015; Yan & Yu, 2017; Sreehari et al., 2019b, 2020; Baby et al., 2020; Majumder et al., 2022; Athulya et al., 2022; Prabhakar et al., 2023; Nandi et al., 2024; Li et al., 2025; Majumder et al., 2025b, and references therein). Empirically, the high-soft state (HSS) is characterized by the multi-temperature black-body emission that is likely to emerge from an optically thick and geometrically thin accretion disc Shakura & Sunyaev (1973). In contrast, the canonical low-hard state (LHS), represented by a power-law profile with high energy cut-off, is found to be dominated by the hard emission produced from the Compton upscattering of disc photons into the ‘hot’ electron cloud (equivalently X-ray corona). Moreover, depending on the relative contributions from the disc and coronal components, hard/soft intermediate states (HIMS/SIMS) are observed.
Needless to mention that the possible geometry of the Comptonizing corona in BH-XRBs remains an unsettled issue to date. Several alternative scenarios featuring the models of the sandwich corona Haardt & Maraschi (1993); Stern et al. (1995), radially elongated corona at the truncated inner accretion disc Eardley et al. (1975); Chakrabarti & Titarchuk (1995); Poutanen et al. (1997); Iyer et al. (2015) and vertically extended corona as the base of the jet Miyamoto & Kitamoto (1991); Markoff et al. (2005); Méndez et al. (2022); Zhang et al. (2023) have been proposed over the years. However, the overall geometry of the disc-corona-jet remains elusive mostly due to the model degeneracies that complicate the interpretation of the observational data.
Notably, the temporal properties of BH-XRBs show rapid X-ray variability over different timescales. This variability is usually observed in the power density spectrum (PDS) and is closely correlated to the spectral states. In particular, transient phenomena like strong and stable Low-Frequency Quasi-periodic Oscillations (LFQPOs) on a wide range of frequencies distinguish the spectral states and act as the precursor of the state transitions in BH-XRBs (Remillard & McClintock, 2006; Done et al., 2007; Nandi et al., 2012; Iyer et al., 2015). For example, the LHS and HIMS are characterized by the appearance of strong, coherent, variable peaked type-C LFQPO of frequency $\sim 0.1-15$ Hz superposed on a flat-top noise (FTN) component in the PDS (Remillard & McClintock, 2006; Nandi et al., 2012, and references therein). The origin of these type-C QPOs is often explained through various mechanisms, including oscillations of radiative shock waves within the accretion disc (Molteni et al., 1996; Chakrabarti et al., 2008; Das et al., 2014), relativistic Lense-Thirring precession of the inner hot flow or the truncated disc (Stella & Vietri, 1998; Ingram et al., 2009), precession of small-scale jets (Ma et al., 2021) and outward drift of the truncated inner disc radius enveloped by the corona (Karpouzas et al., 2020; Bellavita et al., 2022, 2025). The wide array of interpretations introduces a degeneracy among the different disc-corona-jet configurations used in explaining the QPO phenomena (see Ingram & Motta 2019 for a review).
On the other hand, relatively weak type-B/type-A QPOs appear at a narrow frequency range of around $\sim 6-8$ Hz (Casella et al., 2005) during the SIMS. In this state, FTN is absent, and the PDS continuum shows weak red noise characterized by a simple power-law in PDS continuum. These type-B/type-A QPOs are often found to be closely connected with the radio ejections generally observed in the SIMS (Soleri et al., 2008; Fender et al., 2009; Kylafis et al., 2020; Homan et al., 2020; García et al., 2021; Liu et al., 2022; Zhang et al., 2023). Usually, soft states are characterized by less variability in the PDS without the detection of QPO like features (Belloni et al., 1999; Belloni et al., 2005; Nandi et al., 2012; Radhika & Nandi, 2014; Radhika et al., 2016).
Furthermore, X-ray polarimetric study is also considered as a powerful diagnostic tool to infer the accretion geometry of the BH-XRBs. The recent launch of IXPE Weisskopf et al. (2022), a polarimetric mission sensitive to low-energy ($2-8$ keV) X-rays, enables the opportunity to investigate in-depth polarimetric properties of BH-XRBs (see Dovčiak et al. 2024 for a summary). So far, IXPE has observed eleven BH-XRBs, with significant polarized emission detected in nine sources, namely Cyg X11-1- 1 Krawczynski et al. (2022); Jana & Chang (2024); Steiner et al. (2024), 4U $1630-47$ Kushwaha et al. (2023b); Rawat et al. (2023a, b); Rodriguez Cavero et al. (2023); Ratheesh et al. (2024), Cyg X33-3- 3 Veledina et al. (2024a, b), LMC X33-3- 3 Majumder et al. (2024a); Svoboda et al. (2024a), 4U $1957+115$ Kushwaha et al. (2023a); Marra et al. (2024), Swift J1727.816131727.816131727.8-16131727.8 - 1613 Veledina et al. (2023); Ingram et al. (2024); Svoboda et al. (2024b); Podgorný et al. (2024), Swift J151857.0572147151857.0572147151857.0-572147151857.0 - 572147 Mondal et al. (2024), GX $339-4$ Mastroserio et al. (2025) and IGR J17091362417091362417091-362417091 - 3624 Ewing et al. (2025). Notably, no significant polarization is observed in LMC X11-1- 1 Podgorný et al. (2023).
Despite the significant advancements in X-ray polarimetry, interpreting the observed polarization degree (PD) and polarization angle (PA) within the framework of theoretical models remains a formidable challenge. The classical work by Chandrasekhar (Chandrasekhar, 1960) on semi-infinite electron scattering predicts a low PD of $\sim 2\%$ from the accretion disc of highly inclined systems. However, 4U 1630--47 shows a remarkably higher PD of around $8.3\%$ in its disc-dominated thermal state, deviating substantially from these theoretical expectations (Kushwaha et al., 2023b; Ratheesh et al., 2024). Furthermore, PDs of about $4\%$ have been observed in the LHS of Cyg X--1 and Swift J1727.8--1613 (Krawczynski et al., 2022; Ingram et al., 2024), whereas existing models predict only $\sim 1\%$ PD from a wedge-shaped corona in such low-inclined systems (Krawczynski et al., 2022). To reconcile this discrepancy, Poutanen et al. (2023) proposed that an outflowing corona with mildly relativistic motion could produce PD levels consistent with observations. It is worth mentioning that the recent detection of $\sim 9\%$ PD in the LHS of IGR J17091--3624 (Ewing et al., 2025) further indicates the limitations of current theoretical frameworks and highlights the need for more comprehensive models of X-ray polarization.
Moreover, in certain cases, the spectro-temporal findings of BH-XRBs appear to be contradictory with interpretations derived from X-ray polarimetry. For instance, in Swift J1727.8--1613, the evolution of type-C QPOs suggests the presence of a jet-like corona aligned perpendicular to the disc plane (Liao et al., 2024). In contrast, simultaneous polarimetric observations of the same source indicate a radially extended corona resided at the equatorial plane of the disc (Veledina et al., 2023; Ingram et al., 2024). Therefore, it is evident that the polarimetric findings offer challenges to the existing X-ray spectro-temporal models, indicating that a deeper understanding of the accretion dynamics and disc-corona-jet geometry in black hole X-ray binaries is yet to be unveiled.
Being motivated by this, we undertake in-depth timing and spectro-polarimetric analyses of eleven BH-XRBs using quasi-simultaneous archival observations from IXPE, NICER, NuSTAR and AstroSat, covering a broad energy range of $0.5-100$ keV. While doing so, we study the evolution of X-ray (soft/hard) and radio flux densities for all sources using multi-mission long-term monitoring data from MAXI/GSC, Swift/BAT, RATAN and VLITE. Further, we conduct a comprehensive spectro-temporal study of these sources using NICER, NuSTAR and AstroSat data and confirm the presence of distinct spectral states of the BH-XRBs during the observational campaigns. Using alternative model prescriptions relying on different assumptions on the disc-corona configurations, we attempt to explain the spectral features of the sources and investigate the possibility of having inherent degeneracy among the models. Next, we deduce the polarization properties of the sources using IXPE data in the $2-8$ keV energy range, followed by a detailed spectro-polarimetric correlation study. Subsequently, noticeable positive and negative correlations between polarization degree and spectral parameters serve as powerful diagnostics, offering deeper insights into the complex accretion dynamics and geometry of the BH-XRBs under consideration. Finally, we outline the future prospects of X-ray polarimetry in light of the present findings.
The paper is organized as follows. In §2 and §3, we present the source selection with observation details and the data reduction procedures, respectively. The analysis and results of the timing and spectro-polarimetric study are presented in §4. Finally, we discuss the results and present conclusions in §5 and §6, respectively.
Figure 1: Long-term X-ray monitoring of the sources with MAXI/GSC ($2-20$ keV), Swift/BAT ($15-50$ keV) and the corresponding hardness ratio obtained with MAXI/GSC are presented in panel (a), (b) and (c), respectively. Each data point for a given source represents averaged out respective quantity over the exposure of corresponding IXPE observation. In panel (d), the variation of radio flux is shown from the quasi-simultaneous radio observations available from various facilities. The $1\sigma$ errorbars of the respective quantities are small and remain within the markers for most of the cases. For a given source, the data points plotted with MJD, correspond to the respective epochs as mentioned in Table 1. See the text for details.
2 Source Selection and observations
In this work, we analyze all the IXPE (Weisskopf et al., 2022) observations of BH-XRBs conducted so far along with simultaneous/quasi-simultaneous NICER, NuSTAR and AstroSat observations depending on data availability. As of now, IXPE observed eleven BH-XRBs, namely Cyg X11-1- 1, 4U 1630--47, Cyg X33-3- 3, LMC X11-1- 1, 4U 1957$+115$, LMC X33-3- 3, Swift J1727.816131613-1613- 1613, Swift J151857.0572147151857.0572147151857.0-572147151857.0 - 572147, GX $339-4$, IGR J17091362417091362417091-362417091 - 3624 and MAXI J174429417442941744-2941744 - 294, during its first three and a half years of operational period. The IXPE, NICER and NuSTAR observations used in this work are publicly available in the HEASARC [^1] https://heasarc.gsfc.nasa.gov/db-perl/W3Browse/w3browse.pl database and the AstroSat data is archived at the ISSDC [^2] https://webapps.issdc.gov.in/astro_archive/archive/Home.jsp website. The quasi-simultaneous multi-mission observations of the sources are separated by at most three days from the IXPE epoch except LMC X11-1- 1 for which the NuSTAR observation lies four days after the IXPE observation. The sources exhibiting marginal spectro-temporal variability between observations from different missions offer an opportunity for a combined multi-mission study. All the observations analyzed in this study are tabulated in Table 1 along with the date and ObsID. Brief overview of each of the sources under consideration are presented below.
2.1 Cyg X--1
Cyg X-1, the first galactic BH-XRB discovered in 1971 Webster & Murdin (1972), continues to be one of the most extensively studied celestial objects. Recent measurements reveal that it hosts a black hole with a mass of $21.2\pm 2.2$ $M_{\odot}$ Miller-Jones et al. (2021), located at a distance of $2.2\pm 0.2$ kpc Miller-Jones et al. (2021), with a supergiant O-type star as its binary companion. Meanwhile, numerous studies confirm the presence of a maximally rotating ‘hole’ having spin greater than $0.99$ at the center core of the binary system (Tomsick et al., 2013; Zhao et al., 2021; Kushwaha et al., 2021). Cyg X--1 has remained persistently bright, mostly in the LHS, though it has transitioned to the HSS through short-lived intermediate states over the past few decades (Kushwaha et al., 2021).
2.2 4U 1630--47
4U 1630--47 is a recurrent X-ray transient discovered by Uhuru Jones et al. (1976). Since discovery, the source has exhibited more than $20$ quasi-period outbursts Baby et al. (2020); Chatterjee et al. (2022) typically in every $600-700$ days Parmar et al. (1995). Since the optical counterpart of the source is still unknown, the dynamical measurement of its mass and distance remains elusive. Several efforts relying on indirect measurements constrained its mass as $10\pm 0.1$ $M_{\odot}$ Seifina et al. (2014) and distance as $10-11$ kpc Seifina et al. (2014); Kalemci et al. (2018). The presence of dips in the light curves of the $1996$ outburst indicates a high inclination ($i\sim 60-70^{\circ}$) of the system Tomsick et al. (1998); Seifina et al. (2014). Spectral modeling Pahari et al. (2018) and spectro-polarimetric fitting in the high soft state confirm the source’s spin as $\sim 0.92$ Kushwaha et al. (2023b). 4U 16304747-47- 47 is observed in the HSS in most of its outbursts till date, possibly due to fast transition from LHS to HSS Capitanio et al. (2015); Baby et al. (2020); Peng et al. (2024c); Fan et al. (2024); Parra et al. (2025).
2.3 Cyg X--3
Cyg X33-3- 3 was discovered over five decades ago Giacconi et al. (1967) and is a high-mass X-ray binary system hosting a compact object accreting from a Wolf-Rayet donor star van Kerkwijk et al. (1992). Despite extensive studies over the years, the nature of the compact object remains uncertain. Recent polarization measurements suggest that it could be a concealed Galactic ultra-luminous X-ray source (ULX) surpassing the Eddington limit due to anisotropic emissions Veledina et al. (2024a). It is worth mentioning that Cyg X33-3- 3 exhibits exceptional radio activity Martí et al. (2000); Miller-Jones et al. (2004) which is tightly correlated with the spectral states on several occasions including the detection of innermost jet confined within the X-ray funnel-like structure Yang et al. (2023). Recent trigonometric parallax measurements from VLBA observations have precisely constrained the source’s distance as ${9.67}_{-0.48}^{+0.53}$ kpc Reid & Miller-Jones (2023).
2.4 LMC X--1
LMC X11-1- 1 is the first discovered extra-galactic BH-XRB located in the Large Magellanic Cloud (LMC) Mark et al. (1969) at a well-constrained distance of $\sim 48.1$ kpc Pietrzyński et al. (2013). Dynamical measurements have determined the black hole’s mass to be $10.9\pm 1.4$ $M_{\odot}$ with a moderate inclination of $36.4^{{}^{\circ}}\pm 1.9^{{}^{\circ}}$ Orosz et al. (2009). Unlike typical X-ray binaries, LMC X11-1- 1 consistently remains in the HSS, with thermal emission contributing around $80\%$ of the spectral flux Steiner et al. (2012); Bhuvana et al. (2021); Jana et al. (2021) and a luminosity of approximately $2\times 10^{38}$ erg $\rm s^{-1}$ Nowak et al. (2001); Bhuvana et al. (2021). The measurements using continuum fitting method in the thermally dominated state constraints LMC X11-1- 1 spin as $\sim 0.92$ Gou et al. (2009); Bhuvana et al. (2021), indicating the presence of a rapidly rotating black hole.
2.5 4U 1957$+$115
4U 1957$+$115 is a bright and persistent low-mass X-ray binary, first discovered by the Uhuru mission in $1973$ Giacconi et al. (1974). To date, it has consistently been observed in a disc-dominated spectral state, with flux levels varying between $20$ and $70$ mCrab Yaqoob et al. (1993); Maitra et al. (2014); Mudambi et al. (2022). Despite decades of observations with various X-ray missions, key system parameters such as distance, inclination, spin, and mass remain highly uncertain. For instance, a high source distance of $20-40$ kpc and a low inclination ($\sim 13^{\circ}$) have been proposed for a maximum black hole mass of $\sim 6$ $M_{\odot}$ Gomez et al. (2015). In contrast, Maitra et al. (2014) reported a high inclination ($\sim 78^{\circ}$) and a distance between $5-10$ kpc for a black hole mass of less than $10$ $M_{\odot}$. Nevertheless, several studies suggest the presence of a maximally spinning black hole in 4U 1957$+$115 Barillier et al. (2023).
2.6 LMC X--3
LMC X33-3- 3 is a persistent extra-galactic BH-XRB in Large Magellanic Cloud (LMC) at a distance of $48.1$ kpc Orosz et al. (2009), discovered by UHURU in $1971$. A slowly rotating black hole of spin $0.25_{-0.29}^{+0.20}$ Steiner et al. (2010); Bhuvana et al. (2021) and mass $6.98\pm 0.56~{}{\rm M}_{\odot}$ Orosz et al. (2014); Bhuvana et al. (2021) is found to be present at the central core of the system. LMC X33-3- 3 is a relatively high inclined system ($i\sim 69.24^{\circ}\pm 0.72^{\circ}$) Orosz et al. (2014). Being a persistent source, it mostly remains in the HSS Bhuvana et al. (2021) except for a few occasions during which the LHS is also observed Smale & Boyd (2012). An exceptionally anomalous low state of peak luminosity $\sim 10^{35}$ erg $\rm s^{-1}$ is also observed for a few instances Torpin et al. (2017).
2.7 Swift J1727.8--1613
Swift J1727.816131727.816131727.8-16131727.8 - 1613 is a recently discovered BH-XRB transient, first detected by Swift/BAT on August 24, 2023. Immediate monitoring of the source with MAXI/GSC in the $2-20$ keV energy range revealed an ‘unusual’ peak in X-ray flux, increasing from $150$ mCrab to $7$ Crab within a few days of detection Nakajima et al. (2023); Nandi et al. (2024). Radio counterpart of the source, consistent with its optical position, was detected at $5.25$ GHz with VLA and $5$ GHz with ATA, showing a rise in radio flux from $\sim 18$ to $107$ mJy six days after discovery. Recent measurements estimate the source distance as $\sim 2.7$ kpc, with a spin of $\sim 0.98$ and an inclination of $\sim 40^{\circ}$ Mata Sánchez et al. (2024); Peng et al. (2024a). Further, Nandi et al. (2024) reported the detection of evolving type-C QPOs in hard X-rays (up to $\sim 100$ keV) using AstroSat, likely associated with the inner Comptonizing corona. However, NICER observations suggest a vertically elongated, jet-like corona based on the evolution of type-C QPOs Liao et al. (2024).
2.8 GX 339--4
The transient BH-XRB GX $339-4$ was discovered by the OSO77-7- 7 mission in $1973$ Markert et al. (1973) and typically undergoes outbursts every $2-3$ years. Since its discovery, it has remained one of the most extensively studied BH-XRBs due to its rich spectro-temporal variability Méndez & van der Klis (1997); Corbel et al. (2003); Zdziarski et al. (2004); Belloni et al. (2005); Nandi et al. (2012); Aneesha et al. (2024). Detailed timing studies across multiple outbursts have identified distinct temporal features, including type-A, type-B, and type-C QPOs, which are closely associated with the canonical spectral states exhibited by the source (Belloni et al., 2005; Nandi et al., 2012; Mondal et al., 2023; Zhang et al., 2024; Aneesha et al., 2024, and references therein). The black hole mass is constrained to be within $8.28-11.89$ $\rm M_{\odot}$ Sreehari et al. (2019a), while the source distance is estimated as $8.4\pm 0.9$ kpc Parker et al. (2016). Recent studies suggest the source inclination angle between $30^{\circ}-34^{\circ}$, though spin measurements remain highly model-dependent, yielding values ranging from negative to moderate positive spins Zdziarski et al. (2024).
2.9 Swift J151857.0--572147
Swift J151857.0572147151857.0572147151857.0-572147151857.0 - 572147 is a newly discovered Galactic BH-XRB transient, first detected by Swift/BAT during its $2024$ outburst Kennea et al. (2024). Follow-up radio observations with ATCA at $5.5$ and $9$ GHz confirmed the presence of bright radio flares Carotenuto & Russell (2024), generally observed for the BH-XRB transients during the transition from hard to soft spectral states. Subsequent observations with NICER and Insight-HXMT constrained the source distance, inclination, and spin parameter as $\sim 5.8$ kpc, $\sim 2.1^{\circ}$, and $\sim 0.84$, respectively Peng et al. (2024b). Additionally, coherent type-C QPOs were detected with Insight-HXMT in the intermediate state of the outburst, which have been linked to the shock instability model of transonic accretion flow around black holes Chatterjee et al. (2024).
2.10 IGR J17091--3624
IGR J17091362417091362417091-362417091 - 3624 is a galactic BH-XRB transient system, discovered by INTEGRAL/IBIS in $2003$ (Kuulkers et al., 2003). The source exhibits periodic outbursts with a quiescent period of typically four years between the successive outbursts (Capitanio et al., 2009; Krimm et al., 2011; Miller et al., 2016; Katoch et al., 2021). Interestingly, IGR J17091362417091362417091-362417091 - 3624 exhibits several exotic X-ray variability signatures, including the well known “heartbeat” class and therefore, often termed as the twin of GRS $1915+105$ (Capitanio et al., 2012; Radhika et al., 2018; Katoch et al., 2021; Wang et al., 2024; Shui et al., 2024). Despite multiple outbursts over the last two decades, the mass, inclination, and distance of the source are not well constrained due to the null-detection of its optical counterpart. However, several indirect measurements predict the source parameters. Using reflection spectroscopy, the inclination of the source is estimated as $\sim 30^{\circ}-45^{\circ}$ (Xu et al., 2017; Wang et al., 2018). In addition, spectral modeling suggests a black hole of mass $10-12.3$ $M_{\odot}$ likely to be present at the center of IGR J17091362417091362417091-362417091 - 3624 (Iyer et al., 2015; Radhika et al., 2018). Based on the luminosity estimates, the distance of the source is predicted to lie within $11-17$ kpc (Rodriguez et al., 2011).
2.11 MAXI J1744--294
MAXI J174429417442941744-2941744 - 294, located near the Galactic center, is the most recently discovered X-ray transient by MAXI/GSC on January 2, 2025 at a flux level of $\sim 133$ mCrab (Kudo et al., 2025). The energy spectra of the source, obtained from follow-up NuSTAR observations, are found to be well described with the model comprising a thermal disc component of temperature $\sim 0.7$ keV, a power-law component of photon index $\sim 2.3$ and iron line profile at $\sim 6.6$ keV (Mandel et al., 2025). Based on these initial findings, the source is inferred to be a BH-XRB transient observed in the HSS (Mandel et al., 2025). Notably, subsequent NICER monitoring of MAXI J174429417442941744-2941744 - 294 shows an even steeper photon index of $\sim$3 (Jaisawal et al., 2025). Assuming a distance of $8$ kpc, the luminosity of the source is estimated to be $\sim 1.5\times 10^{37}$ erg $\rm s^{-1}$ in $2-10$ keV energy range (Mandel et al., 2025). Further, Grollimund et al. (2025) reported the detection of the radio counterpart of the source at $\sim 1.3$ GHz with MeerKAT.
3 Data reduction
IXPE, NASA’s first space-based low-energy X-ray polarimetry mission, was launched on December 9, 2021 (Weisskopf et al., 2022). It comprises three identical gas-pixel detector units (DUs) designed to measure the polarimetric properties of astrophysical sources using the principle of photo-electron tracking in the $2-8$ keV energy band. For each detected photon, the Stokes parameters (I, Q, U) are determined based on the azimuthal angle of the electric field vector, which is reconstructed by analyzing the photo-electron track within the detector plane. The additive nature of the Stokes parameters enables the summation of individual photon contributions to obtain the resultant I, Q, and U values, which are then used to calculate the polarization degree (PD) and polarization angle (PA).
For the polarimetric analysis, we utilize the cleaned and calibrated level-2 event files from the three DUs of IXPE in the $2-8$ keV energy range. Data reduction is performed using the latest IXPEOBSSIMv31.0.3 software (Baldini et al., 2022), following standard procedures outlined in Kislat et al. 2015; Strohmayer 2017; Kushwaha et al. 2023b; Majumder et al. 2024a, b. The source and background regions are defined as a $60$ arcsec circular region centered at the source coordinates and an annular region with radii between $180$ and $240$ arcsec, respectively (Kushwaha et al., 2023b; Majumder et al., 2024a). The XPSELECT task is used to extract source and background events from these regions. We use XPBIN task with various algorithms, such as PCUBE, PHA1, PHA1Q and PHA1U, to generate the necessary data products for model-dependent polarization as well as spectro-polarimetric studies, respectively. The latest response files provided by the software team [^3] https://ixpeobssim.readthedocs.io/en/latest/ are used during the fitting of Stokes spectra. The IXPE light curves of $2000$ s time bin are extracted from the event files of the DUs using XSELECT.
NICER (Gendreau et al., 2016) is the X-ray mission onboard the International Space Station (ISS), sensitive in the soft X-ray energy band of $0.2-12$ keV and capable of excellent observations for in-depth spectro-temporal studies. The NICER/XTI data is reduced using the standard data extraction software (NICERDASv13444https://heasarc.gsfc.nasa.gov/docs/nicer/nicer_analysis.html) integrated within HEASOFT V6.34555https://heasarc.gsfc.nasa.gov/docs/software/lheasoft/RelNotes_6.34.html with the appropriate calibration database. We use the tool nicerl2 with standard filtering criteria to generate the clean event files. The spectral products are generated using the task nicerl3-spect with background model 3c50 (Remillard et al., 2022). The NICER light curves of desired time resolutions are generated using nicerl3-lc task of the software routine.
We analyze the NuSTAR (Harrison et al., 2013) data using the dedicated software NUSTARDAS [^6] https://heasarc.gsfc.nasa.gov/docs/nustar/analysis/ available in HEASOFT V6.34. We use the task nupipeline to generate the cleaned event files from both FPMA and FPMB instrument onboard NuSTAR. A circular region of $60$ arcsec radii at the source position and away from it is considered to extract the source and background spectra, instrument response and ancillary files, respectively, using the task nuproduct. All the NuSTAR spectra are grouped with $25$ counts per bin to obtain better statistics during the spectral fitting.
AstroSat (Agrawal, 2006) data are extracted using the latest reduction software LaxpcSoftv3.4.4777http://www.tifr.res.in/~astrosat_laxpc/LaxpcSoft.html.The standard procedures of AstroSat/LAXPC data extraction are performed following Antia et al. 2017, 2021; Antia et al. 2022; Sreehari et al. 2019b, 2020; Majumder et al. 2022. We use LAXPC20 data for generating the light curves in different energy ranges for the respective sources except Swift J1727.816131727.816131727.8-16131727.8 - 1613, for which LAXPC10 data in low gain mode was available for analysis (see Nandi et al., 2024, for details). Note that, level-1 data corresponding to single events from all layers of the detector are considered in each observation.
We utilize on-demand [^8] http://maxi.riken.jp/mxondem/ processed MAXI/GSC (Matsuoka et al., 2009) data and scaled map transient analysis data [^9] https://swift.gsfc.nasa.gov/results/transients/ of Swift/BAT (Krimm et al., 2013) for the individual sources in different energy ranges. Note that, MAXI/GSC count rates in $2-20$ keV energy range are converted into the flux unit (Crab) following the relation $1$ Crab $=3.8$ photons $\rm cm^{-2}$ $\rm s^{-1}$ as suggested by the instrument team [^10] http://maxi.riken.jp/top/readme.html. Similarly, for Swift/BAT data in the $15-50$ keV energy range, we obtain the corresponding photon flux [^11] https://swift.gsfc.nasa.gov/results/transients/ in units of mCrab, where $1$ mCrab $=0.000220$ counts $\rm cm^{-2}$ $\rm s^{-1}$.
4 Analysis and Results
In this section, we present the comprehensive analyses, modeling, and observational results of the BH-XRBs under consideration, observed with IXPE. This study incorporates quasi-simultaneous observations from NICER, NuSTAR and AstroSat to enable wide-band spectro-temporal investigations. Furthermore, we explore the polarimetric properties with IXPE and examine the spectro-polarimetric correlations for these sources.
4.1 Temporal Analysis
Figure 2: Results of temporal (top panel) and spectral (bottom panel) analyses of ten BH-XRBs in different spectral states obtained using quasi-simultaneous IXPE, NICER, NuSTAR and AstroSat observations. Top panels: PDS obtained from NICER ($0.5-10$ keV) and/or NuSTAR/AstroSat observations in different energy ranges are depicted for the respective sources in various epochs. The power spectrum of GX $339-4$ from NICER observation is rescaled by a factor of $5$. Bottom panels: Energy spectra of the sources from NICER and/or NuSTAR data are shown during the respective epochs. PDS and energy spectra in LHS, HIMS, SIMS, HSS and SPL states are presented using red, blue, magenta, orange and green colors, respectively. PDS from NuSTAR and AstroSat data are plotted using circle and triangle, irrespective of the spectral states. See the text for details.
4.1.1 Multi-mission Monitoring
We study the variability properties of each source using multi-mission data during the epochs of IXPE observations. In Fig. 1, we present the count rate obtained with MAXI/GSC ($2-20$ keV) and Swift/BAT ($15-50$ keV) monitoring in panels (a) and (b) along with the variation of hardness ratio ($HR$) from MAXI/GSC in panel (c), respectively, for all sources under consideration. Note that, flux values associated with the count rates in soft (hard) energy bands with MAXI/GSC (Swift/BAT) are also mentioned in units of mCrab in the $y$-axis (right side) of the respective panels. Further, radio flux densities obtained from different observational campaigns, such as RATAN (Veledina et al., 2024a; Ingram et al., 2024), E-MERLIN (Williams-Baldwin et al., 2023), VLITE (Peters et al., 2023) and AMI-LA (Steiner et al., 2024) near the IXPE epochs, are shown in panel (d) for the respective sources based on their availability. We observe that all sources show marginal variation in both MAXI/GSC and Swift/BAT count rates over the entire IXPE exposure for a given epoch. Hence, in Fig. 1a-c, we present the average values of the respective quantities over the entire IXPE exposure of each epoch.
Interestingly, for Swift J1727.816131727.816131727.8-16131727.8 - 1613, we observe that both MAXI/GSC and Swift/BAT flux decrease as its outburst progresses. The $HR$ also sharply decreases and anti-correlates with the accompanied radio flux density (see Fig. 1c-d). The variation of both low and high energy fluxes along with $HR$ in presence of radio ejection suggest that the source evolves through spectral state transitions. In particular, we notice exceptionally bright ($\gtrsim 3$ Crab) intermediate states ($HR\sim 0.1-0.5$, E1-E5) followed by relatively faint ($\sim 30$ mCrab) softer states ($HR\sim 0.04-0.1$, E6-E7) before the appearance of dimmed ($\sim 100$ mCrab) hard state ($HR\sim 0.3$, E8) (see Table 2). Moreover, these variability features resemble the ‘canonical’ state transition as HIMS $\rightarrow$ SIMS $\rightarrow$ HSS $\rightarrow$ LHS (decay), commonly exhibited by BH-XRB transients (see Nandi et al., 2012, for details). Similarly, for Cyg X11-1- 1, we observe a marginal increase in the MAXI/GSC count rate and a corresponding decrease in the Swift/BAT count rate during epochs E3 to E7. In the subsequent observations (E8 to E11), the MAXI/GSC count rate decreases while the Swift/BAT count rate rises. In addition, $HR$ shows a significant drop from $0.4$ to $0.1$ followed by an increase in values exceeding $0.4$. This trend possibly indicates the spectral state transitions between LHS (E1-E2, E8-E11) and HSS (E3-E7, E12), respectively (see Table 1).
However, only marginal variations are seen in both low- and high-energy count rates for several sources, such as LMC X11-1- 1, LMC X33-3- 3, 4U $1957+115$, GX $339-4$, Swift J151857.0572147151857.0572147151857.0-572147151857.0 - 572147 and MAXI J174429417442941744-2941744 - 294. For all of them, the hardness ratio remains low ($HR\lesssim 0.1$) and these sources are also detected at flux levels below approximately $100$ mCrab. These possibly suggest that the sources are likely in softer spectral states. In contrast, Cyg X33-3- 3 exhibits a clear sequence of state transitions, evolving from the LHS to the HSS, then to the SIMS, and back to HSS. This evolution is accompanied by moderate changes in count rates and hardness ratio, and is likely associated with strong radio flares ($\sim 100$ mJy) (see Table 1, Fig. 1, and Veledina et al. 2024a). For 4U $1630-47$, HSS (Kushwaha et al., 2023b) and SPL (Rodriguez Cavero et al., 2023) states are observed with reasonable variations in the count rates and $HR$. It is worth noting that IGR $17091-3624$ is observed in the LHS at $\sim 1$ Crab MAXI/GSC flux with $HR\sim 0.4$, very similar to the LHS observations of Cyg X11-1- 1.
4.1.2 Power Density Spectra
We investigate the power density spectra (PDS) of all sources in different spectral states using NICER observations (see Table 1) in $0.5-10$ keV energy range. In addition, PDS properties in hard X-rays ($>10$ keV) are also examined using available NuSTAR and AstroSat observations of the sources. We generate 0.01 s time binned light curves and compute the PDS up to Nyquist frequency of $50$ Hz with $1024$ newbins per interval using powspec inside HEASOFT V6.34. The individual power spectra are averaged out to get the resultant PDS in units of $\rm rms^{2}/Hz$ which is geometrically rebinned in the frequency space with a factor of $1.03$ (see Belloni et al., 2002, 2005; Sreehari et al., 2019b; Majumder et al., 2022). In the upper panels of Fig. 2, we present the PDS of the respective sources in different spectral states. We model each PDS with a combination of constant and multiple Lorentzian components. Further, we compute the total percentage rms amplitude ($rms_{\rm tot}\%$) as a measure of variability in $0.1-50$ Hz and tabulate it in Table 3.
Interestingly, we observe that the PDS in several spectral states exhibit distinct variable features. In particular, PDS of Cyg X11-1- 1 in the LHS show significant variability with $rms_{\rm tot}\%\sim 31.1-52.4$ ($0.5-10$ keV) and a clear power spectral break at $\sim 1.5$ Hz with NICER observations. Further, power spectral properties in hard X-rays ($3-80$ keV) with quasi-simultaneous AstroSat observations exhibit similar break frequencies ($\sim 1.5$ Hz) with $rms_{\rm tot}\%\sim 37$ in LHS (see Table 1 and Fig. 2a). We also observe marginal (significant) PDS variability of $rms_{\rm tot}\%\sim 9.1$ ($33$) with low (high) energy NICER (AstroSat) observations of Cyg X11-1- 1 in the HSS. Note that similar variability properties including power spectral breaks were also reported by Kushwaha et al. (2021) in LHS and HSS of Cyg X11-1- 1.
For Swift J1727.816131727.816131727.8-16131727.8 - 1613, the presence of type-C QPO feature at $\sim 1.4$ Hz is observed during HIMS (epoch 1, hereafter E1) in $0.5-100$ keV with NICER, NuSTAR and AstroSat quasi-simultaneous observations (see also Nandi et al. 2024; Liao et al. 2024). The evolution of the QPO frequency up to $\sim 6.7$ Hz is seen with NICER in the later epochs (E2-E5). However, a power distribution, similar to Cyg X11-1- 1 (LHS), is observed in the LHS of Swift J1727.816131727.816131727.8-16131727.8 - 1613 without the presence of power spectral break, whereas marginal variability with $rms_{\rm tot}\%\sim 1.4$ is noticed during HSS (E6-E7).
The remaining sources (4U $1630-47$, LMC X11-1- 1, LMC X33-3- 3, 4U $1957+115$, Cyg X33-3- 3 and Swift J151857.0572147151857.0572147151857.0-572147151857.0 - 572147) in SIMS and HSS display weak variability signatures mostly represented by constant noise distribution and occasionally marginal power variation at lower frequencies. Conversely, during the recurrent outburst of GX $339-4$, a sharp type-B QPO peak at $\sim 4.5$ Hz ($rms_{\rm QPO}\%\sim 10.2$) and harmonic feature at $\sim 9.5$ Hz along with weak broadband noise ($rms_{\rm tot}\%\sim 15.4$) are observed in SIMS (see Fig. 2 and Aneesha et al. 2024). Note that, the QPO features with enhanced $rms_{\rm QPO}\%$ from $3.76$ ($0.5-10$ keV) to $12.6$ ($3-80$ keV) are observed with NICER and AstroSat, respectively. For IGR J17091362417091362417091-362417091 - 3624, a sharp type-C QPO at $\sim 0.2$ Hz ($rms_{\rm QPO}\%\sim 15$) along with a weak harmonic at $\sim 0.4$ Hz are observed during the onset phase of the recent outburst (see Fig. 2).
Moreover, the overall power spectral properties of the sources, including break frequencies, noise distribution, and the detection of type-C and type-B QPO features confirm the presence of distinct spectral states during the respective IXPE campaigns. These findings further corroborate the results indicating distinct spectral states, obtained from the multi-mission monitoring of the sources (see §4.1.1).
4.2 Wide-band Spectral Analysis
4.2.1 Spectral Modeling
Figure 3: Variation of the data to model ratio of the HSS/HIMS spectra of Cyg X11-1- 1, GX $339-4$ and Swift J1727.816131727.816131727.8-16131727.8 - 1613 fitted with model combination M1 and M3. The improvement in the respective fits by incorporating the reflection model component over M1 is depicted in panels (b), (d) and (f). See the text for details.
[…]
4.2.2 Alternative Spectral Models
We aim to explore the possible alternative disc-corona geometries in explaining the high-energy tail of the observed spectra. To achieve this, we use the comptt (Titarchuk, 1994) component as a proxy for thcomp in modeling thermal Comptonization. Note that, the ‘geometry switch’ selected inside comptt model as ‘$\leq 1$’ enables a disc-like geometry for the Comptonizing corona. Considering this, we proceed in the spectral fitting of all sources with the model combination constant×$\times$×Tbabs×$\times$(diskbb + comptt) (hereafter M2). Interestingly, we find that the model M2 provides a fit equally as good as M1 for all observations under consideration except a few observations of Cyg X11-1- 1 (E6 and E7), Swift J1727.816131727.816131727.8-16131727.8 - 1613 (E4 and E5) and GX $339-4$ (E1). This leads to a possible degeneracy between the selected models in describing the observed spectra and the geometry of the Comptonizing region.
[…]
4.2.3 Spectral Properties
The spectral energy distribution of the individual sources are generally described by the model combination comprising thermal Comptonization, standard disc and occasionally reflection components. Although several models developed based on different coronal geometries are found to describe the observed spectra satisfactorily, we examine the spectral states of the respective sources using model M1 and M3 from the extracted spectral parameters. We present the best-fitted and estimated model parameters for each source in Table 3. The variation of the spectral parameters over different observational epochs indicates the presence of distinct spectral states of the sources under consideration.
Figure 4: Confidence contours ($3\sigma$) in the PD–PA plane obtained from the model-independent polarization measurements using $2-8$ keV IXPE data, combining all DUs across various spectral states of the sources. The epoch of each observation is indicated in parentheses next to the corresponding spectral state. For Swift J151857.0572147151857.0572147151857.0-572147151857.0 - 572147, LMC X11-1- 1 and MAXI J174429417442941744-2941744 - 294, null-detection of polarization is observed, whereas all other sources exhibit significant polarization. The inset highlights a zoomed-in region of the PD–PA space for better clarity. See the text for details.
We find that the effects of Comptonization are reasonably high during the LHS of Cyg X11-1- 1 with covering fraction $cov_{\rm frac}\sim 0.61_{-0.04}^{+0.04}-0.77_{-0.02}^{+0.02}$, where $cov_{\rm frac}$ measures the fraction of seed photons that are being up-scattered by thermal electrons of the corona. In addition, a high Comptonized flux contribution of $73-80\%$ is also observed in the spectra (see Table 3). Indeed, such a high $cov_{\rm frac}$ results in low photon index ($\Gamma_{\rm th}\lesssim 1.64_{-0.02}^{+0.02}$) as observed during the LHS observations of Cyg X11-1- 1. In the HIMS and LHS of Swift J1727.816131727.816131727.8-16131727.8 - 1613, we find $cov_{\rm frac}\sim 0.35_{-0.01}^{+0.01}-0.82_{-0.02}^{+0.02}$ and $\Gamma_{\rm th}\sim 1.71_{-0.02}^{+0.03}-2.11_{-0.04}^{+0.05}$ with up to $\sim 64\%$ Comptonized emission. Similarly, LHS of IGR J17091362417091362417091-362417091 - 3624 which is dominated by the Comptonized emission ($\sim 87\%$ flux contribution) yields harder ($\Gamma_{\rm th}\sim 1.62_{-0.01}^{+0.02}$) spectral characteristics. On the other hand, the HSS/SIMS spectra of Cyg X11-1- 1 and Swift J1727.816131727.816131727.8-16131727.8 - 1613 results in low Comptonization with $cov_{\rm frac}$ in the range $0.005_{-0.002}^{+0.002}-0.38_{-0.03}^{+0.04}$ and thermal disc emission with $39-98\%$ flux contribution along with a steeper spectral index ($\Gamma_{\rm th}\sim 2.32_{-0.04}^{+0.06}-5.14_{-0.39}^{+0.36}$). Moreover, HSS with $76-98\%$ disc contribution are observed for LMC X11-1- 1, 4U $1630-47$, LMC X33-3- 3 and 4U $1957+115$ with negligible Comptonized emission ($cov_{\rm frac}<0.1$) as shown in Table 3. We find that GX $339-4$ (Swift J151857.0572147151857.0572147151857.0-572147151857.0 - 572147) exhibits intermediate spectral characteristics with $cov_{\rm frac}\sim 0.34$ ($0.21$) and $\Gamma_{\rm th}\sim 2$ ($2.5$) with $\sim 50\%$ ($78\%$) disc emission during SIMS (HSS).
The reflection spectroscopy of Cyg X11-1- 1, Swift J1727.816131727.816131727.8-16131727.8 - 1613 and GX $339-4$ with model M3 during the HSS/SIMS results in disc ionization parameter ($\log\xi$) in the range of $3.31_{-0.05}^{+0.05}-4.29_{-0.21}^{+0.14}$ erg cm $\rm s^{-1}$, iron abundance ($A_{\rm fe}$) between $\sim 4.61_{-0.31}^{+0.36}$ and $6.35_{-1.46}^{+2.08}$ $A_{\odot}$, and reflection fraction ($R_{\rm f}$) ranging $\sim 0.59_{-0.11}^{+0.05}$ to $3.16_{-0.33}^{+0.35}$. Here, $A_{\odot}$ represents the solar iron abundance. The best-fitted parameters of the relxill component obtained from model M3 are tabulated in Table 4, whereas the parameters associated with other spectral components (diskbb and thcomp) of M3 are presented in Table 3 for the respective sources. Notably, a high ionization ($\log\xi\sim 4$) observed in HSS/SIMS generally suggests the reflection of significant coronal emission producing strong reflection signatures.
In addition, we find that the model component relxilllp (Dauser et al., 2013) also delineates the reflection features assuming the primary source emissivity profile from a lamppost corona. Moreover, another reflection flavor relxillCp, in which the spectrum of the coronal emission is computed from the nthComp (Zdziarski et al., 1996) model of spherical geometry, also fits the data for Cyg X11-1- 1 (E6 and E7) and Swift J1727.816131727.816131727.8-16131727.8 - 1613 (E4). Therefore, it is evident that the reflection spectroscopy also results in degeneracy among several models relying on contemporary coronal geometries.
Overall, the wide-band spectro-temporal analysis (see §4.1 and §4.2) confirms the presence of distinct spectral states in BH-XRBs, but falls short of resolving the inherent degeneracy among the possible coronal geometries. This highlights the importance of incorporating X-ray polarimetric observations into the study of BH-XRBs. Subsequently, we explore how polarization measurements can offer deeper insights into this issue.
4.3 Polarization Measurements
We study the in-depth polarimetric properties of the eleven BH-XRBs observed with IXPE during multiple campaigns. In general, the polarization measurements are performed using two distinct methods. The first one relies on the model-independent approach based on the analysis of the polarization cube (Kislat et al., 2015) using the dedicated python package IXPEOBSSIMv31.0.3121212https://ixpeobssim.readthedocs.io/en/latest/ (Baldini et al., 2022). This provides the polarimetric properties by computing the polarization degree (PD) and polarization angle (PA) from the normalized Stokes parameters without any prior assumption on the emission mechanism. The alternative method depends on the simultaneous modeling of the I, Q, and U Stokes spectra in XPSEC. In this approach, the source intensity spectrum (I) is modeled using the appropriate physical components responsible for the emission mechanism and a polarization model accounting for the Q and U spectra (Strohmayer, 2017). Accordingly, we carry out both model-independent and model-dependent analyses to ensure the consistency of the polarimetric measurements.
We perform model-independent polarization cube analysis using XPBIN tool inside the software routine with PCUBE algorithm. Following Kushwaha et al. 2023b; Majumder et al. 2024a, b, we consider all events from the three DUs of IXPE in the analysis. Subsequently, the normalized Q and U Stokes parameters, the polarization degree $({\rm PD}=\sqrt{(Q/I)^{2}+(U/I)^{2}})$ and polarization angle $({\rm PA}=\frac{1}{2}\tan^{-1}(U/Q))$ are computed in $2-8$ keV energy band for all observations of the respective sources. Following the guidelines of IAU [^13] https://aas.org/posts/news/2015/12/iau-calls-consistency-use-polarization-angle (Contopoulos & Jappel, 1974), we adopt the convention that PA increases counter-clockwise from North to East direction in the sky. We also calculate the minimum detectable polarization at $99\%$ confidence level ($\rm MDP_{99}\%$) and the significance (SIGNIF in units of $\sigma$) for each measurement. It may be noted that an observed $\rm PD>MDP_{99}\%$ indicates a polarization level that would arise from random statistical fluctuations with only $1\%$ chance probability (Kislat et al., 2015). All the calculated polarimetric parameters and the measurement statistics are tabulated in Table 5. In Fig. 4, we present $3\sigma$ confidence contours associated with the measurements of PD and PA for the respective sources in different spectral states.
We find significant degree of polarized emission ($\rm PD>\rm MDP_{\rm 99}\%$) in all sources except LMC X11-1- 1, Swift J151857.0572147151857.0572147151857.0-572147151857.0 - 572147 and MAXI J1744--294. More precisely, we find that PD varies within $1.22\pm 0.35\%-21.41\pm 0.41\%$ including all sources in different spectral states within $2-8$ keV energy band. We note that Cyg X33-3- 3 shows the maximum degree of polarization, whereas GX $339-4$ exhibits the lowest PD ($>\rm MDP_{99}\%$) among all sources under consideration. Further, a significant variation of PD is observed over different spectral states of the sources. For example, Cyg X11-1- 1 manifests a maximum PD of $\sim 4.8\%$ in the LHS, which decreases to $\sim 1.4\%$ in the HSS. Interestingly, a similar variation of PD is observed over several spectral states during the outburst phase of Swift J1727.816131727.816131727.8-16131727.8 - 1613 in which the polarization degree drops down to $\sim 0.4\%$ (HSS) from $\sim 4.7\%$ (HIMS) and again increases to $\sim 3.2\%$ in the LHS. Most astonishingly, 4U $1630-47$ exhibits exceedingly high PD in the HSS ($\sim 8.3\%$) and SPL ($\sim 6.8\%$), respectively. Moreover, we find an unprecedentedly high PD of $\sim 9\%$ in the LHS of the recent outburst of IGR J17091362417091362417091-362417091 - 3624, which is the maximum reported for any confirmed BH-XRB observed with IXPE to date (see also Ewing et al. 2025). In contrast, 4U $1957+115$ and LMC X33-3- 3 are found to show $\sim 1.9\%$ and $\sim 2.4-3\%$ polarization in the HSS, respectively. It may be noted that, for GX $339-4$, significant PD ($\sim 1.2\%$, E1) is obtained in $3-8$ keV energy range (see also Mastroserio et al. 2025), whereas a null-detection is seen considering the entire IXPE energy band of $2-8$ keV. Moreover, we find that the PA of the sources in different spectral states shows marginal variation (see Table 5).
  • Null-detection of Polarization: It is worth mentioning that we re-confirm the null-detection of polarized emission ($\rm PD<\rm MDP_{\rm 99}\%$) for a few sources under consideration. In particular, LMC X11-1- 1 remains the first source that fails to manifest significant polarization with IXPE (see Table 5 and Podgorný et al. 2023). Further, Swift J151857.0572147151857.0572147151857.0-572147151857.0 - 572147 exhibits a PD of $\sim 0.25\%$ in $2-8$ keV which remains significantly low from the corresponding $\rm MDP_{\rm 99}\sim 0.78\%$, confirming a null-detection. This finding contradicts the polarization detection with PD $\sim 1.3\%$ as reported by Mondal et al. (2024). This discrepancy possibly arises as the present analysis is carried out using the latest version of the software IXPEOBSSIMv31.0.3, fixing the bugs related to the visualization of polarization cubes in the previous releases [^14] https://ixpeobssim.readthedocs.io/en/latest/release_notes.html. We emphasize that our results are consistent with the recent polarization measurements of Swift J151857.0572147151857.0572147151857.0-572147151857.0 - 572147 Ling et al. (2024). Further, the null-detection of polarization is also observed in a few observations of Swift J1727.816131727.816131727.8-16131727.8 - 1613 (E6, E7), GX $339-4$ (E2) and LMC X33-3- 3 (E3). In addition, for the first time, we report the null-detection of X-ray polarization with PD ($\sim 0.7\%$) $<\rm MDP_{\rm 99}(1.3\%)$ in the recently discovered BH-XRB candidate MAXI J174429417442941744-2941744 - 294 during HSS.
4.4 Energy-dependent Polarization
Figure 5: Variation of polarization degree (PD) with energy. In each panel, source name and spectral state (observed during different epoch) are marked. The histograms in colors represent the minimum detectable polarization at $99\%$ confidence ($\rm MDP_{99}\%$). See the text for details.
We further explore energy-dependent polarization properties of the selected sources that confirm the presence of significant polarized emission in $2-8$ keV energy range of IXPE. For each source except 4U $1957+115$, IGR J17091362417091362417091-362417091 - 3624 and GX $339-4$, we divide $2-8$ keV energy range into six linearly spaced energy bins and estimate polarization using the model-independent approach (see also Svoboda et al. 2024a; Veledina et al. 2023). For 4U $1957+115$, IGR J17091362417091362417091-362417091 - 3624 and GX $339-4$, only four energy bins are considered for the energy-dependent study to ensure relatively improved photon statistics per energy bin (see also Marra et al. 2024). Further, we estimate the significance of the energy variation of PD over these bins following Krawczynski et al. (2022); Majumder et al. (2024b). In doing so, we consider PD in $2-8$ keV energy range (see Table 5) as the null hypothesis against which the significance of energy variations in PD is computed for each epoch of a given source. With this, following Majumder et al. (2024b), we calculate the probability of rejecting the null hypothesis using $(N-1)$ degrees of freedom, taking into account the variation of PD over $N$ energy bins for the respective cases. This essentially indicates the confidence level at which the PD variation is significant over different energy bins. The obtained significance levels of the energy variation of PD are mentioned in Table 5 for the respective cases. We observe that this significance varies from $\lesssim 1\sigma$ to $>4\sigma$ across different epochs of the sources in various spectral states. In Fig. 5, we present the variation of PD with energy over different observational epochs using filled circles in the respective panels. The histograms in colors represent the $\rm MDP_{99}\%$ associated with the measurements in the selected energy bins for the corresponding epochs. Note that, for a given source, when multi-epoch observations are available, we present the results for two epochs of different spectral states only, except LMC X--3.
Interestingly, we observe that Cyg X11-1- 1 and Swift J1727.816131613-1613- 1613 show similar energy variation of PD within $\sim 2-6\%$ in different spectral states (LHS, HSS, HIMS and SIMS). Moreover, for Cyg X11-1- 1, the variation in PD with energy remains most significant at a confidence level exceeding $4\sigma$ in HSS, whereas it remains $2.8\sigma$ for Swift J1727.816131727.816131727.8-16131727.8 - 1613 in HIMS, when all observational epochs are considered (see Fig. 5). Note that, PD $<\rm MDP_{99}\%$ at higher energy bins in a few epochs of several sources, indicates a null-detection. Interestingly, unlike Cyg X11-1- 1 and Swift J1727.816131727.816131727.8-16131727.8 - 1613, a weak energy dependency of PD is noticed for IGR $17091-3624$ in the LHS, for which PD corresponding to two out of four energy bins remain below the $\rm MDP_{99}\%$ (see Fig. 5). We find a significant increase in PD with energy up to $\sim 11\%$ in HSS ($>4\sigma$) and SPL ($>4\sigma$) state of 4U 16304747-47- 47 (see Fig. 5). Similarly, distinct variations of PD over several energy bins are observed for Cyg X33-3- 3 during LHS ($>4\sigma$) and SIMS ($>4\sigma$). LMC X33-3- 3 shows constant polarization over the initial three energy bins with a sudden jump in PD up to $\sim 7.8\%$ at $\sim 5.4$ keV in both epochs (see Fig. 5). Further, we only observe a marginal variation ($1.7\sigma$) of PD in 4U 1957$+115$ up to $\sim 3\%$. Interestingly, significant ($3.3\sigma$) energy variation of polarization degree is observed for GX $339-4$. Although, PD $<\rm MDP_{99}\%$ is obtained in all four energy bins (see Fig. 5). Overall, we state that a significant ($>3\sigma$) variation of polarization degree with energy is observed only for Cyg X11-1- 1, 4U 16304747-47- 47 and Cyg X33-3- 3, whereas a moderate ($2\sigma$ to $3\sigma$) variation is seen for Swift J1727.816131613-1613- 1613 and LMC X33-3- 3. However, Veledina et al. 2024a, b reported an energy-independent variation of PD across 12 energy bins for Cyg X--3. We mention that the strong ($>4\sigma$) energy dependence of PD observed in our analysis possibly resulted due to relatively small uncertainties obtained in PD estimates over the energy bins and a significant drop in PD around the unpolarized iron line complex at $\sim 6.4$ keV. It is worth mentioning that the energy variation of PA for all sources under consideration remains statistically insignificant ($<2\sigma$). However, we note that PA drops from $-30^{\circ}$ to $-60^{\circ}$ at $\sim 6$ keV in 4U $1957+115$, which is marginally significant at $1.6\sigma$ confidence level.
4.5 Spectro-polarimetric Modeling
We explore the model-dependent spectro-polarimetric properties of the sources from the simultaneous fitting of the I, Q and U Stokes spectra of IXPE in $2-8$ keV energy range. Due to the calibration issues Krawczynski et al. (2022); Veledina et al. (2024a); Steiner et al. (2024), only DU1 data is considered for the spectro-polarimetric study of Cyg X11-1- 1, Cyg X33-3- 3 and Swift J1727.816131613-1613- 1613, whereas data combining all the DUs are used for the remaining sources. For simplicity, we first adopt the model combination M4: Tbabs×$\times$×polconst×$\times$(diskbb + powerlaw) for the spectro-polarimetric fitting. Here, polconst represents a constant polarization model, where the degree and angle of polarization are treated as the model parameters. The best-fit results yield $\chi_{\rm red}^{2}$ ranging from $0.91-1.79$ across all selected sources, with PD varying between $\sim 1.16-13.06\%$ and PA spanning from $-89.13^{\circ}$ to $85.15^{\circ}$. Note that the spectro-polarimetric modeling of Swift J1727.816131727.816131727.8-16131727.8 - 1613 does not constrain the hydrogen column density, likely due to low-energy threshold of IXPE ($\sim 2$ keV; see Veledina et al., 2023; Ingram et al., 2024).
Next, we investigate the energy-dependent polarization properties using model-dependent approach for all the sources that demonstrate the signature of X-ray polarization. In doing so, we replace the polconst component with the energy-dependent polarization model polpow [^15] https://heasarc.gsfc.nasa.gov/xanadu/xspec/manual/node218.html. Empirically, polpow assumes a power-law variation in the polarization parameters of the form PD($E$) $=A_{\rm norm}\times E^{-A_{\rm index}}$ (in fraction) and PA($E$) $=\psi_{\rm norm}\times E^{-\psi_{\rm index}}$ (in ) while modeling the Q and U Stokes spectra. We find that the best-fitted $\psi_{\rm index}$ remains consistent with zero within its $1\sigma$ errors and hence, we freeze it to zero during the spectral fitting of all observations. Accordingly, we obtain the best spectral fit for all sources with $\chi_{\rm red}^{2}\sim 0.97-1.46$. Furthermore, integrating PD($E$) in $2-8$ keV energy range using the best-fitted model parameters ($A_{\rm index}$ and $A_{\rm norm}$), we obtain PD as $3.23-14.33\%$. Moreover, with $\psi_{\rm index}$ frozen to zero, the polarization angles are obtained as the best-fitted $\psi_{\rm norm}$ values that vary from $-88.92^{\circ}$ to $85.11^{\circ}$. Note that, in a few observations of Cyg X11-1- 1, 4U 1957$+115$ and GX $339-4$ (see Table 6), the data is unable to constrain the parameters of polpow component even at $1\sigma$ level, and hence, we refrain from the modeling with polpow for these observations. In Table 6, we present the best-fitted model parameters obtained from the spectro-polarimetric study of all sources. We find that the model-dependent polarimetric results are in agreement with the findings of the model-independent study except Cyg X33-3- 3, confirming the robustness of the polarization measurements for the individual sources. However, for Cyg X--3, we obtain significantly low PD compared to the results of the model-independent study, likely due to multiple low-energy unpolarized emission lines including strong iron complex around $\sim 6.4$ keV (see Veledina et al. 2024a), dominating the Stokes spectra.
4.6 Spectro-polarimetric Correlation
Figure 6: Variation of PD ($2-8$ keV), obtained from model-independent analysis, as a function of the disc-to-Comptonized flux ratio ($F_{\rm ratio}$) derived from wide-band spectral modelling. The color bar represents the covering fraction ($cov_{\rm frac}$), while each data point is labeled with the corresponding spectral state and observational epoch. Data points within the shaded region (light blue) are used to compute the Pearson correlation coefficient. Downward arrows indicate null-detections of polarization for the respective sources. See the text for details.
We examine the spectro-polarimetric correlation properties for each source based on the results obtained from wide-band spectral modeling using model M1 (see §4.2) and the polarimetric measurements (see §4.3). Specifically, we explore the correlation between the observed PD obtained from model-independent study, and best-fitted spectral parameters, namely the ratio of disc to Comptonized spectral flux ($F_{\rm ratio}$) in $1-100$ keV energy range, and the covering fraction ($cov_{\rm frac}$). In Fig. 6, we present the variation of PD with $F_{\rm ratio}$ (see Table 3) using different markers, where the covering fraction ($cov_{\rm frac}$, in percent) is represented by the colors of respective markers. The color bar on the right indicates the range of $cov_{\rm frac}$. The spectral state and the corresponding epoch (in parentheses) for each observation are marked in the figure.
[…]
Furthermore, we find a strong correlation (anti-correlation) between PD and $cov_{\rm frac}$ ($F_{\rm ratio}$) with $\rho\sim 0.9$ ($-0.9$) for Swift J1727.816131727.816131727.8-16131727.8 - 1613. In particular, a sharp drop in PD is noticed from $\sim 4.7\%$ (HIMS) to $\sim 0.4\%$ (HSS) (see also Svoboda et al. 2024b) during which $F_{\rm ratio}$ increases from $\sim 0.7$ to $\sim 66$ and $cov_{\rm frac}$ noticeably reduces down to $13\%$ from a maximum value of $82\%$. A similar trend is observed for Cyg X11-1- 1 with $\rho=0.9$ ($-0.9$) for the respective positive (negative) correlations between PD and $cov_{\rm frac}$ ($F_{\rm ratio}$). As before, a low (high) $F_{\rm ratio}\sim 0.3$ ($11.5$) and high (low) $cov_{\rm frac}\sim 61-77\%$ ($0.5-31\%$) are observed in the LHS (HSS) of Cyg X11-1- 1 with PD $\sim 3.7-4.8\%$ ($1.4-2.1\%$) (see Fig. 6). Moreover, we also find a similar behavior in the LHS of IGR J17091362417091362417091-362417091 - 3624, for which a very high PD ($\sim 9\%$) is seen with $cov_{\rm frac}\sim 87\%$ and $F_{\rm ratio}\sim 0.2$. These findings suggest that a significant contribution in the polarization degree comes from the Comptonized emission for Swift J1727.816131727.816131727.8-16131727.8 - 1613, Cyg X11-1- 1 and IGR J17091362417091362417091-362417091 - 3624 in the HIMS and/or LHS. Furthermore, we notice a low polarization degree (PD $\lesssim 2\%$) for all the sources except 4U 16304747-47- 47 and LMC X33-3- 3 in the HSS/SIMS with $F_{\rm ratio}\gtrsim 2$ and $cov_{\rm frac}\lesssim 30\%$. Interestingly, 4U 16304747-47- 47 being the source of extreme soft nature ($F_{\rm ratio}\sim 52$, $cov_{\rm frac}\sim 5\%$) manifests a high polarization degree of $\sim 8.3\%$ in the HSS, whereas it reduces to $6.8\%$ in the SPL state with $F_{\rm ratio}\sim 3.6$ and $cov_{\rm frac}\sim 38\%$. We observe that 4U $1957+115$ (GX $339-4$) shows PD $\sim 1.2\%$ ($1.9\%$) with $cov_{\rm frac}\sim 2.7\%$ ($34\%$) and $F_{\rm ratio}\sim 18$ ($2.4$) in the HSS (SIMS). Moreover, LMC X33-3- 3 exhibits moderate polarization with PD in the range of $2.4-3\%$ for $F_{\rm ratio}\gtrsim 4$ and $cov_{\rm frac}\sim 0.5\%$ during the HSS.
5 Discussion
In this study, we present a comprehensive spectro-polarimetric investigation of eleven BH-XRBs using quasi-simultaneous observations from IXPE, NICER, AstroSat and NuSTAR. IXPE data are analyzed to examine the polarization properties in $2-8$ keV energy range, while combined NICER, AstroSat and NuSTAR observations provide insights into the spectro-temporal behavior across the wide energy range ($0.5-100$ keV). Subsequently, the spectro-polarimetric findings are used to investigate the accretion-ejection dynamics and geometry surrounding the sources.
5.1 Spectro-temporal Properties and Spectral States
Spectro-temporal studies using NICER, NuSTAR and AstroSat observations in $0.5-100$ keV energy range impart distinct variability properties in various spectral states of the sources (see §4.1 and §4.2). The obtained results are in line with the spectro-temporal characteristics generally seen in the respective spectral states of BH-XRBs (Belloni et al., 2005; Remillard & McClintock, 2006; Kalemci et al., 2006; McClintock et al., 2009; Nandi et al., 2012; Radhika & Nandi, 2014; Radhika et al., 2016; Sreehari et al., 2019b; Connors et al., 2019; Méndez et al., 2022; Athulya et al., 2022; Aneesha et al., 2024; Shui et al., 2024; Li et al., 2025). These findings reinforce the presence of distinct spectral states as observed during the multi-mission campaigns of the sources.
In particular, we observe a high fractional variability amplitude $rms_{\rm tot}\%\sim 30-52$ in the PDS of LHS, whereas significantly low variability with $rms_{\rm tot}\%\sim 6-14$ are seen in the intermediate states (HIMS and SIMS) of the sources (see Table 3). Negligible power distribution at lower frequencies results in marginal variability ($rms_{\rm tot}\%\sim 1-10$) for the sources in the HSS (see Table 3 and Fig. 2). In addition, we find the presence of a strong type-C QPO feature at $\sim 1.4$ Hz (E1), which evolves to $\sim 6.7$ Hz (E2-E5), during the HIMS and SIMS of Swift J1727.816131613-1613- 1613 over a wide energy range of $0.5-100$ keV with NICER, NuSTAR and AstroSat observations (see Fig. 2). The origin of this QPO is attributed to a strong Comptonizing corona surrounding the central source (Chakrabarti & Titarchuk, 1995; Chakrabarti et al., 2008; Nandi et al., 2024); however, based on the correlation between the evolution of QPO frequency and quasi-simultaneous radio jet ejections, Liao et al. (2024) suggest a jet-like, vertically elongated corona for Swift J1727.816131727.816131727.8-16131727.8 - 1613. A similar type-C QPO at $\sim 0.2$ Hz is detected in LHS of IGR J17091362417091362417091-362417091 - 3624 (see Fig. 2). Moreover, we also observe a type-B QPO in the disc-dominated SIMS of GX $339-4$ in the $0.5-80$ keV energy range from NICER and AstroSat data. Similar type-B QPOs have been observed during $2024$ and earlier outbursts of GX $339-4$, indicating the presence of a compact corona located very close to the black hole, accompanied by dominant disc emission (Nandi et al., 2012; Aneesha et al., 2024).
Admittedly, the wide-band energy spectral analysis ($0.7-70$ keV) reveals distinct spectral states of each source. The observed spectra of all sources are well described by a standard accretion disc component (diskbb) combined with thermal Comptonization of soft photons by a spherical corona (thcomp) (see Table 3 and Fig. 2). The spectral properties exhibit a strong correlation with the spectral states of the sources. In particular, during LHS and HIMS, the spectra are dominated by Comptonized emission, contributing $\sim 64-87\%$ of the total flux, accompanied by a high covering fraction ($cov_{\rm frac}\sim 61-87$). In contrast, the HSS and SIMS are characterized by a prominent disc component, contributing up to $39-98\%$ with marginal Comptonization effects ($cov_{\rm frac}\sim 0.5-38\%$) across all sources. We observe the present findings to be broadly consistent with the previously reported results for the respective sources. For example, Svoboda et al. 2024b found a steeper $\Gamma_{\rm th}\sim 4.9$ and low $cov_{\rm frac}\sim 20\%$ in the HSS of Swift J1727.816131727.816131727.8-16131727.8 - 1613, which agrees well with our findings of $\Gamma_{\rm th}\sim 4.6-5.1$ and $cov_{\rm frac}\sim 17\%$. In addition, the LHS results for Cyg X--1 (E1) are consistent with those of Krawczynski et al. (2022), except for their measurement of a high electron temperature, which is likely because of the differences in spectral energy coverage and/or the choice of model components.
We also notice that the Comptonization model with a disc-like coronal geometry (comptt) yields a comparable fit to the observed spectra similar to the spherical corona model (thcomp). This evidently indicates the degeneracy in determining the corona geometry (see §4.2). Furthermore, strong reflection signatures detected in the soft state observations (HSS and SIMS) of Cyg X11-1- 1, Swift J1727.816131727.816131727.8-16131727.8 - 1613, and GX $339-4$ allow the detailed investigation of alternative coronal geometries using various reflection models (see §4.2). We also observe that the reflected emission, either from a lamppost corona, modeled with relxilllp, or from a spherical corona as described in relxillCp, can be reproduced satisfactorily considering a highly ionized accretion disc ($\log\xi\sim 4$). All these findings highlight the inherent degeneracy among different disc-corona geometry models in explaining the observed spectral characteristics.
5.2 Model Prescriptions and Limitations of X-ray Polarization in BH-XRBs
X-ray polarization measurements with IXPE reveal significant polarized emission across several spectral states of the sources in $2-8$ keV energy range, except for LMC X--1 (see §4.3 and Podgorný et al. 2023), Swift J151857.057214151857.057214151857.0-57214151857.0 - 57214 (see §4.3 and Ling et al. 2024) and MAXI J174429417442941744-2941744 - 294 (see §4.3). We find that Cyg X--1 exhibits PD within $3-4.8\%$ in LHS, which decreases to $1.4-2.5\%$ in HSS. A similar noticeable reduction in the PD from $4.7\%$ (HIMS) to $0.4\%$ (HSS) is also observed for the exceptionally bright transient Swift J1727.8--1613. The most recent IXPE observation of this source confirms the recovery of PD $\sim 3.2\%$ in the faint LHS. We mention that these results are in agreement with the previously reported polarimetric findings of Cyg X--1 (Krawczynski et al., 2022; Jana & Chang, 2024; Steiner et al., 2024) and Swift J1727.8--1613 (Veledina et al., 2023; Ingram et al., 2024; Svoboda et al., 2024b; Podgorný et al., 2024).
Intriguingly, an exceptionally high PD of $\sim 8.3\%$ is observed in the HSS of 4U 16304747-47- 47, which drops to $\sim 6.8\%$ during the SPL state exhibiting strong ($>4\sigma$) energy dependency (see Table 5 and Fig. 5; also Kushwaha et al. 2023b; Rodriguez Cavero et al. 2023; Rawat et al. 2023b, a; Ratheesh et al. 2024). In contrast, low PD values are observed in several sources: LMC X33-3- 3 ($\sim 2.4-3\%$), 4U 1957$+115$ ($\sim 1.9\%$) and GX $339-4$ ($\sim 1.2\%$), being consistent with the previous measurements (Majumder et al., 2024a; Svoboda et al., 2024a; Garg et al., 2024; Marra et al., 2024; Mastroserio et al., 2025). In the present work, we report polarization detections analyzing the most recent IXPE observations of Cyg X--3 (PD $\sim 21.4\%$, SIMS), LMC X--3 (PD $\sim 2.4\%$, HSS) and IGR J17091--3624 (PD $\sim 9\%$, LHS), apart from the earlier measurements of these sources. In addition, for the first time to the best of our knowledge, we report the null-detection of X-ray polarization in LMC X33-3- 3 (E3) and a newly discovered BH-XRB candidate MAXI J174429417442941744-2941744 - 294.
Meanwhile, several models have been proposed to account for the observed polarization in BH-XRBs. For instance, remarkably high PD ($\sim 8.3\%$) of 4U 1630--47 is attributed to the emission from a partially ionized thin/slim accretion disc viewed at inclination $i\gtrsim 60^{\circ}$ (Ratheesh et al., 2024). Similar explanation is also suggested for LMC X33-3- 3 in the HSS that exhibits PD $\sim 3\%$ at $i\sim 70^{\circ}$ (Majumder et al., 2024a; Svoboda et al., 2024a). Further, scattering of photons in strong disc winds of different opening angles ($\alpha_{\rm w}\sim 30^{\circ}-40^{\circ}$) has also been suggested as another possible mechanism to explain high PD values of 4U 1630--47 (Kushwaha et al., 2023b; Rawat et al., 2023b; Nitindala et al., 2025). On the other hand, in high-spin black hole systems, strong relativistic effects cause intense gravitational lensing that bends emitted photons back onto the disc. This enhances the reflection features and significantly affects the observed polarization (see §4.2, Schnittman & Krolik, 2009). Such an effect appears consistent with the polarimetric observations (PD $\sim 1.2-2.5\%$) in the soft state as reported for rapidly spinning systems like Cyg X11-1- 1 ($a_{*}>0.998$) and 4U $1957+115$ ($a_{*}\sim 0.988$) (Marra et al., 2024; Steiner et al., 2024).
Furthermore, it has been suggested that $\sim 4\%$ (Cyg X11-1- 1, LHS) and $\sim 9\%$ (IGR J17091362417091362417091-362417091 - 3624, LHS) polarization can be obtained from a wedge-shaped corona with mildly relativistic outflowing plasma having inclination in the range $30^{\circ}\lesssim i\lesssim 60^{\circ}$ (Poutanen et al., 2023; Ewing et al., 2025). A similar model of bulk Comptonization involving a mildly relativistic outflow predicts an even higher polarization degree up to $\lesssim 10\%$ (Dexter & Begelman, 2024). Notably, observed $\sim 4.7\%$ PD at $i\sim 40^{\circ}$ in HIMS of Swift J1727.816131727.816131727.8-16131727.8 - 1613 also seems to be consistent with the predictions from a wedge-shaped corona with outflowing plasma velocity $\lesssim 0.2c$ (Poutanen et al., 2023). However, QPOs are detected in Swift J1727.816131727.816131727.8-16131727.8 - 1613 ($0.5-100$ keV) and IGR J17091362417091362417091-362417091 - 3624 ($3-79$ keV) during observations simultaneous with IXPE, although their physical origin remains unclear within the framework of the outflowing corona geometry. On the other hand, Kumar (2024) predicted PD $\sim 0.5-4\%$ assuming a simple static spherical corona for $30^{\circ}\lesssim i\lesssim 60^{\circ}$. This appears broadly consistent with the low PD of $\sim 1.2\%$ observed at $i\sim 30^{\circ}$ in the SIMS of GX $339-4$.
All the above interpretations of the polarimetric findings point toward a persistent degeneracy in understanding the different accretion-ejection dynamics. More precisely, predictions regarding the disc-corona-jet geometry become even more challenging when the simultaneous origin of temporal features (QPOs), spectral distributions and polarization (PD) is considered. This emphasizes the need for a unified framework of accretion-ejection processes in BH-XRBs that can consistently explain both timing and spectro-polarimetric observations.
5.3 Accretion-Ejection Scenarios in BH-XRBs
Figure 7: Schematic representation of the possible disc-corona-jet configurations for seven BH-XRB sources based on the framework of two-component accretion flow scenario. The Keplerian and sub-Keplerian components, as indicated in panel (b), are also present in the other two configurations. Colored vertical arrows in the corona regions of panels (a) and (c) indicate the direction of polarization angle for the respective sources. In panels (a) and (c), the jet configuration applies only to the sources in distinct spectral states, which are marked by their respective serial numbers. See the text for details.
Following the discussion of model degeneracies and the need for a unified model framework, we next explore physically motivated accretion-ejection geometries capable of explaining the commonly observed spectro-temporal features, while also offering a promising basis for interpreting the spectro-polarimetric signatures in BH-XRBs. Towards this, a plausible disc-corona-jet configuration based on the two-component accretion flow model Chakrabarti & Titarchuk (1995) is illustrated in Fig. 7 and discussed in detail below.
The spectro-polarimetric correlation study (see §4.6 and Fig. 6) suggests that the polarization signatures (PD $\sim 3-9\%$) in the LHS/HIMS of Cyg X11-1- 1, Swift J1727.816131727.816131727.8-16131727.8 - 1613 and IGR J17091362417091362417091-362417091 - 3624 are primarily governed by the dominant Comptonization process characterized by $F_{\rm ratio}<1$ and $cov_{\rm frac}\sim 35-87\%$. Interestingly, the radio jet position angle of Cyg X11-1- 1 and Swift J1727.816131727.816131727.8-16131727.8 - 1613 roughly aligns along the observed PA throughout the respective IXPE campaigns (Miller-Jones et al., 2021; Wood et al., 2024). This suggests that PA is possibly oriented perpendicular to the disc-corona geometry. Further, it may be noted that PA resulting from multiple Compton up-scatterings generally tends to align along the minor axis of the corona (Ingram et al., 2024). These pieces of evidence suggest a disc-corona configuration comprising a radial corona located close to the black hole, along with a truncated accretion disc at larger radius during LHS/HIMS (see Fig. 7a) of Swift J1727.816131727.816131727.8-16131727.8 - 1613 and Cyg X11-1- 1. The above configuration potentially explains the QPOs in Swift J1727.816131727.816131727.8-16131727.8 - 1613 (see §4.2 and Fig. 2c) in terms of the oscillating boundary of the Comptonizing region (see Fig. 7a and Molteni et al. 1996; Nandi et al. 2024). It is worth mentioning that the coronal geometry in IGR J17091362417091362417091-362417091 - 3624 remains inconclusive from the polarimetric results due to the unresolved position angle of the radio jet. Although the presence of QPO in the LHS of IGR J17091362417091362417091-362417091 - 3624 (see Fig. 2g) perhaps indicates a similar disc-corona configuration presented in Fig. 7a. However, the absence of QPO in Cyg X11-1- 1 implies that either the coronal boundary fails to satisfy the necessary resonance condition for modulation, or the oscillation is too subtle to be detected as predicted by Majumder et al. (2025a).
For 4U $1630-47$, an absorption feature around $\sim 7$ keV is observed, suggesting the presence of strong disc winds in the HSS ($F_{\rm ratio}\sim 52$ and $cov_{\rm frac}\sim 5\%$) spectra (see §4.2). Accordingly, a disc-dominated accretion configuration along with a weak corona, accompanied by strong disc winds with varying opening angles (see also Nitindala et al. 2025), is likely favored for this source (see Fig. 7b). However, a relatively strong corona ($cov_{\rm frac}\sim 38\%$) may be present in the SPL state as compared to the HSS (see Fig. 7b).
For GX $339-4$, a low PD ($\sim 1.2\%$) is observed within $3-8$ keV energy range during SIMS ($cov_{\rm frac}\sim 34\%$ and $F_{\rm ratio}\sim 2.4$), while null-detection of PD is seen in the HSS (Table 5). This possibly associates the polarization signature of the source with Comptonized emission. We notice $\rm PD<\rm MDP_{99}\%$ in the entire IXPE energy range of $2-8$ keV during SIMS, which possibly resulted from the depolarization of the overall radiation by unpolarized and/or oppositely polarized (as compared to corona) disc emission (Ingram et al., 2024) at lower energies ($2-3$ keV). The polarization angle of GX $339-4$ aligns with the direction of the discrete radio jet knot (Mastroserio et al., 2025). These findings predict the GX $339-4$ configuration comprising a radially extended weak compact corona coexisting with a strong disc ($F_{\rm ratio}\sim 2.4$, $\Gamma_{\rm th}\sim 2$), as illustrated in Fig. 7c. Moreover, a type-B QPO (see Fig. 2e) observed in this source could be explained from this configuration using an analogous argument presented above for Fig. 7a (see also Nandi et al. 2012; Aneesha et al. 2024).
Furthermore, we conjecture that a similar disc-corona configuration (see Fig. 7c) is also preferred in the disc-dominated HSS/SIMS of other sources (LMC X33-3- 3, 4U $1957+115$, Cyg X11-1- 1 and Swift J1727.816131727.816131727.8-16131727.8 - 1613) with major contributions from thermal disc and/or various general relativistic effects (Majumder et al., 2024a; Svoboda et al., 2024a; Marra et al., 2024; Steiner et al., 2024; Svoboda et al., 2024b) in the observed PD ($1.4-3\%$). It is noteworthy that although the presence of a radial corona is favored for 4U $1630-47$, LMC X33-3- 3, and 4U $1957+115$ in Fig. 7, the coronal geometry remains uncertain due to the lack of resolved radio jet orientation for these sources. Moreover, high PD (up to $\sim 21\%$) dominated by reflected emission in Cyg X33-3- 3 appears inconsistent with the disc-corona configurations presented in Fig. 7. It has been suggested that a disc-corona geometry featuring an optically thick funnel in the innermost region of the accretion disc, which obscures most of the primary emission, is favored for Cyg X33-3- 3 Veledina et al. (2024a, b). Finally, we mention that, although the two-component accretion flow configuration aligns with the disc-corona geometry inferred from polarimetric studies for most of the BH-XRBs, a detailed radiative transfer computation is needed within this framework for the predictions on X-ray polarization.
5.4 Frontiers of X-ray Polarimetry: A Promising Era Ahead
Figure 8: (a) Time binned ($50$ s) light curve of Swift J1727.816131727.816131727.8-16131727.8 - 1613 observed with IXPE in the hard-intermediate state (HIMS). (b) Best-fitted power density spectra ($0.1-50$ Hz) obtained from the quasi-simultaneous NICER ($0.5-10$ keV), NuSTAR ($3-20$ keV) and AstroSat ($20-100$ keV) observations. (c) Best-fitted energy spectra from the quasi-simultaneous NICER and NuSTAR observations in $0.7-40$ keV energy range. (d) IXPE image ($2-8$ keV) of Swift J1727.816131727.816131727.8-16131727.8 - 1613. The red circular boundary denotes $60$ arcsec source region considered for polarimetric analysis. (e) Variation of the polarization degree (PD) with energy obtained from the model-independent polarimetric analysis. The gray histograms represent the $\rm MDP_{\rm 99}\%$ level. See the text for details.
Indeed, the first X-ray polarization measurements were conducted in the 1970s by OSO-8, revealing significant polarization in Cyg X11-1- 1 with PD $\sim 2.4\%$ at $2.6$ keV and $\sim 5.3\%$ at $5.2$ keV (Long et al., 1980). Thereafter, higher polarization was observed in hard X-rays for Cyg X11-1- 1 with INTEGRAL as $\lesssim 20\%$ in the $130-230$ keV band (Jourdain et al., 2012) and $\sim 75\%$ above $400$ keV (Rodriguez et al., 2015). Recently, Chattopadhyay et al. (2024) detected $\sim 23\%$ polarization in the HIMS of Cyg X11-1- 1 using AstroSat/CZTI. Despite these advances, low-energy X-ray polarimetry remained relatively underexplored until the launch of IXPE, which has made substantial progress in this domain during its first three and a half years of operation. Nevertheless, several fundamental questions remain unanswered, paving the way for future research.
At present, Swift J1727.816131727.816131727.8-16131727.8 - 1613 is the only BH-XRB transient observed during its outburst with IXPE exhibiting polarization over four spectral states (HIMS, SIMS, HSS and LHS) along with QPO features. To establish robust spectro-polarimetric correlations across different spectral states, coordinated IXPE observations along with simultaneous spectro-temporal coverage from other missions are indispensable. Further, since the effects of Comptonization become prominent beyond the coverage of IXPE up to $8$ keV, complementary wide-band X-ray spectro-polarimetry and timing studies are crucial to constrain the coronal geometry. Thus, the four windows (see Fig. 8) of X-ray astronomy (imaging, timing, spectroscopy and polarimetry) with wide energy coverage is crucial to unravel the complete understanding of BH-XRBs (see Fig. 8).
In this context, the recently launched XPoSat [^16] https://www.isro.gov.in/XPoSat.html mission and the upcoming XL-Calibur mission (Abarr et al., 2021) are set to play a crucial role in investigating hard X-ray polarimetric properties of BH-XRBs. In particular, POLIX ($8-30$ keV) and XSPECT ($0.8-15$ keV) onboard on XPoSat are suitable for mid-energy polarimetry and simultaneous spectro-temporal studies, respectively. The balloon-borne XL-Calibur mission, operating in $15-80$ keV energy range, completed a successful flight [^17] https://sscspace.com/nasa-xl-calibur-balloon-launched/ in July 2024, aiming to measure polarization in Cyg X11-1- 1. Additionally, the future soft X-ray spectro-polarimetric mission eXTP [^18] https://heasarc.gsfc.nasa.gov/docs/heasarc/missions/extp.html ($2-10$ keV), expected to launch in $2027$, will offer improved sensitivity for low-energy polarization measurements. Needless to mention that these dedicated missions will significantly advance our understanding of the accretion dynamics and coronal geometries in BH-XRBs. However, simultaneous observations combining wide-band timing and spectro-polarimetric measurement are still lacking, which emphasizes the need for future missions with broader energy ranges and enhanced capabilities.
6 Conclusions
In this paper, we perform a detailed timing and spectro-polarimetric study of eleven BH-XRBs using quasi-simultaneous IXPE, NICER, NuSTAR and AstroSat observations. Our analyses provide valuable insights into the accretion-ejection dynamics and the geometry of BH-XRBs under consideration. The key findings from our study along with their implications are summarized below.
  • The combined spectro-temporal results in the wide-band energy range ($0.5-100$ keV) reveal the presence of distinct canonical spectral states of BH-XRBs, which are closely connected to the temporal characteristics and emission mechanism of the sources. The timing and spectral features result in degeneracy among the different viable disc-corona geometries of the sources.
  • The detection of X-ray polarization is reconfirmed in eight out of eleven BH-XRBs with moderate to strong energy dependence in PD. A comprehensive spectro-polarimetric correlation study reveals significant positive (negative) correlations between PD and $cov_{\rm frac}$ ($F_{\rm ratio}$) indicating contributions from various surrounding components (disc and corona) to the observed polarization.
  • Timing and spectro-polarimetric results, combined with known radio-jet angles, suggest that a two-component disc-corona model may be plausible. In this framework, a radially extended strong corona is likely to be present at the inner part of a truncated accretion disc during harder states, while a comparatively weaker corona persists in softer states for Swift J1727.8--1613, Cyg X--1, GX 339--4, and IGR J17091--3624.
  • For 4U 1630--47, LMC X--3, and 4U 1957$+$115, a thermally dominated accretion disc with a weak corona seems to be preferred. However, the coronal geometry remains unclear due to the minimal contribution of Comptonized emission in the IXPE band and the lack of complementary radio observations, though this analysis does support the presence of disc winds in 4U 1630--47. We also note that the disc-corona geometry outlined in this work bears limitations to explain the observed features of Cyg X--3, for which an alternative configuration involving an optically thick inner funnel has been proposed (Veledina et al., 2024a, b).
7 Acknowledgments
SM and SD thank the Department of Physics, IIT Guwahati, India for providing the facilities to complete this work. SM also thanks Luca Baldini for the discussion in identifying the bugs in IXPEOBSSIM software and releasing the latest version IXPEOBSSIM V31.0.3. AK, SS, KJ and AN thank GH, SAG; DD, PDMSA, and Director, URSC for encouragement and continuous support to carry out this research. This publication uses data from the IXPE, NICER, NuSTAR and AstroSat missions. This research has made use of the MAXI data provided by RIKEN, JAXA and the MAXI team (Matsuoka et al., 2009). The Swift/BAT transient monitor results provided by the Swift/BAT team are also used (Krimm et al., 2013). We thank each instrument team for processing the data and providing the necessary software tools for the analysis.
8 Data Availability
Data used for this publication from IXPE, NICER and NuSTAR missions are currently available at the HEASARC browse (https://heasarc.gsfc.nasa.gov/db-perl/W3Browse/w3browse.pl). The AstroSat data is available at the ISSDC data archive (https://webapps.issdc.gov.in/astro_archive/archive/Home.jsp). The MAXI/GSC on-demand data is available at http://maxi.riken.jp/mxondem/ and Swift/BAT data is taken from https://swift.gsfc.nasa.gov/results/transients/.
Table 1: Details of the multi-mission observations of the selected sources over different epochs. In the table, blue shades represent the epochs for which quasi-simultaneous IXPE, NICER, and NuSTAR observations are available. The gray shades denote the quasi-simultaneous observations with IXPE and NICER only. The magenta colored shade indicates the epoch for which only quasi-simultaneous IXPE and NuSTAR observations are available. The orange colored shade denotes the available AstroSat observations close to the IXPE epochs. See the text for details.
Source Epoch Instrument Obs. ID Start Date MJD Exposure Mean Rate MAXI/GSC Swift/BAT Radio Flux Spectral
Start (ks) (cts/s) Flux (mCrab) Flux (mCrab) (mJy) State
($2-20$ keV) ($15-50$ keV)
Cyg X11-1- 1 E1 IXPE 01002901 2022-05-15 59714.64 242 $9.82\pm 0.21$ $401\pm 11$ $928\pm 34$- LHS
NICER 5100320101 2022-05-15 59714.26 8.2 $2938\pm 25$- -- -- --
NuSTAR 30702017002 2022-05-18 59717.57 16 $324\pm 15$- -- -- --
AstroSat A11_080T01_9000005146 2022-05-15 59714.03 $47.3$ $1505\pm 37$- -- -- --
E2 IXPE 01250101 2022-06-18 59748.86 86 $10.32\pm 0.12$ $420\pm 14$ $885\pm 33$- LHS
NICER 5100320108 2022-06-20 59750.60 3 $3001\pm 25$- -- -- --
NuSTAR 90802013002 2022-06-20 59750.50 13 $338\pm 14$- -- -- --
E3 IXPE 02008201 2023-05-02 60066.96 21 $26.93\pm 0.18$ $888\pm 12$ $546\pm 21$ $9.1\pm 0.3^{\boxdot}$ HSS
E4 IXPE 02008301 2023-05-09 60073.44 31 $33.93\pm 0.23$ $701\pm 12$ $613\pm 23$ $11.7\pm 0.2^{\boxdot}$ HSS
E5 IXPE 02008401 2023-05-24 60088.82 25 $44.26\pm 0.30$ $809\pm 20$ $569\pm 35$ $2.4\pm 0.2^{\boxdot}$ HSS
AstroSat T05_105T01_9000005662 2023-05-24 60088.79 60.5 $1354\pm 36$- -- -- --
E6 IXPE 02008501 2023-06-13 60108.96 29 $39.33\pm 0.28$ $765\pm 27$ $593\pm 45$ $8.9\pm 0.2^{\boxdot}$ HSS
NuSTAR 80902318004 2023-06-14 60109.02 9.8 $694\pm 22$- -- -- HSS
E7 IXPE 02008601 2023-06-20 60115.04 34 $48.09\pm 0.24$ $803\pm 24$ $859\pm 46$ $9.4\pm 0.4^{\boxdot}$ HSS
NICER 6643010104 2023-06-20 60115.97 7.6 $23227\pm 201$- -- -- --
NuSTAR 80902318006 2023-06-20 60115.82 10.2 $878\pm 26$- -- -- --
E8 IXPE 03002201 2024-04-12 60412.02 55.8 $7.28\pm 0.12$ $333\pm 9$ $736\pm 31$- LHS
NICER 7100320104 2024-04-11 60411.10 0.7 $1508\pm 18$- -- -- --
E9 IXPE 03003101 2024-05-06 60436.37 53.9 $7.14\pm 0.11$ $283\pm 8$ $919\pm 42$- LHS
E10 IXPE 03010001 2024-05-26 60456.04 57.5 $7.29\pm 0.12$ $286\pm 18$ $747\pm 28$- LHS
E11 IXPE 03010101 2024-06-14 60475.65 55.8 $5.72\pm 0.09$ $293\pm 11$ $725\pm 27$- LHS
NICER 7706010104 2024-06-14 60475.99 6.1 $1698\pm 54$- -- -- --
E12 IXPE 03002599 2024-10-10 60593.22 110 $22.1\pm 2.2$ $600\pm 13$ $795\pm 27$- HSS
4U 16304747-47- 47 E1 IXPE 01250401 2022-08-23 59814.95 459 $4.56\pm 0.08$ $247\pm 8$ $10\pm 2$- HSS
NICER 5501010102 2022-08-23 59814.01 2 $479\pm 10$- -- -- --
NuSTAR 80802313002 2022-08-25 59816.19 17 $348\pm 14$- -- -- --
E2 IXPE 02250601 2023-03-10 60013.78 138 $10.16\pm 0.12$ $550\pm 16$ $246\pm 20$- SPL
NICER 6557010101 2023-03-10 60013.76 5 $1009\pm 15$- -- -- --
NuSTAR 80801327002 2023-03-09 60012.36 12 $884\pm 29$- -- -- --
AstroSat A12_056T02_9000005538 2023-03-10 60013.08 29.4 $1938\pm 43$- -- -- --
Cyg X33-3- 3 E1 IXPE 02001899 2022-10-14 59866.04 538 $1.33\pm 0.04$ $106\pm 20$ $129\pm 26$ $106\pm 24^{*}$ LHS
NICER 5142010105 2022-10-14 59866.28 5 $44.79\pm 3.12$- -- -- --
NuSTAR 90802323002 2022-10-13 59865.62 18 $197\pm 1$- -- -- --
E2 IXPE 02250301 2022-12-25 59938.41 199 $4.55\pm 0.09$ $295\pm 31$ $144\pm 35$ $107\pm 36^{*}$ HSS
NuSTAR 90801336002 2022-12-25 59938.33 36 $467\pm 2$- -- -- --
E3 IXPE 02009101 2023-11-17 60265.83 291 $1.43\pm 0.05$ $131\pm 8$ $159\pm 13$- SIMS
NICER 6692010101 2023-11-17 60265.77 1.9 $125\pm 7$- -- -- --
E4 IXPE 03250301 2024-06-02 60463.77 50 $7.45\pm 0.12$ $411\pm 16$ $14\pm 9$- HSS
LMC X11-1- 1 E1 IXPE 02001901 2022-10-19 59871.61 563 $0.96\pm 0.03$ $14\pm 9$ $13\pm 3$- HSS
NICER 5100070101 2022-10-19 59871.79 2.8 $193\pm 6$- -- -- --
NuSTAR 90801324002 2022-10-24 59876.16 19 $15\pm 2$- -- -- --
4U 1957$+115$ E1 IXPE 02006601 2023-05-12 60076.10 572 $1.09\pm 0.04$ $19\pm 4$ $2\pm 1$- HSS
NICER 6100400101 2023-05-12 60076.55 0.6 $265\pm 8$- -- -- --
NuSTAR 30902042002 2023-05-15 60079.33 19 $29\pm 4$- -- -- --
LMC X33-3- 3 E1 IXPE 02006599 2023-07-07 60132.77 562 $0.97\pm 0.03$ $16\pm 4$ $18\pm 3$- HSS
NICER 6101010117 2023-07-08 60133.42 0.1 $358\pm 23$- -- -- --
NuSTAR 30902041002 2023-07-09 60134.51 28 $15\pm 3$- -- -- --
E2 IXPE 03004899 2024-10-03 60586.51 385 $1.62\pm 0.04$ $40\pm 5$ $14\pm 9$- HSS
NICER 7704010101 2024-10-03 60586.63 2.4 $606\pm 11$- -- -- --
E3 IXPE 03004901 2024-11-21 60635.94 382 $0.42\pm 0.02$ $8\pm 5$- -- HSS
NICER 7704010201 2024-11-23 60637.10 1.7 $246\pm 7$- -- -- --
  • Data are not available in public domain. AMI-LA ($15.5$ GHz, Steiner et al. 2024). RATAN ($4.7$ GHz, Veledina et al. 2024a).
Table 2: Same as Table 1.
Source Epoch Instrument Obs. ID Start Date MJD Exposure Mean Rate MAXI/GSC Swift/BAT Radio Flux Spectral
Start (ks) (cts/s) Flux (mCrab) Flux (mCrab) (mJy) State
($2-20$ keV) ($15-50$ keV)
Swift J1727.816131613-1613- 1613 E1 IXPE 02250901 2023-09-07 60194.81 19 $42.21\pm 0.28$ $6053\pm 71$ $4898\pm 173$ $120\pm 12^{**}$ HIMS
NICER 6203980115 2023-09-08 60195.61 0.2 $64955\pm 274$- -- -- --
NuSTAR 80902333006 2023-09-07 60194.78 0.7 $6536\pm 8$- -- -- --
AstroSat T05_145T01_9000005836 2023-09-08 60195.07 $32$ $541\pm 23$- -- -- --
E2 IXPE 02251001 2023-09-16 60203.70 37 $41.29\pm 0.25$ $5058\pm 31$ $2472\pm 98$ $144\pm 4.32^{*}$ HIMS
NICER 6203980119 2023-09-13 60200.18 4.9 $65392\pm 122$- -- -- --
NuSTAR 80902313002 2023-09-16 60203.81 0.6 $5478\pm 26$- -- -- --
E3 IXPE 02251101 2023-09-27 60214.92 21 $36.85\pm 0.22$ $3395\pm 49$ $300\pm 10$ $88\pm 2.64^{*}$ HIMS
E4 IXPE 02251201 2023-10-04 60221.54 17 $41.31\pm 0.24$ $4689\pm 46$ $206\pm 38$ $11\pm 0.33^{*}$ SIMS
NICER 6557020401 2023-10-04 60221.18 6.5 $66263\pm 62$- -- -- --
NuSTAR 80902313008 2023-10-04 60221.54 0.5 $4111\pm 4$- -- -- --
E5 IXPE 02251301 2023-10-10 60227.47 18 $34.62\pm 0.20$ $3446\pm 38$ $206\pm 38$ $35\pm 1.05^{*}$ SIMS
NICER 6203980136 2023-10-09 60226.02 1.9 $57789\pm 149$- -- -- --
NuSTAR 80902313016 2023-10-10 60227.82 1.1 $3196\pm 4$- -- -- --
E6 IXPE 03005701 2024-02-11 60351.39 67 $8.09\pm 0.13$ $80\pm 7$ $21\pm 12$- HSS
NICER 7708010101 2024-02-11 60351.39 3.1 $6415\pm 37$- -- -- --
E7 IXPE 03006001 2024-02-20 60360.07 151 $5.59\pm 0.09$ $67\pm 8$ $26\pm 22$- HSS
NICER 7708010106 2024-02-19 60359.07 0.6 $5354\pm 34$- -- -- --
E8 IXPE 03005801 2024-04-03 60403.66 202 $7.87\pm 0.11$ $76\pm 11$ $109\pm 14$- LHS
NICER 7708010109 2024-04-03 60403.29 2.5 $675\pm 12$- -- -- --
GX $339-4$ E1 IXPE 03005101 2024-02-14 60354.93 95 $17.44\pm 0.21$- $79\pm 25$- SIMS
NICER 7702010112 2024-02-14 60354.98 3.8 $5263\pm 34$- -- -- --
NuSTAR 91002306002 2024-02-14 60354.70 16 $308\pm 13$- -- -- --
AstroSat A05_166T01_9000006070 2024-02-14 60354.31 $50.7$ $629\pm 26$- -- -- --
E2 IXPE 03005301 2024-03-08 60377.68 98 $7.37\pm 0.11$ $103\pm 6$ $12\pm 8$- HSS
AstroSat A13_028T01_9000006122 2024-03-10 60379.44 -- -- -- -- -- --
Swift J151857.0572147151857.0572147151857.0-572147151857.0 - 572147 E1 IXPE 03250201 2024-03-18 60387.15 96 $12.18\pm 0.12$ $422\pm 8$ $152\pm 17$- HSS
NICER 7204220111 2024-03-18 60387.55 4.2 $1468\pm 18$- -- -- --
NuSTAR 91001311004 2024-03-18 60387.64 9.2 $444\pm 17$- -- -- --
IGR J17091362417091362417091-362417091 - 3624 E1 IXPE 04250201 2025-03-07 60741.30 163 $0.27\pm 0.02$ $755\pm 11$ $82\pm 27$- LHS
NuSTAR 81002342008 2025-03-07 60741.58 21 $27\pm 15$- -- -- --
MAXI J174429417442941744-2941744 - 294 E1 IXPE 04250301 2025-04-05 60770.09 149 $0.17\pm 0.11$ $45\pm 28$- -- HSS
  • Data are not available in public domain. RATAN ($4.7$ GHz, Veledina et al. 2023; Ingram et al. 2024); ∗∗E-MERLIN ($5.1$ GHz, Williams-Baldwin et al. 2023).
[…]
Source Epoch $n_{\rm H}$ $kT_{\rm in}$ $kT_{\rm e}$ $\Gamma_{\rm th}$ $cov_{\rm frac}$ $\chi^{2}/d.o.f$ $F_{\rm disc}^{\boxtimes}$ $F_{\rm Comp}^{\boxtimes}$ $F_{\rm bol}^{\boxtimes}$ $L_{\rm bol}^{\boxtimes}$ Spectral Energy $rms_{\rm tot}^{\dagger}$
[…]
Cyg X11-1- 1 E1 $0.44_{-0.02}^{+0.02}$ $0.41_{-0.03}^{+0.03}$ $30.35_{-2.22}^{+2.87}$ $1.61_{-0.01}^{+0.02}$ $0.77_{-0.02}^{+0.02}$ $1517/1468$ $0.68$ $2.45$ $3.07$ $1.4$ LHS $0.7-60$ $32.5$
($d=2.2$ kpc) E2 $0.34_{-0.02}^{-0.02}$ $0.38_{-0.01}^{+0.01}$ $22.57_{-1.13}^{+1.50}$ $1.61_{-0.01}^{+0.01}$ $0.73_{-0.03}^{+0.04}$ $1479/1403$ $0.74$ $2.53$ $3.29$ $1.5$ LHS $0.7-60$ $31$
E6 $0.51^{*}$ $0.52_{-0.02}^{+0.02}$ $10^{*}$ $1.01^{*}$ $0.005_{-0.002}^{+0.002}$ $2122/1913$ $1.75$ $0.22$ $4.43$ $2$ HSS $3-60$-
E7 $0.52_{-0.01}^{+0.01}$ $0.43_{-0.01}^{+0.02}$ $10^{*}$ $3.21_{-0.14}^{+0.11}$ $0.31_{-0.02}^{+0.03}$ $2185/2167$ $3.78$ $0.33$ $6.18$ $2.8$ HSS $0.7-60$ $9.1$
E8 $0.43_{-0.03}^{+0.03}$ $0.27_{-0.03}^{+0.03}$ $20^{*}$ $1.64_{-0.02}^{+0.02}$ $0.65_{-0.07}^{+0.06}$ $115/133$ $0.37$ $1.15$ $1.53$ $0.7$ LHS $0.7-10$ $41.5$
E11 $0.39_{-0.04}^{+0.02}$ $0.32_{-0.04}^{+0.03}$ $20^{*}$ $1.62_{-0.03}^{+0.03}$ $0.61_{-0.04}^{+0.04}$ $152/131$ $0.39$ $1.18$ $1.58$ $0.7$ LHS $0.7-10$ $52.4$
4U 16304747-47- 47 E1 $6.69_{-0.30}^{+0.24}$ $1.36_{-0.01}^{+0.01}$ $20^{*}$ $3.34_{-0.11}^{+0.12}$ $0.052_{-0.007}^{+0.008}$ $1134/928$ $1.56$ $0.03$ $1.59$ $15.1$ HSS $0.7-40$ $6.7$
($d=10$ kpc) E2 $4.55_{-0.05}^{+0.07}$ $1.46_{-0.02}^{+0.02}$ $20^{*}$ $2.47_{-0.04}^{+0.05}$ $0.38_{-0.01}^{+0.02}$ $2366/1980$ $2.76$ $0.77$ $3.65$ $34.7$ SPL $0.7-60$ $4.8$
Cyg X33-3- 3 E1 $4.04_{-0.11}^{+0.26}$ $1.21_{-0.02}^{+0.02}$ $6.14_{-0.34}^{+0.28}$ $1.55_{-0.04}^{+0.03}$ $0.83_{-0.01}^{+0.02}$ $2921/1842$ $0.21$ $0.41$ $0.64$ $5.7$ LHS $3-50$-
($d=9.7$ kpc) E2 $5.54_{-0.13}^{+0.07}$ $1.01_{-0.02}^{+0.01}$ $7.95_{-0.14}^{+0.22}$ $2.21_{-0.01}^{+0.01}$ $0.63_{-0.01}^{+0.01}$ $2709/1903$ $0.94$ $0.62$ $1.58$ $14.1$ HSS $3-50$-
LMC X11-1- 1 E1 $0.50_{-0.01}^{+0.01}$ $0.98_{-0.01}^{+0.01}$ $10^{*}$ $2.58_{-0.09}^{+0.10}$ $0.091_{-0.010}^{+0.012}$ $744/613$ $0.065$ $0.004$ $0.074$ $16$ HSS $0.8-30$ $10.1$
($d=48.1$ kpc)
4U 1957$+115$ E1 $0.08_{-0.01}^{+0.01}$ $1.40_{-0.01}^{+0.01}$ $10^{*}$ $1.86_{-0.10}^{+0.12}$ $0.027_{-0.004}^{+0.004}$ $728/701$ $0.092$ $0.005$ $0.096$ $0.9$ HSS $0.7-40$ $5$
($d=10$ kpc)
LMC X33-3- 3 E1 $0.04^{*}$ $1.10_{-0.01}^{+0.01}$ $10^{*}$ $1.78_{-0.35}^{+0.45}$ $0.005_{-0.002}^{+0.003}$ $567/543$ $0.061$ $0.001$ $0.062$ $13.4$ HSS $0.7-20$-
($d=48.1$ kpc) E2 $0.04^{*}$ $1.09_{-0.01}^{+0.01}$ $10^{*}$ $2.51_{-0.07}^{+0.08}$ $0.3^{*}$ $153/128$ $0.11$ $0.03$ $0.14$ $30.8$ HSS $1-10$-
E3 $0.04^{*}$ $0.6_{-0.03}^{+0.04}$ $1.04_{-0.05}^{+0.08}$ $1.72_{-0.08}^{+0.11}$ $0.4^{*}$ $118/117$ $0.03$ $0.006$ $0.035$ $7.7$ HSS $1-10$-
Swift J1727.816131613-1613- 1613 E1 $0.23_{-0.01}^{+0.01}$ $0.29_{-0.02}^{+0.01}$ $8.44_{-0.11}^{+0.16}$ $1.98_{-0.02}^{+0.04}$ $0.74_{-0.01}^{+0.01}$ $1181/931$ $10.07$ $15.01$ $25.24$ $17.5$ HIMS $0.7-40$ $12$
($d=2.7$ kpc) E2 $0.20_{-0.01}^{+0.01}$ $0.43_{-0.02}^{+0.02}$ $9.34_{-0.23}^{+0.22}$ $2.11_{-0.04}^{+0.05}$ $0.82_{-0.02}^{+0.02}$ $912/847$ $11.35$ $13.58$ $25.11$ $17.4$ HIMS $0.7-40$-
E4 $0.25_{-0.01}^{+0.01}$ $0.85_{-0.02}^{+0.03}$ $10^{*}$ $2.74_{-0.04}^{+0.04}$ $0.38_{-0.03}^{+0.04}$ $1373/1447$ $15.89$ $3.63$ $25.04$ $17.3$ SIMS $0.7-40$ $6.1$
E5 $0.29_{-0.01}^{+0.02}$ $0.88_{-0.02}^{+0.02}$ $10^{*}$ $2.32_{-0.04}^{+0.06}$ $0.19_{-0.02}^{+0.02}$ $1572/1601$ $10.12$ $2.09$ $16.78$ $11.6$ SIMS $0.7-60$ $6.4$
E6 $0.17_{-0.01}^{+0.01}$ $0.48_{-0.01}^{+0.01}$ $10^{*}$ $5.14_{-0.39}^{+0.36}$ $0.17_{-0.05}^{+0.03}$ $127/132$ $0.91$ $0.02$ $0.93$ $0.6$ HSS $0.7-10$ $1.4$
E7 $0.18_{-0.01}^{+0.01}$ $0.45_{-0.01}^{+0.01}$ $10^{*}$ $4.57_{-0.60}^{+0.65}$ $0.13_{-0.05}^{+0.09}$ $102/111$ $0.66$ $0.01$ $0.67$ $0.5$ HSS $0.7-10$ $1.3$
E8 $0.12^{*}$ $0.24_{-0.01}^{+0.01}$ $20^{*}$ $1.71_{-0.02}^{+0.03}$ $0.35_{-0.01}^{+0.01}$ $126/130$ $0.08$ $0.12$ $0.19$ $0.1$ LHS $0.7-10$ $30.2$
GX $339-4$ E1 $0.51_{-0.01}^{+0.01}$ $0.60_{-0.02}^{+0.02}$ $0.97_{-0.01}^{+0.02}$ $2.02_{-0.52}^{+0.44}$ $0.34_{-0.05}^{+0.08}$ $1603/1538$ $0.84$ $0.36$ $1.71$ $11.5$ SIMS $0.7-50$ $14.2$
($d=8.4$ kpc)
Swift J151857.0572147151857.0572147151857.0-572147151857.0 - 572147 E1 $3.83_{-0.01}^{+0.02}$ $0.94_{-0.01}^{+0.01}$ $20^{*}$ $2.50_{-0.01}^{+0.01}$ $0.21_{-0.01}^{+0.01}$ $1572/1367$ $2.11$ $0.34$ $2.71$ $8.7$ HSS $0.7-50$ $12.6$
($d=5.8$ kpc)
IGR J17091362417091362417091-362417091 - 3624 E1 $1.1^{*}$ $0.88_{-0.38}^{+0.27}$ $23.54_{-2.91}^{+4.01}$ $1.62_{-0.01}^{+0.02}$ $0.87_{-0.03}^{+0.03}$ $1465/1486$ $0.02$ $0.13$ $0.15$ $1.7$ LHS $3-60$ $24.3$
($d=11$ kpc)
  • $L_{\rm Edd}=1.26\times 10^{39}$ erg $\rm s^{-1}$ for $10M_{\odot}$ BH. Calculated in $1-100$ keV energy range. Computed in $0.1-50$ Hz using NICER observations in $0.5-10$ keV energy range. Frozen values. Results are obtained using model M3.
[…]
Source Epoch Spin Inclination ($i$) $\Gamma$ $A_{\rm fe}$ $\log\xi$ $R_{\rm f}$ $F_{\rm relxill}^{\boxdot}$ Spectral
($a_{*}$) () ($A_{\odot}$) ($\log$ erg ${\rm cm}$ ${\rm s}^{-1}$) ($\tiny{10^{-8}}$ erg ${\rm cm}^{-2}$ ${\rm s}^{-1}$) State
aCyg X11-1- 1 E6 $0.98^{*}$ $27^{*}$ $2.32_{-0.06}^{+0.04}$ $6.35_{-1.46}^{+2.08}$ $3.86_{-0.08}^{+0.09}$ $0.67_{-0.04}^{+0.04}$ $2.33$ HSS
E7 $0.98^{*}$ $27^{*}$ $2.04_{-0.01}^{+0.01}$ $4.61_{-0.31}^{+0.36}$ $3.31_{-0.05}^{+0.05}$ $1.11_{-0.03}^{+0.03}$ $2.14$ HSS
bGX $339-4$ E1 $0.85^{*}$ $50^{*}$ $2.27_{-0.03}^{+0.03}$ $0.5^{\boxtimes}$ $3.92_{-0.05}^{+0.12}$ $3.16_{-0.33}^{+0.35}$ $0.54$ SIMS
cSwift J1727.816131727.816131727.8-16131727.8 - 1613 E4 $0.98^{*}$ $40^{*}$ $2.19_{-0.02}^{+0.03}$ $5.02^{\boxtimes}$ $4.11_{-0.13}^{+0.12}$ $0.68_{-0.14}^{+0.17}$ $5.44$ SIMS
E5 $0.98^{*}$ $40^{*}$ $2.48_{-0.03}^{+0.02}$ $5.1^{*}$ $4.29_{-0.21}^{+0.14}$ $0.59_{-0.11}^{+0.05}$ $4.38$ SIMS
  • Frozen values. Fixed to the best-fitted values. Calculated in $1-100$ keV energy range. $A_{\odot}$ is the solar iron abundance. $r_{\rm g}=GM/\rm c^{2}$. aKushwaha et al. (2021), bMastroserio et al. (2025), cPeng et al. (2024a).
Table 5: Model-independent polarimetric results obtained using the PCUBE algorithm in the $2-8$ keV energy band from IXPE observations of eleven BH-XRBs. PD and PA represent the polarization degree and angle, respectively. Q///I and U///I are the normalized Stokes parameters. MDP denotes the minimum detectable polarization, and SIGNIF is the detection significance in units of $\sigma$. PDE represents the significance of the energy variation of PD in units of $\sigma$. Gray-shaded entries indicate null-detection of polarization. See text for details.
Source Epoch Obs. ID PD PA Q///I U///I $\rm MDP_{99}$ SIGNIF Spectral Remarks PDE
(%) () (%) (%) (%) ($\sigma$) State Sig. ($\sigma$)
Cyg X11-1- 1 E1 01002901 $3.70\pm 0.19$ $-20.75\pm 1.50$ $2.77\pm 0.19$ $-2.45\pm 0.19$ $0.59$ $19$ LHS Detection $2.8$
E2 01250101 $3.79\pm 0.32$ $-25.54\pm 2.39$ $2.38\pm 0.32$ $-2.95\pm 0.32$ $0.96$ $12$ LHS Detection $<1$
E3 02008201 $2.50\pm 0.40$ $-19.23\pm 4.63$ $1.95\pm 0.40$ $-1.55\pm 0.40$ $1.22$ $5.7$ HSS Detection $<1$
E4 02008301 $2.46\pm 0.30$ $-22.89\pm 3.45$ $1.71\pm 0.30$ $-1.76\pm 0.30$ $0.90$ $8$ HSS Detection $3$
E5 02008401 $2.01\pm 0.29$ $-25.82\pm 4.11$ $1.25\pm 0.29$ $-1.58\pm 0.29$ $0.88$ $6.6$ HSS Detection $>4$
E6 02008501 $1.44\pm 0.28$ $-25.38\pm 5.66$ $0.91\pm 0.28$ $-1.12\pm 0.28$ $0.86$ $4.5$ HSS Detection $1.2$
E7 02008601 $2.07\pm 0.23$ $-36.37\pm 3.19$ $0.61\pm 0.23$ $-1.98\pm 0.23$ $0.70$ $9$ HSS Detection $1.1$
E8 03002201 $3.75\pm 0.47$ $-24.22\pm 3.61$ $2.49\pm 0.47$ $-2.80\pm 0.47$ $1.43$ $7.6$ LHS Detection $1.8$
E9 03003101 $3.04\pm 0.48$ $-18.35\pm 4.54$ $2.44\pm 0.48$ $-1.82\pm 0.48$ $1.46$ $5.9$ LHS Detection $<1$
E10 03010001 $4.65\pm 0.46$ $-27.77\pm 2.85$ $2.63\pm 0.46$ $-3.83\pm 0.46$ $1.4$ $10.1$ LHS Detection $<1$
E11 03010101 $4.76\pm 0.53$ $-32.64\pm 3.18$ $1.99\pm 0.53$ $-4.32\pm 0.53$ $1.60$ $9$ LHS Detection $1.5$
E12 03002599 $2.79\pm 0.19$ $-21.97\pm 1.97$ $2.01\pm 0.19$ $-1.94\pm 0.19$ $0.58$ $14.5$ HSS Detection $>4$
4U 16304747-47- 47 E1 01250401 $8.34\pm 0.17$ $17.80\pm 0.60$ $6.79\pm 0.17$ $4.86\pm 0.17$ $0.53$ $48$ HSS Detection $>4$
E2 02250601 $6.77\pm 0.21$ $21.36\pm 0.90$ $4.97\pm 0.21$ $4.59\pm 0.21$ $0.65$ $31.7$ SPL Detection $>4$
Cyg X33-3- 3 E1 02001899 $20.60\pm 0.31$ $-89.79\pm 0.43$ $-20.60\pm 0.31$ $-0.15\pm 0.31$ $0.94$ $66.3$ LHS Detection $>4$
E2 02250301 $10.58\pm 0.28$ $-87.39\pm 0.75$ $-10.54\pm 0.28$ $-0.96\pm 0.28$ $0.85$ $38$ HSS Detection $>4$
E3 02009101 $21.41\pm 0.41$ $-88.09\pm 0.55$ $-21.36\pm 0.41$ $-1.43\pm 0.41$ $1.24$ $52.7$ SIMS Detection $>4$
E4 03250301 $11.98\pm 0.43$ $-85.22\pm 1.03$ $-11.82\pm 0.43$ $-1.99\pm 0.43$ $1.31$ $27.8$ HSS Detection $>4$
LMC X11-1- 1 E1 02001901 $1.04\pm 0.40$ $53.97\pm 11.09$ $-0.32\pm 0.40$ $0.99\pm 0.40$ $1.23$ $1.8$ HSS Null-detection --
4U 1957$+115$ E1 02006601 $1.85\pm 0.37$ $-42.09\pm 5.75$ $0.19\pm 0.37$ $-1.84\pm 0.37$ $1.13$ $4.5$ HSS Detection $1.7$
LMC X33-3- 3 E1 02006599 $3.04\pm 0.40$ $-44.24\pm 3.77$ $0.08\pm 0.40$ $-3.04\pm 0.40$ $1.21$ $7.2$ HSS Detection $1.8$
E2 03004899 $2.39\pm 0.37$ $-38.43\pm 4.45$ $0.54\pm 0.37$ $-2.33\pm 0.37$ $1.13$ $6$ HSS Detection $2.3$
E3 03004901 $2.19\pm 0.77$ $-39.42\pm 10.01$ $0.42\pm 0.77$ $-2.15\pm 0.77$ $2.32$ $2.4$ HSS Null-detection --
Swift J1727.816131613-1613- 1613 E1 02250901 $4.72\pm 0.28$ $2.50\pm 1.69$ $4.71\pm 0.28$ $0.41\pm 0.28$ $0.84$ $17$ HIMS Detection $<1$
E2 02251001 $4.48\pm 0.20$ $2.28\pm 1.31$ $4.46\pm 0.20$ $0.36\pm 0.20$ $0.62$ $22$ HIMS Detection $2.8$
E3 02251101 $4.39\pm 0.29$ $0.95\pm 1.87$ $4.38\pm 0.29$ $0.15\pm 0.29$ $0.87$ $15.3$ HIMS Detection $2.6$
E4 02251201 $3.81\pm 0.29$ $-0.49\pm 2.22$ $3.81\pm 0.29$ $-0.07\pm 0.29$ $0.89$ $12.8$ SIMS Detection $1.1$
E5 02251301 $3.35\pm 0.32$ $-2.01\pm 2.75$ $3.34\pm 0.32$ $-0.23\pm 0.32$ $0.98$ $10.4$ SIMS Detection $2.5$
E6 03005701 $1.18\pm 0.44$ $-2.74\pm 10.55$ $1.18\pm 0.44$ $-0.11\pm 0.44$ $1.32$ $2$ HSS Null-detection --
E7 03006001 $0.40\pm 0.36$ $6.05\pm 25.83$ $0.39\pm 0.36$ $0.08\pm 0.36$ $1.08$ $<1$ HSS Null-detection --
E8 03005801 $3.22\pm 0.60$ $4.18\pm 5.37$ $3.19\pm 0.60$ $0.47\pm 0.60$ $1.83$ $4.8$ LHS Detection $1.6$
[…]
E2 03005301 $0.47\pm 0.36$ $-25.15\pm 22.07$ $0.30\pm 0.36$ $-0.36\pm 0.36$ $1.09$ $0.8$ HSS Null-detection --
Swift J151857.0572147151857.0572147151857.0-572147151857.0 - 572147 E1 03250201 $0.25\pm 0.26$ $-22.35\pm 29.98$ $0.17\pm 0.26$ $-0.17\pm 0.26$ $0.78$ $0.5$ HSS Null-detection --
IGR J17091362417091362417091-362417091 - 3624 E1111 04250201 $9\pm 1.41$ $83.71\pm 4.48$ $-8.78\pm 1.41$ $1.96\pm 1.41$ $4.27$ $6.1$ LHS Detection $1.2$
MAXI J174429417442941744-2941744 - 294 E1111 04250301 $0.71\pm 0.42$ $-15.37\pm 17.21$ $-0.60\pm 0.42$ $-0.36\pm 0.42$ $1.27$ $1.2$ HSS Null-Detection --
  • Measurements in $3-8$ keV. Null-detection in $2-8$ keV. Computed considering $6$ linear energy bins for all sources except GX $339-4$, 4U 1957$+115$ and IGR J17091362417091362417091-362417091 - 3624, where $4$ linear bins are used.
Table 6: Results from the spectro-polarimetric modeling of IXPE Stokes spectra across different observation epochs of the sources in $2-8$ keV energy band. The parameters, with their standard definitions, are obtained from the constant polarization model (polconst) and the energy-dependent polarization model (polpow). See text for details.
Source Epoch $A_{\rm norm}$ $A_{\rm index}$ $\psi_{\rm norm}$ $\chi^{2}_{\rm red}$ $\rm PD_{polpow}$ $\rm PA_{polpow}$ $\rm PD_{polconst}$ $\rm PA_{polconst}$ $\chi^{2}_{\rm red}$
[…]
Cyg X11-1- 1a E1 $0.020_{-0.006}^{+0.008}$ $-0.43\pm 0.25$ $-19.83\pm 2.22$ $1.01$ $3.93\pm 0.95$ $-19.83\pm 2.22$ $3.40\pm 0.26$ $-19.73\pm 2.24$ $0.91$
E2 -- -- -- -- -- -- $3.41\pm 0.44$ $-25.13\pm 3.73$ $1.09$
E3 -- -- -- -- -- -- $2.24\pm 0.61$ $-11.87\pm 7.95$ $1.15$
E4 $0.008_{-0.004}^{+0.005}$ $-0.90\pm 0.56$ $-17.74\pm 5.45$ $1.13$ $3.39\pm 0.40$ $-17.74\pm 5.45$ $2.29\pm 0.45$ $-18.33\pm 5.68$ $1.13$
E5 $0.003_{-0.002}^{+0.003}$ $-1.57\pm 0.57$ $-8.63\pm 6.16$ $1.10$ $3.95\pm 0.31$ $-8.63\pm 6.16$ $1.91\pm 0.47$ $-9.48\pm 7.18$ $1.10$
E6 $0.003\pm 0.002$ $-1.52\pm 0.68$ $-16.26\pm 5.74$ $1.12$ $3.63\pm 0.47$ $-16.26\pm 5.74$ $2.06\pm 0.45$ $-16.64\pm 6.27$ $1.13$
E7 $0.009\pm 0.005$ $-0.80\pm 0.47$ $-25.03\pm 4.56$ $1.11$ $3.23\pm 0.99$ $-25.03\pm 4.56$ $2.19\pm 0.37$ $-25.77\pm 4.69$ $1.11$
E8b -- -- -- -- -- -- $3.04\pm 0.66$ $-27.09\pm 6.24$ $1.07$
E9b -- -- -- -- -- -- $3.35\pm 0.67$ $-25.73\pm 5.79$ $1.25$
E10b -- -- -- -- -- -- $5.83\pm 0.64$ $-32.72\pm 3.15$ $0.96$
E11b -- -- -- -- -- -- $4\pm 0.7$ $-34.98\pm 5.27$ $1.05$
E12 $0.009_{-0.004}^{+0.005}$ $-0.9_{-0.4}^{+0.4}$ $-24.1\pm 3.26$ $1.03$ $3.38\pm 0.48$ $-24.1\pm 3.26$ $2.52\pm 0.48$ $-23.03\pm 3.34$ $1.05$
4U 16304747-47- 47 E1c $0.028\pm 0.003$ $-0.75\pm 0.07$ $17.90\pm 0.52$ $1.28$ $9.25\pm 0.02$ $17.90\pm 0.52$ $7.89\pm 0.14$ $17.94\pm 0.53$ $1.36$
E2 $0.022\pm 0.004$ $-0.74\pm 0.11$ $21.45\pm 0.77$ $1.13$ $8.02\pm 0.02$ $21.45\pm 0.77$ $6.27\pm 0.17$ $21.30\pm 0.78$ $1.17$
Cyg X33-3- 3a E1c $0.028\pm 0.004$ $-0.73\pm 0.09$ $-88.92\pm 0.93$ $1.46$ $8.89\pm 0.11$ $-88.92\pm 0.93$ $7.09\pm 0.24$ $-89.13\pm 0.95$ $1.79$
E2c $0.018\pm 0.002$ $-0.58\pm 0.11$ $-88.48\pm 0.97$ $1.31$ $4.43\pm 0.16$ $-88.48\pm 0.97$ $3.58\pm 0.21$ $-88.58\pm 1.01$ $1.36$
E3c $0.024\pm 0.005$ $-0.87\pm 0.12$ $-86.98\pm 3.45$ $1.32$ $9.75\pm 0.21$ $-86.98\pm 3.45$ $7.33\pm 0.31$ $-87.84\pm 1.19$ $1.63$
E4c $0.11\pm 0.03$ $-0.17\pm 0.11$ $-86.35\pm 1.34$ $1.31$ $14.33\pm 0.44$ $-86.35\pm 1.34$ $13.06\pm 1.01$ $-86.29\pm 2.19$ $1.31$
4U 1957$+115$ E1 -- -- -- -- -- -- $1.83\pm 0.33$ $-43.73\pm 5.19$ $1.09$
LMC X33-3- 3 E1 $0.011_{-0.006}^{+0.007}$ $-0.82\pm 0.5$ $-44.29\pm 3.59$ $1.12$ $4.08\pm 0.26$ $-44.29\pm 3.59$ $2.94\pm 0.37$ $-44.94\pm 3.63$ $1.12$
E2 $0.003_{-0.001}^{+0.002}$ $-1.67\pm 0.51$ $-40.47\pm 4.02$ $1.07$ $4.71\pm 0.58$ $-40.47\pm 4.02$ $2.24\pm 0.56$ $-38.21\pm 7.17$ $1.08$
Swift J1727.816131613-1613- 1613a E1 $0.022_{-0.009}^{+0.015}$ $-0.42\pm 0.36$ $2.21\pm 2.74$ $1.08$ $4.26\pm 0.21$ $2.21\pm 2.74$ $3.94\pm 0.38$ $2.02\pm 2.75$ $1.08$
E2 $0.018_{-0.006}^{+0.008}$ $-0.55\pm 0.26$ $1.53\pm 2.08$ $1.38$ $4.29\pm 0.85$ $1.53\pm 2.08$ $3.85\pm 0.28$ $1.41\pm 2.10$ $1.39$
E3 $0.021_{-0.010}^{+0.018}$ $-0.37\pm 0.23$ $-3.55\pm 3.24$ $1.04$ $3.75\pm 0.88$ $-3.55\pm 3.24$ $3.53\pm 0.39$ $-3.63\pm 3.25$ $1.04$
E4 $0.007_{-0.004}^{+0.005}$ $-1.08\pm 0.54$ $-1.21\pm 3.76$ $1.05$ $4.00\pm 0.36$ $-1.21\pm 3.76$ $3.11\pm 0.41$ $1.06$
E5 $0.004_{-0.002}^{+0.003}$ $-1.49\pm 0.54$ $-7.21\pm 4.11$ $0.97$ $4.59\pm 0.41$ $-7.21\pm 4.11$ $2.94\pm 0.45$ $-8.38\pm 4.44$ $0.98$
E8 $0.006_{-0.003}^{+0.005}$ $-1.28\pm 0.47$ $0.76\pm 4.21$ $0.97$ $4.81\pm 0.17$ $0.76\pm 4.21$ $3.22\pm 0.51$ $2.68\pm 4.54$ $0.97$
GX $339-4$d E1 -- -- -- -- -- -- $1.16\pm 0.31$ $-72.52\pm 7.78$ $1.03$
IGR J17091362417091362417091-362417091 - 3624b E1 $0.025_{-0.013}^{+0.016}$ $-0.89\pm 0.43$ $85.11\pm 3.88$ $1.04$ $10.41\pm 0.51$ $85.11\pm 3.88$ $8.24\pm 1.16$ $85.15\pm 4.03$ $1.05$
  • aOnly IXPE spectra of DU1 is fitted. Not constrained. bdiskbb is not required. cpowerlaw is not required. dSpectral fitting performed within $3-8$ keV range.

### v2

@@
 Abstract
-We present a comprehensive spectro-polarimetric and timing analysis of eleven black hole X-ray binaries, namely Cyg X11-1- 1, 4U $1630-47$, Cyg X33-3- 3, LMC X11-1- 1, 4U $1957+115$, LMC X33-3- 3, Swift J1727.816131727.816131727.8-16131727.8 - 1613, GX $339-4$, Swift J151857.0572147151857.0572147151857.0-572147151857.0 - 572147, IGR J17091362417091362417091-362417091 - 3624, and MAXI J174429417442941744-2941744 - 294, using quasi-simultaneous observations from IXPE, NICER, NuSTAR, and AstroSat.
-Timing analyses reveal characteristic type-B and type-C Quasi-periodic Oscillations (QPOs) across different spectral states, often associated with episodic radio ejections.
-Broadband ($0.7-60$ keV) spectral modelling, employing disc, Comptonization, and reflection components, reveals degeneracies in constraining disc-corona geometries.
-Polarimetric measurements in $2-8$ keV band detect significant polarization degrees (PDs) ranging from $3-20.6\%$ ($1.2-21.4\%$) in harder (softer) states, with moderate to strong energy dependence, except for LMC X11-1- 1, Swift J151857.0572147151857.0572147151857.0-572147151857.0 - 572147, and MAXI J174429417442941744-2941744 - 294, where no significant polarization is detected.
-We report the polarization detections of Cyg X33-3- 3 (PD $\sim 21.4\%$, SIMS), LMC X33-3- 3 (PD $\sim 2.4\%$, HSS) and IGR J17091362417091362417091-362417091 - 3624 (PD $\sim 9\%$, LHS) using the recent IXPE observations.
-A positive correlation is found between PD and the Comptonized photon fraction ($cov_{\rm frac}$), while an anti-correlation is observed with the disc-to-Comptonized flux ratio ($F_{\rm ratio}$) across spectral states.
-The combined timing, spectral, and polarimetric results, together with constraints from radio jet observations, suggest a radially extended corona within a truncated disc for Cyg X11-1- 1, Swift J1727.816131727.816131727.8-16131727.8 - 1613, IGR J17091362417091362417091-362417091 - 3624, and GX $339-4$, whereas the disc-corona geometry remains unconstrained for 4U $1957+115$, LMC X33-3- 3, and 4U $1630-47$.
+We present a comprehensive spectro-polarimetric and timing analysis of twelve black hole X-ray binaries, namely Cyg X-1, 4U 1630-47, Cyg X-3, LMC X-1, 4U 1957++115, LMC X-3, Swift J1727.8-1613, GX 339-4, Swift J151857.0-572147, IGR J17091-3624, MAXI J1744-294 and GRS 1915++105, using quasi-simultaneous observations from IXPE, NICER, NuSTAR, and AstroSat.
+Timing analyses reveal type-B and type-C Quasi-periodic Oscillations across different spectral states, often associated with episodic radio ejections.
+Broadband (0.7600.7-60 keV) spectral modeling, employing disc, Comptonization, and reflection components, reveals degeneracies in constraining disc-corona geometries.
+Polarimetric measurements in 282-8 keV band detect significant polarization degrees (PDs) ranging from $3-20.6\%$ ($1.2-21.4\%$) in harder (softer) states, with moderate to strong energy dependence, except for LMC X1-1, Swift J151857.0572147151857.0-572147, and MAXI J1744-294, where no significant polarization is detected.
+We report polarization detections of Cyg X3-3 (PD $\sim 21.4\%$, SIMS), LMC X3-3 (PD $\sim 2.4\%$, HSS) and IGR J17091362417091-3624 (PD $\sim 9\%$, LHS) using recent IXPE observations.
+A positive correlation is found between PD and Comptonized photon fraction ($cov_{\rm frac}$), while an anti-correlation is observed with disc-to-Comptonized flux ratio ($F_{\rm ratio}$) across spectral states.
+Combined timing, spectral, and polarimetric results, together with constraints from radio jet observations, suggest a radially extended corona within a truncated disc for Cyg X1-1, Swift J1727.816131727.8-1613, IGR J17091362417091-3624, and GX 3394339-4, whereas the disc-corona geometry remains poorly constrained for 4U $1957+115$, LMC X3-3, and 4U 1630471630-47.
 We discuss the implications of these findings for understanding accretion geometries and highlight prospects for future X-ray polarimetric studies.
@@
 accretion, accretion disc – black hole physics – polarization – techniques: polarimetric – radiation mechanisms: general – X-rays: binaries – stars: individual
-pubyear: 2025
 1 Introduction
 Black hole X-ray binaries (BH-XRBs) are believed to be the ideal candidates for understanding the physical processes that govern the radiation mechanism around the compact objects.
-BH-XRBs often exhibit distinct spectral states over timescales ranging from days to months, strongly associated with the underlying accretion dynamics of the system (Morgan et al., 1997; Zhang et al., 1997; Paul et al., 1998; Sobczak et al., 1999; Done & Zycki, 1999; Chakrabarti & Manickam, 2000; Homan et al., 2001; Corbel et al., 2001; Titarchuk & Shrader, 2002; Pottschmidt et al., 2003; Rodriguez et al., 2003; Corbel et al., 2004; Belloni et al., 2005; Kalemci et al., 2006; Remillard & McClintock, 2006; Yu & Yan, 2009; Shaposhnikov et al., 2010; Tarana et al., 2011; Nandi et al., 2012; Sriram et al., 2013; Seifina et al., 2014; Iyer et al., 2015; Yan & Yu, 2017; Sreehari et al., 2019b, 2020; Baby et al., 2020; Majumder et al., 2022; Athulya et al., 2022; Prabhakar et al., 2023; Nandi et al., 2024; Li et al., 2025; Majumder et al., 2025b, and references therein).
+BH-XRBs often exhibit distinct spectral states over timescales ranging from days to months, strongly associated with the underlying accretion dynamics of the system (Morgan et al., 1997; Done & Zycki, 1999; Chakrabarti & Manickam, 2000; Homan et al., 2001; Belloni et al., 2005; Remillard & McClintock, 2006; Nandi et al., 2012; Iyer et al., 2015; Sreehari et al., 2019b; Baby et al., 2020; Majumder et al., 2022; Athulya et al., 2022; Prabhakar et al., 2023; Nandi et al., 2024; Majumder et al., 2025b, and references therein).
 Empirically, the high-soft state (HSS) is characterized by the multi-temperature black-body emission that is likely to emerge from an optically thick and geometrically thin accretion disc Shakura & Sunyaev (1973).
@@
 The origin of these type-C QPOs is often explained through various mechanisms, including oscillations of radiative shock waves within the accretion disc (Molteni et al., 1996; Chakrabarti et al., 2008; Das et al., 2014), relativistic Lense-Thirring precession of the inner hot flow or the truncated disc (Stella & Vietri, 1998; Ingram et al., 2009), precession of small-scale jets (Ma et al., 2021) and outward drift of the truncated inner disc radius enveloped by the corona (Karpouzas et al., 2020; Bellavita et al., 2022, 2025).
-The wide array of interpretations introduces a degeneracy among the different disc-corona-jet configurations used in explaining the QPO phenomena (see Ingram & Motta 2019 for a review).
+The wide array of interpretations introduces a degeneracy among different disc-corona-jet configurations used in explaining QPO phenomena (see Ingram & Motta 2019 for a review).
 On the other hand, relatively weak type-B/type-A QPOs appear at a narrow frequency range of around $\sim 6-8$ Hz (Casella et al., 2005) during the SIMS.
@@
 Usually, soft states are characterized by less variability in the PDS without the detection of QPO like features (Belloni et al., 1999; Belloni et al., 2005; Nandi et al., 2012; Radhika & Nandi, 2014; Radhika et al., 2016).
-Furthermore, X-ray polarimetric study is also considered as a powerful diagnostic tool to infer the accretion geometry of the BH-XRBs.
+Furthermore, X-ray polarimetric study is also considered a powerful diagnostic tool to infer the accretion geometry of the BH-XRBs.
 The recent launch of IXPE Weisskopf et al.
-(2022), a polarimetric mission sensitive to low-energy ($2-8$ keV) X-rays, enables the opportunity to investigate in-depth polarimetric properties of BH-XRBs (see Dovčiak et al.
-2024 for a summary).
-So far, IXPE has observed eleven BH-XRBs, with significant polarized emission detected in nine sources, namely Cyg X11-1- 1 Krawczynski et al.
-(2022); Jana & Chang (2024); Steiner et al.
-(2024), 4U $1630-47$ Kushwaha et al.
-(2023b); Rawat et al.
-(2023a, b); Rodriguez Cavero et al.
-(2023); Ratheesh et al.
-(2024), Cyg X33-3- 3 Veledina et al.
-(2024a, b), LMC X33-3- 3 Majumder et al.
-(2024a); Svoboda et al.
-(2024a), 4U $1957+115$ Kushwaha et al.
-(2023a); Marra et al.
-(2024), Swift J1727.816131727.816131727.8-16131727.8 - 1613 Veledina et al.
-(2023); Ingram et al.
-(2024); Svoboda et al.
-(2024b); Podgorný et al.
-(2024), Swift J151857.0572147151857.0572147151857.0-572147151857.0 - 572147 Mondal et al.
-(2024), GX $339-4$ Mastroserio et al.
-(2025) and IGR J17091362417091362417091-362417091 - 3624 Ewing et al.
-(2025).
-Notably, no significant polarization is observed in LMC X11-1- 1 Podgorný et al.
-(2023).
+(2022), a polarimetric mission sensitive to low-energy (282-8 keV) X-rays, enables the opportunity to investigate in-depth polarimetric properties of BH-XRBs.
+So far, IXPE has observed twelve BH-XRBs, namely Cyg X-1, 4U 1630-47, Cyg X-3, LMC X-3, LMC X-1, 4U 1957++115, Swift J1727.8-1613, Swift J151857.0-572147, GX 339-4, IGR J17091-3624, MAXI J1744-294 and GRS 1915++105 with significant polarized emission detected in eight sources (see also Dovčiak et al., 2024).
 Despite the significant advancements in X-ray polarimetry, interpreting the observed polarization degree (PD) and polarization angle (PA) within the framework of theoretical models remains a formidable challenge.
 The classical work by Chandrasekhar (Chandrasekhar, 1960) on semi-infinite electron scattering predicts a low PD of $\sim 2\%$ from the accretion disc of highly inclined systems.
-However, 4U 1630--47 shows a remarkably higher PD of around $8.3\%$ in its disc-dominated thermal state, deviating substantially from these theoretical expectations (Kushwaha et al., 2023b; Ratheesh et al., 2024).
-Furthermore, PDs of about $4\%$ have been observed in the LHS of Cyg X--1 and Swift J1727.8--1613 (Krawczynski et al., 2022; Ingram et al., 2024), whereas existing models predict only $\sim 1\%$ PD from a wedge-shaped corona in such low-inclined systems (Krawczynski et al., 2022).
+However, 4U 1630-47 shows a remarkably higher PD of around $8.3\%$ in its disc-dominated thermal state, deviating substantially from these theoretical expectations (Kushwaha et al., 2023; Ratheesh et al., 2024).
+Similarly, soft state observations of Cyg X1-1 show polarization perpendicular to the disc plane, whereas theoretical models predict it to be parallel.
+This possibly indicates a rapidly rotating black hole in Cyg X1-1, responsible for the returning radiation effects seen in the spectro-polarimetric data (Steiner et al., 2024; Krawczynski et al., 2024).
+Furthermore, PDs of about $4\%$ have been observed in the LHS of Cyg X-1 and Swift J1727.8-1613 (Krawczynski et al., 2022; Ingram et al., 2024), whereas existing models predict only $\sim 1\%$ PD from a wedge-shaped corona in such low-inclined systems (Krawczynski et al., 2022).
 To reconcile this discrepancy, Poutanen et al.
-(2023) proposed that an outflowing corona with mildly relativistic motion could produce PD levels consistent with observations.
-It is worth mentioning that the recent detection of $\sim 9\%$ PD in the LHS of IGR J17091--3624 (Ewing et al., 2025) further indicates the limitations of current theoretical frameworks and highlights the need for more comprehensive models of X-ray polarization.
+(2023) invoked the model of Beloborodov (1998) of an outflowing corona with mildly relativistic motion to explain the high PD levels.
+It is worth mentioning that the recent detection of $\sim 9\%$ PD in the LHS of IGR J17091-3624 (Ewing et al., 2025) further indicates the limitations of current theoretical frameworks and highlights the need for more comprehensive models of X-ray polarization.
 Moreover, in certain cases, the spectro-temporal findings of BH-XRBs appear to be contradictory with interpretations derived from X-ray polarimetry.
-For instance, in Swift J1727.8--1613, the evolution of type-C QPOs suggests the presence of a jet-like corona aligned perpendicular to the disc plane (Liao et al., 2024).
+For instance, in Swift J1727.8-1613, the evolution of type-C QPOs suggests the presence of a jet-like corona aligned perpendicular to the disc plane (Liao et al., 2024).
 In contrast, simultaneous polarimetric observations of the same source indicate a radially extended corona resided at the equatorial plane of the disc (Veledina et al., 2023; Ingram et al., 2024).
-Therefore, it is evident that the polarimetric findings offer challenges to the existing X-ray spectro-temporal models, indicating that a deeper understanding of the accretion dynamics and disc-corona-jet geometry in black hole X-ray binaries is yet to be unveiled.
-Being motivated by this, we undertake in-depth timing and spectro-polarimetric analyses of eleven BH-XRBs using quasi-simultaneous archival observations from IXPE, NICER, NuSTAR and AstroSat, covering a broad energy range of $0.5-100$ keV.
-While doing so, we study the evolution of X-ray (soft/hard) and radio flux densities for all sources using multi-mission long-term monitoring data from MAXI/GSC, Swift/BAT, RATAN and VLITE.
-Further, we conduct a comprehensive spectro-temporal study of these sources using NICER, NuSTAR and AstroSat data and confirm the presence of distinct spectral states of the BH-XRBs during the observational campaigns.
-Using alternative model prescriptions relying on different assumptions on the disc-corona configurations, we attempt to explain the spectral features of the sources and investigate the possibility of having inherent degeneracy among the models.
-Next, we deduce the polarization properties of the sources using IXPE data in the $2-8$ keV energy range, followed by a detailed spectro-polarimetric correlation study.
-Subsequently, noticeable positive and negative correlations between polarization degree and spectral parameters serve as powerful diagnostics, offering deeper insights into the complex accretion dynamics and geometry of the BH-XRBs under consideration.
-Finally, we outline the future prospects of X-ray polarimetry in light of the present findings.
+Therefore, it is evident that the polarimetric findings offer challenges to the existing X-ray spectro-temporal models, indicating that a deeper understanding of the accretion dynamics and disc-corona-jet geometry in BH-XRBs is yet to be unveiled.
+Being motivated by this, we carry out in-depth timing and spectro-polarimetric analyses of eleven BH-XRBs using quasi-simultaneous archival observations from IXPE, NICER, NuSTAR and AstroSat, covering a broad energy range of 0.51000.5-100 keV.
+We conduct a comprehensive spectro-temporal study of these sources using NICER, NuSTAR and AstroSat data and confirm the presence of distinct spectral states of the BH-XRBs during the observational campaigns.
+Further, we deduce the polarization properties of the sources using IXPE data in the 282-8 keV energy range, followed by a detailed spectro-polarimetric correlation study.
+Notably, the timing and spectro-polarimetric results serve as powerful diagnostics offering deeper insights into the complex accretion dynamics and geometry of BH-XRBs under consideration.
 The paper is organized as follows.
@@
 Finally, we discuss the results and present conclusions in §5 and §6, respectively.
-Figure 1: Long-term X-ray monitoring of the sources with MAXI/GSC ($2-20$ keV), Swift/BAT ($15-50$ keV) and the corresponding hardness ratio obtained with MAXI/GSC are presented in panel (a), (b) and (c), respectively.
+Figure 1: Long-term X-ray monitoring of the sources with MAXI/GSC (2202-20 keV), Swift/BAT (155015-50 keV) and the corresponding hardness ratio obtained with MAXI/GSC are presented in panel (a), (b) and (c), respectively.
 Each data point for a given source represents averaged out respective quantity over the exposure of corresponding IXPE observation.
@@
 For a given source, the data points plotted with MJD, correspond to the respective epochs as mentioned in Table 1.
-See the text for details.
 2 Source Selection and observations
 In this work, we analyze all the IXPE (Weisskopf et al., 2022) observations of BH-XRBs conducted so far along with simultaneous/quasi-simultaneous NICER, NuSTAR and AstroSat observations depending on data availability.
-As of now, IXPE observed eleven BH-XRBs, namely Cyg X11-1- 1, 4U 1630--47, Cyg X33-3- 3, LMC X11-1- 1, 4U 1957$+115$, LMC X33-3- 3, Swift J1727.816131613-1613- 1613, Swift J151857.0572147151857.0572147151857.0-572147151857.0 - 572147, GX $339-4$, IGR J17091362417091362417091-362417091 - 3624 and MAXI J174429417442941744-2941744 - 294, during its first three and a half years of operational period.
+As of now, IXPE observed twelve BH-XRBs during its first three and a half years of operational period.
 The IXPE, NICER and NuSTAR observations used in this work are publicly available in the HEASARC [^1] https://heasarc.gsfc.nasa.gov/db-perl/W3Browse/w3browse.pl database and the AstroSat data is archived at the ISSDC [^2] https://webapps.issdc.gov.in/astro_archive/archive/Home.jsp website.
-The quasi-simultaneous multi-mission observations of the sources are separated by at most three days from the IXPE epoch except LMC X11-1- 1 for which the NuSTAR observation lies four days after the IXPE observation.
+The quasi-simultaneous multi-mission observations of the sources are separated by at most three days from the IXPE epoch except LMC X1-1 for which the NuSTAR observation lies four days after the IXPE observation.
 The sources exhibiting marginal spectro-temporal variability between observations from different missions offer an opportunity for a combined multi-mission study.
-All the observations analyzed in this study are tabulated in Table 1 along with the date and ObsID.
-Brief overview of each of the sources under consideration are presented below.
-2.1 Cyg X--1
-Cyg X-1, the first galactic BH-XRB discovered in 1971 Webster & Murdin (1972), continues to be one of the most extensively studied celestial objects.
-Recent measurements reveal that it hosts a black hole with a mass of $21.2\pm 2.2$ $M_{\odot}$ Miller-Jones et al.
-(2021), located at a distance of $2.2\pm 0.2$ kpc Miller-Jones et al.
-(2021), with a supergiant O-type star as its binary companion.
-Meanwhile, numerous studies confirm the presence of a maximally rotating ‘hole’ having spin greater than $0.99$ at the center core of the binary system (Tomsick et al., 2013; Zhao et al., 2021; Kushwaha et al., 2021).
-Cyg X--1 has remained persistently bright, mostly in the LHS, though it has transitioned to the HSS through short-lived intermediate states over the past few decades (Kushwaha et al., 2021).
-2.2 4U 1630--47
-4U 1630--47 is a recurrent X-ray transient discovered by Uhuru Jones et al.
+We mention that IXPE monitoring of GRS $1915+105$ was carried out on $10^{\rm th}$ May 2025 for about $\sim 142$ ks.
+However, the source remained below the detection level with IXPE, showing a MAXI/GSC flux of $\sim 6$ mCrab (see Fig.
+1).
+Therefore, GRS $1915+105$ is excluded from the detailed analyses in the following sections.
+All the observations analyzed in this study are tabulated in Table 1.
+A brief overview of each of the sources under consideration are presented below.
+$\bullet$ Cyg X-1: This is the first galactic BH-XRB discovered in 1971 Webster & Murdin (1972), continues to be one of the most extensively studied celestial objects.
+Recent measurements reveal that it hosts a black hole with a mass of $21.2\pm 2.2$ $M_{\odot}$, located at a distance of $2.2\pm 0.2$ kpc, with a supergiant O-type star as its binary companion (Miller-Jones et al., 2021).
+Meanwhile, numerous studies confirm the presence of a maximally rotating black hole having spin greater than $0.99$ at the center core of the binary system (Tomsick et al., 2013; Gou et al., 2014; Zhao et al., 2021; Kushwaha et al., 2021).
+Cyg X-1 has remained persistently bright, mostly in the LHS, though it has transitioned to the HSS through short-lived intermediate states over the past few decades (Kushwaha et al., 2021).
+Interestingly, significant polarized emission is detected in both LHS and HSS of the source (Krawczynski et al., 2022; Jana & Chang, 2024; Steiner et al., 2024).
+$\bullet$ 4U 1630-47: This is a recurrent X-ray transient discovered by Uhuru Jones et al.
 (1976).
@@
 (2020); Chatterjee et al.
-(2022) typically in every $600-700$ days Parmar et al.
+(2022) typically in every 600700600-700 days Parmar et al.
 (1995).
-Since the optical counterpart of the source is still unknown, the dynamical measurement of its mass and distance remains elusive.
 Several efforts relying on indirect measurements constrained its mass as $10\pm 0.1$ $M_{\odot}$ Seifina et al.
-(2014) and distance as $10-11$ kpc Seifina et al.
+(2014), distance as 101110-11 kpc Seifina et al.
 (2014); Kalemci et al.
-(2018).
-The presence of dips in the light curves of the $1996$ outburst indicates a high inclination ($i\sim 60-70^{\circ}$) of the system Tomsick et al.
+(2018) and high inclination $i\sim 60-70^{\circ}$ Tomsick et al.
 (1998); Seifina et al.
@@
 Spectral modeling Pahari et al.
-(2018) and spectro-polarimetric fitting in the high soft state confirm the source’s spin as $\sim 0.92$ Kushwaha et al.
-(2023b).
-4U 16304747-47- 47 is observed in the HSS in most of its outbursts till date, possibly due to fast transition from LHS to HSS Capitanio et al.
-(2015); Baby et al.
-(2020); Peng et al.
-(2024c); Fan et al.
-(2024); Parra et al.
+(2018) and spectro-polarimetric fitting in the high soft state confirm its spin as $\sim 0.92$ Kushwaha et al.
 (2023).
-2.3 Cyg X--3
-Cyg X33-3- 3 was discovered over five decades ago Giacconi et al.
+Notably, 4U 1630-47 is the second source observed by IXPE for which a remarkably high polarization degree ($\sim 6-8\%$) has been detected in both thermal and steep powerlaw (SPL) states (Ratheesh et al., 2024; Kushwaha et al., 2023; Rawat et al., 2023b; Rodriguez Cavero et al., 2023).
+$\bullet$ Cyg X-3: It was discovered over five decades ago Giacconi et al.
 (1967) and is a high-mass X-ray binary system hosting a compact object accreting from a Wolf-Rayet donor star van Kerkwijk et al.
@@
 (2024a).
-It is worth mentioning that Cyg X33-3- 3 exhibits exceptional radio activity Martí et al.
-(2000); Miller-Jones et al.
-(2004) which is tightly correlated with the spectral states on several occasions including the detection of innermost jet confined within the X-ray funnel-like structure Yang et al.
-(2023).
-Recent trigonometric parallax measurements from VLBA observations have precisely constrained the source’s distance as ${9.67}_{-0.48}^{+0.53}$ kpc Reid & Miller-Jones (2023).
-2.4 LMC X--1
-LMC X11-1- 1 is the first discovered extra-galactic BH-XRB located in the Large Magellanic Cloud (LMC) Mark et al.
+Recent measurements from VLBA observations have precisely constrained the distance of the source as ${9.67}_{-0.48}^{+0.53}$ kpc Reid & Miller-Jones (2023).
+$\bullet$ LMC X-1: This is the first discovered extra-galactic persistent BH-XRB located in the Large Magellanic Cloud (LMC) Mark et al.
 (1969) at a well-constrained distance of $\sim 48.1$ kpc Pietrzyński et al.
@@
 (2009).
-Unlike typical X-ray binaries, LMC X11-1- 1 consistently remains in the HSS, with thermal emission contributing around $80\%$ of the spectral flux Steiner et al.
-(2012); Bhuvana et al.
-(2021); Jana et al.
-(2021) and a luminosity of approximately $2\times 10^{38}$ erg $\rm s^{-1}$ Nowak et al.
-(2001); Bhuvana et al.
-(2021).
-The measurements using continuum fitting method in the thermally dominated state constraints LMC X11-1- 1 spin as $\sim 0.92$ Gou et al.
+The measurements using the continuum fitting method in the thermally dominated state constrain LMC X1-1 spin as $\sim 0.92$ Gou et al.
 (2009); Bhuvana et al.
 (2021), indicating the presence of a rapidly rotating black hole.
-2.5 4U 1957$+$115
-4U 1957$+$115 is a bright and persistent low-mass X-ray binary, first discovered by the Uhuru mission in $1973$ Giacconi et al.
+LMC X-1 is the first source for which a null-detection of X-ray polarization has been reported from IXPE data (Podgorný et al., 2023).
+$\bullet$ 4U 1957++115: This is a bright and persistent low-mass X-ray binary, first discovered by the Uhuru mission in $1973$ Giacconi et al.
 (1974).
-To date, it has consistently been observed in a disc-dominated spectral state, with flux levels varying between $20$ and $70$ mCrab Yaqoob et al.
-(1993); Maitra et al.
-(2014); Mudambi et al.
+The source distance is found to be in the range of 5405-40 kpc, with inclination within $\sim 13^{\circ}-78^{\circ}$ and mass in $\sim 6-10M_{\odot}$ (Maitra et al., 2014; Gomez et al., 2015).
+Nevertheless, several studies suggest the presence of a maximally spinning black hole in 4U 1957++115 Barillier et al.
 (2023).
-Despite decades of observations with various X-ray missions, key system parameters such as distance, inclination, spin, and mass remain highly uncertain.
-For instance, a high source distance of $20-40$ kpc and a low inclination ($\sim 13^{\circ}$) have been proposed for a maximum black hole mass of $\sim 6$ $M_{\odot}$ Gomez et al.
-(2015).
-In contrast, Maitra et al.
-(2014) reported a high inclination ($\sim 78^{\circ}$) and a distance between $5-10$ kpc for a black hole mass of less than $10$ $M_{\odot}$.
-Nevertheless, several studies suggest the presence of a maximally spinning black hole in 4U 1957$+$115 Barillier et al.
-(2023).
-2.6 LMC X--3
-LMC X33-3- 3 is a persistent extra-galactic BH-XRB in Large Magellanic Cloud (LMC) at a distance of $48.1$ kpc Orosz et al.
+Polarimetric study of the source reveals several relativistic effects close to the black hole, leading to low PD of $\sim 1.9\%$ (Marra et al., 2024).
+$\bullet$ LMC X-3: This is a persistent extra-galactic BH-XRB in Large Magellanic Cloud (LMC) at a distance of $48.1$ kpc Orosz et al.
 (2009), discovered by UHURU in $1971$.
@@
 (2021) is found to be present at the central core of the system.
-LMC X33-3- 3 is a relatively high inclined system ($i\sim 69.24^{\circ}\pm 0.72^{\circ}$) Orosz et al.
+LMC X3-3 is a relatively high inclined system ($i\sim 69.24^{\circ}\pm 0.72^{\circ}$) Orosz et al.
 (2014).
@@
 (2017).
-2.7 Swift J1727.8--1613
-Swift J1727.816131727.816131727.8-16131727.8 - 1613 is a recently discovered BH-XRB transient, first detected by Swift/BAT on August 24, 2023.
-Immediate monitoring of the source with MAXI/GSC in the $2-20$ keV energy range revealed an ‘unusual’ peak in X-ray flux, increasing from $150$ mCrab to $7$ Crab within a few days of detection Nakajima et al.
+Interestingly, LMC X-3 is another BH-XRB for which a relatively high polarization degree ($\sim 3\%$) has been reported from soft state observations (Majumder et al., 2024a; Svoboda et al., 2024a).
+$\bullet$ Swift J1727.8-1613: This is a recently discovered BH-XRB transient, first detected by Swift/BAT on August 24, 2023.
+Immediate monitoring of the source with MAXI/GSC in the 2202-20 keV energy range revealed an ‘unusual’ peak in X-ray flux, increasing from $150$ mCrab to $7$ Crab within a few days of detection Nakajima et al.
 (2023); Nandi et al.
 (2024).
-Radio counterpart of the source, consistent with its optical position, was detected at $5.25$ GHz with VLA and $5$ GHz with ATA, showing a rise in radio flux from $\sim 18$ to $107$ mJy six days after discovery.
 Recent measurements estimate the source distance as $\sim 2.7$ kpc, with a spin of $\sim 0.98$ and an inclination of $\sim 40^{\circ}$ Mata Sánchez et al.
@@
 (2024a).
-Further, Nandi et al.
-(2024) reported the detection of evolving type-C QPOs in hard X-rays (up to $\sim 100$ keV) using AstroSat, likely associated with the inner Comptonizing corona.
-However, NICER observations suggest a vertically elongated, jet-like corona based on the evolution of type-C QPOs Liao et al.
-(2024).
-2.8 GX 339--4
-The transient BH-XRB GX $339-4$ was discovered by the OSO77-7- 7 mission in $1973$ Markert et al.
-(1973) and typically undergoes outbursts every $2-3$ years.
-Since its discovery, it has remained one of the most extensively studied BH-XRBs due to its rich spectro-temporal variability Méndez & van der Klis (1997); Corbel et al.
-(2003); Zdziarski et al.
-(2004); Belloni et al.
-(2005); Nandi et al.
-(2012); Aneesha et al.
-(2024).
-Detailed timing studies across multiple outbursts have identified distinct temporal features, including type-A, type-B, and type-C QPOs, which are closely associated with the canonical spectral states exhibited by the source (Belloni et al., 2005; Nandi et al., 2012; Mondal et al., 2023; Zhang et al., 2024; Aneesha et al., 2024, and references therein).
-The black hole mass is constrained to be within $8.28-11.89$ $\rm M_{\odot}$ Sreehari et al.
+X-ray polarimetric constraints with IXPE in various spectral states of Swift J1727.816131727.8-1613 also aim to unveil the disc-corona geometry of the source (Veledina et al., 2023; Ingram et al., 2024; Svoboda et al., 2024b; Podgorný et al., 2024).
+$\bullet$ GX 339-4: It was discovered by the OSO7-7 mission in $1973$ Markert et al.
+(1973) and typically undergoes outbursts every 232-3 years.
+The black hole mass is constrained to be within 8.2811.898.28-11.89 $\rm M_{\odot}$ Sreehari et al.
 (2019a), while the source distance is estimated as $8.4\pm 0.9$ kpc Parker et al.
@@
 (2024).
-2.9 Swift J151857.0--572147
-Swift J151857.0572147151857.0572147151857.0-572147151857.0 - 572147 is a newly discovered Galactic BH-XRB transient, first detected by Swift/BAT during its $2024$ outburst Kennea et al.
+Further, Mastroserio et al.
+(2025) reported the alignment of X-ray polarization direction with that of radio jet orientation in GX 3394339-4.
+$\bullet$ Swift J151857.0-572147: This is a newly discovered Galactic BH-XRB transient, first detected by Swift/BAT during its $2024$ outburst Kennea et al.
 (2024).
-Follow-up radio observations with ATCA at $5.5$ and $9$ GHz confirmed the presence of bright radio flares Carotenuto & Russell (2024), generally observed for the BH-XRB transients during the transition from hard to soft spectral states.
 Subsequent observations with NICER and Insight-HXMT constrained the source distance, inclination, and spin parameter as $\sim 5.8$ kpc, $\sim 2.1^{\circ}$, and $\sim 0.84$, respectively Peng et al.
@@
 (2024).
-2.10 IGR J17091--3624
-IGR J17091362417091362417091-362417091 - 3624 is a galactic BH-XRB transient system, discovered by INTEGRAL/IBIS in $2003$ (Kuulkers et al., 2003).
-The source exhibits periodic outbursts with a quiescent period of typically four years between the successive outbursts (Capitanio et al., 2009; Krimm et al., 2011; Miller et al., 2016; Katoch et al., 2021).
-Interestingly, IGR J17091362417091362417091-362417091 - 3624 exhibits several exotic X-ray variability signatures, including the well known “heartbeat” class and therefore, often termed as the twin of GRS $1915+105$ (Capitanio et al., 2012; Radhika et al., 2018; Katoch et al., 2021; Wang et al., 2024; Shui et al., 2024).
-Despite multiple outbursts over the last two decades, the mass, inclination, and distance of the source are not well constrained due to the null-detection of its optical counterpart.
-However, several indirect measurements predict the source parameters.
+Notably, Mondal et al.
+(2024) reported low polarization for this source, whereas Ling et al.
+(2024) found a null-detection.
+$\bullet$ IGR J17091-3624: This is a galactic BH-XRB transient system, discovered by INTEGRAL/IBIS in $2003$ (Kuulkers et al., 2003).
+Interestingly, IGR J17091362417091-3624 exhibits several exotic X-ray variability signatures, including the well known “heartbeat” class and therefore, often termed as the twin of GRS $1915+105$ (Capitanio et al., 2012; Radhika et al., 2018; Katoch et al., 2021; Wang et al., 2024; Shui et al., 2024).
 Using reflection spectroscopy, the inclination of the source is estimated as $\sim 30^{\circ}-45^{\circ}$ (Xu et al., 2017; Wang et al., 2018).
-In addition, spectral modeling suggests a black hole of mass $10-12.3$ $M_{\odot}$ likely to be present at the center of IGR J17091362417091362417091-362417091 - 3624 (Iyer et al., 2015; Radhika et al., 2018).
-Based on the luminosity estimates, the distance of the source is predicted to lie within $11-17$ kpc (Rodriguez et al., 2011).
-2.11 MAXI J1744--294
-MAXI J174429417442941744-2941744 - 294, located near the Galactic center, is the most recently discovered X-ray transient by MAXI/GSC on January 2, 2025 at a flux level of $\sim 133$ mCrab (Kudo et al., 2025).
+In addition, spectral modeling suggests a black hole of mass 1012.310-12.3 $M_{\odot}$ likely to be present at the center of IGR J17091362417091-3624 (Iyer et al., 2015; Radhika et al., 2018).
+Based on the luminosity estimates, the distance of the source is predicted to lie within 111711-17 kpc (Rodriguez et al., 2011).
+Further, exceptionally high polarization degree has been reported for this source in LHS (Ewing et al., 2025; Debnath et al., 2025).
+$\bullet$ MAXI J1744-294: The sources is located near the Galactic center, is the most recently discovered X-ray transient by MAXI/GSC on January 2, 2025 at a flux level of $\sim 133$ mCrab (Kudo et al., 2025).
 The energy spectra of the source, obtained from follow-up NuSTAR observations, are found to be well described with the model comprising a thermal disc component of temperature $\sim 0.7$ keV, a power-law component of photon index $\sim 2.3$ and iron line profile at $\sim 6.6$ keV (Mandel et al., 2025).
 Based on these initial findings, the source is inferred to be a BH-XRB transient observed in the HSS (Mandel et al., 2025).
-Notably, subsequent NICER monitoring of MAXI J174429417442941744-2941744 - 294 shows an even steeper photon index of $\sim$3 (Jaisawal et al., 2025).
-Assuming a distance of $8$ kpc, the luminosity of the source is estimated to be $\sim 1.5\times 10^{37}$ erg $\rm s^{-1}$ in $2-10$ keV energy range (Mandel et al., 2025).
-Further, Grollimund et al.
-(2025) reported the detection of the radio counterpart of the source at $\sim 1.3$ GHz with MeerKAT.
+Assuming a distance of $8$ kpc, the luminosity of the source is estimated to be $\sim 1.5\times 10^{37}$ erg $\rm s^{-1}$ in 2102-10 keV energy range (Mandel et al., 2025).
 3 Data reduction
 IXPE, NASA’s first space-based low-energy X-ray polarimetry mission, was launched on December 9, 2021 (Weisskopf et al., 2022).
-It comprises three identical gas-pixel detector units (DUs) designed to measure the polarimetric properties of astrophysical sources using the principle of photo-electron tracking in the $2-8$ keV energy band.
+It comprises three identical gas-pixel detector units (DUs) designed to measure the polarimetric properties of astrophysical sources using the principle of photo-electron tracking in the 282-8 keV energy band.
 For each detected photon, the Stokes parameters (I, Q, U) are determined based on the azimuthal angle of the electric field vector, which is reconstructed by analyzing the photo-electron track within the detector plane.
-The additive nature of the Stokes parameters enables the summation of individual photon contributions to obtain the resultant I, Q, and U values, which are then used to calculate the polarization degree (PD) and polarization angle (PA).
-For the polarimetric analysis, we utilize the cleaned and calibrated level-2 event files from the three DUs of IXPE in the $2-8$ keV energy range.
+The additive nature of the Stokes parameters enables the summation of individual photon contributions to obtain the resultant I, Q, and U, which are used to calculate the polarization degree (PD) and polarization angle (PA).
+For the polarimetric analysis, we utilize the cleaned and calibrated level-2 event files from the three DUs of IXPE in the 282-8 keV energy range.
 Data reduction is performed using the latest IXPEOBSSIMv31.0.3 software (Baldini et al., 2022), following standard procedures outlined in Kislat et al.
 2015; Strohmayer 2017; Kushwaha et al.
-2023b; Majumder et al.
+2023; Majumder et al.
 2024a, b.
-The source and background regions are defined as a $60$ arcsec circular region centered at the source coordinates and an annular region with radii between $180$ and $240$ arcsec, respectively (Kushwaha et al., 2023b; Majumder et al., 2024a).
+The source and background regions are defined as a $60$ arcsec circular region centered at the source coordinates and an annular region with radii between $180$ and $240$ arcsec, respectively (Kushwaha et al., 2023; Majumder et al., 2024a).
 The XPSELECT task is used to extract source and background events from these regions.
 We use XPBIN task with various algorithms, such as PCUBE, PHA1, PHA1Q and PHA1U, to generate the necessary data products for model-dependent polarization as well as spectro-polarimetric studies, respectively.
-The latest response files provided by the software team [^3] https://ixpeobssim.readthedocs.io/en/latest/ are used during the fitting of Stokes spectra.
-The IXPE light curves of $2000$ s time bin are extracted from the event files of the DUs using XSELECT.
-NICER (Gendreau et al., 2016) is the X-ray mission onboard the International Space Station (ISS), sensitive in the soft X-ray energy band of $0.2-12$ keV and capable of excellent observations for in-depth spectro-temporal studies.
-The NICER/XTI data is reduced using the standard data extraction software (NICERDASv13444https://heasarc.gsfc.nasa.gov/docs/nicer/nicer_analysis.html) integrated within HEASOFT V6.34555https://heasarc.gsfc.nasa.gov/docs/software/lheasoft/RelNotes_6.34.html with the appropriate calibration database.
+The latest response files provided by the software team are used in the fitting of Stokes spectra.
+The IXPE light curves of $2000$ s time bin are extracted from the event files using XSELECT.
+NICER (Gendreau et al., 2016) is the X-ray mission onboard the International Space Station (ISS), sensitive in the soft X-ray energy band of 0.2120.2-12 keV and capable of excellent observations for in-depth spectro-temporal studies.
+The NICER/XTI data is reduced using the standard data extraction software (NICERDASv13) integrated within HEASOFT V6.34 with the appropriate calibration database.
 We use the tool nicerl2 with standard filtering criteria to generate the clean event files.
 The spectral products are generated using the task nicerl3-spect with background model 3c50 (Remillard et al., 2022).
-The NICER light curves of desired time resolutions are generated using nicerl3-lc task of the software routine.
-We analyze the NuSTAR (Harrison et al., 2013) data using the dedicated software NUSTARDAS [^6] https://heasarc.gsfc.nasa.gov/docs/nustar/analysis/ available in HEASOFT V6.34.
+The NICER light curves of desired time resolutions are generated using nicerl3-lc task.
+We analyze the NuSTAR (Harrison et al., 2013) data using the dedicated software NUSTARDAS available in HEASOFT V6.34.
 We use the task nupipeline to generate the cleaned event files from both FPMA and FPMB instrument onboard NuSTAR.
@@
 All the NuSTAR spectra are grouped with $25$ counts per bin to obtain better statistics during the spectral fitting.
-AstroSat (Agrawal, 2006) data are extracted using the latest reduction software LaxpcSoftv3.4.4777http://www.tifr.res.in/~astrosat_laxpc/LaxpcSoft.html.The standard procedures of AstroSat/LAXPC data extraction are performed following Antia et al.
+AstroSat (Agrawal, 2006) data are extracted using the latest reduction software LaxpcSoftv3.4.4.The standard procedures of AstroSat/LAXPC data extraction are performed following Antia et al.
 2017, 2021; Antia et al.
@@
 2022.
-We use LAXPC20 data for generating the light curves in different energy ranges for the respective sources except Swift J1727.816131727.816131727.8-16131727.8 - 1613, for which LAXPC10 data in low gain mode was available for analysis (see Nandi et al., 2024, for details).
+We use LAXPC20 data for generating the light curves in different energy ranges for the respective sources except Swift J1727.816131727.8-1613, for which LAXPC10 data in low gain mode was available for analysis (see Nandi et al., 2024, for details).
 Note that, level-1 data corresponding to single events from all layers of the detector are considered in each observation.
 We utilize on-demand [^3] http://maxi.riken.jp/mxondem/ processed MAXI/GSC (Matsuoka et al., 2009) data and scaled map transient analysis data [^4] https://swift.gsfc.nasa.gov/results/transients/ of Swift/BAT (Krimm et al., 2013) for the individual sources in different energy ranges.
-Note that, MAXI/GSC count rates in $2-20$ keV energy range are converted into the flux unit (Crab) following the relation $1$ Crab $=3.8$ photons $\rm cm^{-2}$ $\rm s^{-1}$ as suggested by the instrument team [^10] http://maxi.riken.jp/top/readme.html.
-Similarly, for Swift/BAT data in the $15-50$ keV energy range, we obtain the corresponding photon flux [^11] https://swift.gsfc.nasa.gov/results/transients/ in units of mCrab, where $1$ mCrab $=0.000220$ counts $\rm cm^{-2}$ $\rm s^{-1}$.
+Note that, MAXI/GSC count rates in 2202-20 keV energy range are converted into the flux unit (Crab) following the relation $1$ Crab =3.8=3.8 photons $\rm cm^{-2}$ $\rm s^{-1}$ as suggested by the instrument team [^5] http://maxi.riken.jp/top/readme.html.
+Similarly, for Swift/BAT data in the 155015-50 keV energy range, we obtain the corresponding photon flux [^6] https://swift.gsfc.nasa.gov/results/transients/ in units of mCrab, where $1$ mCrab =0.000220=0.000220 counts $\rm cm^{-2}$ $\rm s^{-1}$.
 4 Analysis and Results
-In this section, we present the comprehensive analyses, modeling, and observational results of the BH-XRBs under consideration, observed with IXPE.
-This study incorporates quasi-simultaneous observations from NICER, NuSTAR and AstroSat to enable wide-band spectro-temporal investigations.
-Furthermore, we explore the polarimetric properties with IXPE and examine the spectro-polarimetric correlations for these sources.
 4.1 Temporal Analysis
-Figure 2: Results of temporal (top panel) and spectral (bottom panel) analyses of ten BH-XRBs in different spectral states obtained using quasi-simultaneous IXPE, NICER, NuSTAR and AstroSat observations.
-Top panels: PDS obtained from NICER ($0.5-10$ keV) and/or NuSTAR/AstroSat observations in different energy ranges are depicted for the respective sources in various epochs.
-The power spectrum of GX $339-4$ from NICER observation is rescaled by a factor of $5$.
+Figure 2: Results of temporal (top panel) and spectral (bottom panel) analyses of ten BH-XRBs in different spectral states (color coded) obtained using quasi-simultaneous IXPE, NICER, NuSTAR and AstroSat observations.
+Top panels: PDS obtained from NICER (0.5100.5-10 keV) and/or NuSTAR/AstroSat observations in different energy ranges are depicted for the respective sources in various epochs.
+The power spectrum of GX 3394339-4 from NICER observation is rescaled by a factor of $5$.
 Bottom panels: Energy spectra of the sources from NICER and/or NuSTAR data are shown during the respective epochs.
@@
 In Fig.
-1, we present the count rate obtained with MAXI/GSC ($2-20$ keV) and Swift/BAT ($15-50$ keV) monitoring in panels (a) and (b) along with the variation of hardness ratio ($HR$) from MAXI/GSC in panel (c), respectively, for all sources under consideration.
+1, we present the count rate obtained with MAXI/GSC (2202-20 keV) and Swift/BAT (155015-50 keV) monitoring in panels (a) and (b) along with the variation of hardness ratio ($HR$) from MAXI/GSC in panel (c), respectively, for all sources under consideration.
+Here, $HR$ is defined as the flux ration in 6206-20 KeV and 262-6 keV.
 Note that, flux values associated with the count rates in soft (hard) energy bands with MAXI/GSC (Swift/BAT) are also mentioned in units of mCrab in the $y$-axis (right side) of the respective panels.
@@
 1a-c, we present the average values of the respective quantities over the entire IXPE exposure of each epoch.
-Interestingly, for Swift J1727.816131727.816131727.8-16131727.8 - 1613, we observe that both MAXI/GSC and Swift/BAT flux decrease as its outburst progresses.
+Interestingly, for Swift J1727.816131727.8-1613, we observe that both MAXI/GSC and Swift/BAT flux decrease as its outburst progresses.
 The $HR$ also sharply decreases and anti-correlates with the accompanied radio flux density (see Fig.
@@
 Moreover, these variability features resemble the ‘canonical’ state transition as HIMS $\rightarrow$ SIMS $\rightarrow$ HSS $\rightarrow$ LHS (decay), commonly exhibited by BH-XRB transients (see Nandi et al., 2012, for details).
-Similarly, for Cyg X11-1- 1, we observe a marginal increase in the MAXI/GSC count rate and a corresponding decrease in the Swift/BAT count rate during epochs E3 to E7.
+Similarly, for Cyg X1-1, we observe a marginal increase in the MAXI/GSC count rate and a corresponding decrease in the Swift/BAT count rate during epochs E3 to E7.
 In the subsequent observations (E8 to E11), the MAXI/GSC count rate decreases while the Swift/BAT count rate rises.
@@
 This trend possibly indicates the spectral state transitions between LHS (E1-E2, E8-E11) and HSS (E3-E7, E12), respectively (see Table 1).
-However, only marginal variations are seen in both low- and high-energy count rates for several sources, such as LMC X11-1- 1, LMC X33-3- 3, 4U $1957+115$, GX $339-4$, Swift J151857.0572147151857.0572147151857.0-572147151857.0 - 572147 and MAXI J174429417442941744-2941744 - 294.
+However, only marginal variations are seen in both low- and high-energy count rates for several sources, such as LMC X1-1, LMC X3-3, 4U $1957+115$, GX 3394339-4, Swift J151857.0572147151857.0-572147 and MAXI J17442941744-294.
 For all of them, the hardness ratio remains low ($HR\lesssim 0.1$) and these sources are also detected at flux levels below approximately $100$ mCrab.
 These possibly suggest that the sources are likely in softer spectral states.
-In contrast, Cyg X33-3- 3 exhibits a clear sequence of state transitions, evolving from the LHS to the HSS, then to the SIMS, and back to HSS.
+In contrast, Cyg X3-3 exhibits a clear sequence of state transitions, evolving from the LHS to the HSS, then to the SIMS, and back to HSS.
 This evolution is accompanied by moderate changes in count rates and hardness ratio, and is likely associated with strong radio flares ($\sim 100$ mJy) (see Table 1, Fig.
@@
 2024a).
-For 4U $1630-47$, HSS (Kushwaha et al., 2023b) and SPL (Rodriguez Cavero et al., 2023) states are observed with reasonable variations in the count rates and $HR$.
-It is worth noting that IGR $17091-3624$ is observed in the LHS at $\sim 1$ Crab MAXI/GSC flux with $HR\sim 0.4$, very similar to the LHS observations of Cyg X11-1- 1.
+For 4U 1630471630-47, HSS (Kushwaha et al., 2023) and SPL (Rodriguez Cavero et al., 2023) states are observed with reasonable variations in the count rates and $HR$.
+It is worth noting that IGR 17091362417091-3624 is observed in the LHS at $\sim 1$ Crab MAXI/GSC flux with $HR\sim 0.4$, very similar to the LHS observations of Cyg X1-1.
+Furthermore, GRS $1915+105$ is observed with IXPE in its present faint and obscured state with a flux level of $\sim 6$ mCrab only (see Fig.
+1).
+We find that observations in harder spectral states (LHS and HIMS) and relatively softer spectral states (HSS, SIMS and SPL) of the sources are separated in two distinct regions of $HR$ variations (see Fig.
+1).
+In particular, $HR\sim 0.22$ separates two regions, with $HR>0.22$ indicating harder and $HR<0.22$ corresponds to softer spectral states with few exceptions.
 4.1.2 Power Density Spectra
-We investigate the power density spectra (PDS) of all sources in different spectral states using NICER observations (see Table 1) in $0.5-10$ keV energy range.
-In addition, PDS properties in hard X-rays ($>10$ keV) are also examined using available NuSTAR and AstroSat observations of the sources.
+We investigate the power density spectra (PDS) of all sources in different spectral states using NICER observations (see Table 1) in 0.5100.5-10 keV energy range.
+In addition, PDS properties in hard X-rays (>10>10 keV) are also examined using available NuSTAR and AstroSat observations of the sources.
 We generate 0.01 s time binned light curves and compute the PDS up to Nyquist frequency of $50$ Hz with $1024$ newbins per interval using powspec inside HEASOFT V6.34.
@@
 We model each PDS with a combination of constant and multiple Lorentzian components.
-Further, we compute the total percentage rms amplitude ($rms_{\rm tot}\%$) as a measure of variability in $0.1-50$ Hz and tabulate it in Table 3.
+Further, we compute the total percentage rms amplitude ($rms_{\rm tot}\%$) as a measure of variability in 0.1500.1-50 Hz and tabulate it in Table 3.
 Interestingly, we observe that the PDS in several spectral states exhibit distinct variable features.
-In particular, PDS of Cyg X11-1- 1 in the LHS show significant variability with $rms_{\rm tot}\%\sim 31.1-52.4$ ($0.5-10$ keV) and a clear power spectral break at $\sim 1.5$ Hz with NICER observations.
-Further, power spectral properties in hard X-rays ($3-80$ keV) with quasi-simultaneous AstroSat observations exhibit similar break frequencies ($\sim 1.5$ Hz) with $rms_{\rm tot}\%\sim 37$ in LHS (see Table 1 and Fig.
+In particular, PDS of Cyg X1-1 in the LHS show significant variability with $rms_{\rm tot}\%\sim 31.1-52.4$ (0.5100.5-10 keV) and a clear power spectral break at $\sim 1.5$ Hz with NICER observations.
+Further, power spectral properties in hard X-rays (3803-80 keV) with quasi-simultaneous AstroSat observations exhibit similar break frequencies ($\sim 1.5$ Hz) with $rms_{\rm tot}\%\sim 37$ in LHS (see Table 1 and Fig.
 2a).
-We also observe marginal (significant) PDS variability of $rms_{\rm tot}\%\sim 9.1$ ($33$) with low (high) energy NICER (AstroSat) observations of Cyg X11-1- 1 in the HSS.
+We also observe marginal (significant) PDS variability of $rms_{\rm tot}\%\sim 9.1$ ($33$) with low (high) energy NICER (AstroSat) observations of Cyg X1-1 in the HSS.
 Note that similar variability properties including power spectral breaks were also reported by Kushwaha et al.
-(2021) in LHS and HSS of Cyg X11-1- 1.
-For Swift J1727.816131727.816131727.8-16131727.8 - 1613, the presence of type-C QPO feature at $\sim 1.4$ Hz is observed during HIMS (epoch 1, hereafter E1) in $0.5-100$ keV with NICER, NuSTAR and AstroSat quasi-simultaneous observations (see also Nandi et al.
+(2021) in LHS and HSS of Cyg X1-1.
+For Swift J1727.816131727.8-1613, the presence of type-C QPO feature at $\sim 1.4$ Hz is observed during HIMS (epoch 1, hereafter E1) in 0.51000.5-100 keV with NICER, NuSTAR and AstroSat quasi-simultaneous observations (see also Nandi et al.
 2024; Liao et al.
@@
 The evolution of the QPO frequency up to $\sim 6.7$ Hz is seen with NICER in the later epochs (E2-E5).
-However, a power distribution, similar to Cyg X11-1- 1 (LHS), is observed in the LHS of Swift J1727.816131727.816131727.8-16131727.8 - 1613 without the presence of power spectral break, whereas marginal variability with $rms_{\rm tot}\%\sim 1.4$ is noticed during HSS (E6-E7).
-The remaining sources (4U $1630-47$, LMC X11-1- 1, LMC X33-3- 3, 4U $1957+115$, Cyg X33-3- 3 and Swift J151857.0572147151857.0572147151857.0-572147151857.0 - 572147) in SIMS and HSS display weak variability signatures mostly represented by constant noise distribution and occasionally marginal power variation at lower frequencies.
-Conversely, during the recurrent outburst of GX $339-4$, a sharp type-B QPO peak at $\sim 4.5$ Hz ($rms_{\rm QPO}\%\sim 10.2$) and harmonic feature at $\sim 9.5$ Hz along with weak broadband noise ($rms_{\rm tot}\%\sim 15.4$) are observed in SIMS (see Fig.
+However, a power distribution, similar to Cyg X1-1 (LHS), is observed in the LHS of Swift J1727.816131727.8-1613 without the presence of power spectral break, whereas marginal variability with $rms_{\rm tot}\%\sim 1.4$ is noticed during HSS (E6-E7).
+The remaining sources (4U 1630471630-47, LMC X1-1, LMC X3-3, 4U $1957+115$, Cyg X3-3 and Swift J151857.0572147151857.0-572147) in SIMS and HSS display weak variability signatures mostly represented by constant noise distribution and occasionally marginal power variation at lower frequencies.
+Conversely, during the recurrent outburst of GX 3394339-4, a sharp type-B QPO peak at $\sim 4.5$ Hz ($rms_{\rm QPO}\%\sim 10.2$) and harmonic feature at $\sim 9.5$ Hz along with weak broadband noise ($rms_{\rm tot}\%\sim 15.4$) are observed in SIMS (see Fig.
 2 and Aneesha et al.
 2024).
-Note that, the QPO features with enhanced $rms_{\rm QPO}\%$ from $3.76$ ($0.5-10$ keV) to $12.6$ ($3-80$ keV) are observed with NICER and AstroSat, respectively.
-For IGR J17091362417091362417091-362417091 - 3624, a sharp type-C QPO at $\sim 0.2$ Hz ($rms_{\rm QPO}\%\sim 15$) along with a weak harmonic at $\sim 0.4$ Hz are observed during the onset phase of the recent outburst (see Fig.
+Note that, the QPO features with enhanced $rms_{\rm QPO}\%$ from $3.76$ (0.5100.5-10 keV) to $12.6$ (3803-80 keV) are observed with NICER and AstroSat, respectively.
+For IGR J17091362417091-3624, a sharp type-C QPO at $\sim 0.2$ Hz ($rms_{\rm QPO}\%\sim 15$) along with a weak harmonic at $\sim 0.4$ Hz are observed during the recent outburst (see Fig.
 2).
@@
 4.2.1 Spectral Modeling
-Figure 3: Variation of the data to model ratio of the HSS/HIMS spectra of Cyg X11-1- 1, GX $339-4$ and Swift J1727.816131727.816131727.8-16131727.8 - 1613 fitted with model combination M1 and M3.
+Figure 3: Variation of the data to model ratio of the HSS/HIMS spectra of Cyg X1-1, GX 3394339-4 and Swift J1727.816131727.8-1613 fitted with model combination M1 and M3.
 The improvement in the respective fits by incorporating the reflection model component over M1 is depicted in panels (b), (d) and (f).
@@
 4.2.2 Alternative Spectral Models
-We aim to explore the possible alternative disc-corona geometries in explaining the high-energy tail of the observed spectra.
-To achieve this, we use the comptt (Titarchuk, 1994) component as a proxy for thcomp in modeling thermal Comptonization.
-Note that, the ‘geometry switch’ selected inside comptt model as ‘$\leq 1$’ enables a disc-like geometry for the Comptonizing corona.
-Considering this, we proceed in the spectral fitting of all sources with the model combination constant×$\times$×Tbabs×$\times$(diskbb + comptt) (hereafter M2).
-Interestingly, we find that the model M2 provides a fit equally as good as M1 for all observations under consideration except a few observations of Cyg X11-1- 1 (E6 and E7), Swift J1727.816131727.816131727.8-16131727.8 - 1613 (E4 and E5) and GX $339-4$ (E1).
-This leads to a possible degeneracy between the selected models in describing the observed spectra and the geometry of the Comptonizing region.
+[…]
 […]
@@
 The variation of the spectral parameters over different observational epochs indicates the presence of distinct spectral states of the sources under consideration.
-Figure 4: Confidence contours ($3\sigma$) in the PD–PA plane obtained from the model-independent polarization measurements using $2-8$ keV IXPE data, combining all DUs across various spectral states of the sources.
+Figure 4: Confidence contours ($3\sigma$) in the PD–PA plane obtained from the model-independent polarization measurements using 282-8 keV IXPE data, combining all DUs across various spectral states of the sources.
 The epoch of each observation is indicated in parentheses next to the corresponding spectral state.
-For Swift J151857.0572147151857.0572147151857.0-572147151857.0 - 572147, LMC X11-1- 1 and MAXI J174429417442941744-2941744 - 294, null-detection of polarization is observed, whereas all other sources exhibit significant polarization.
-The inset highlights a zoomed-in region of the PD–PA space for better clarity.
-See the text for details.
-We find that the effects of Comptonization are reasonably high during the LHS of Cyg X11-1- 1 with covering fraction $cov_{\rm frac}\sim 0.61_{-0.04}^{+0.04}-0.77_{-0.02}^{+0.02}$, where $cov_{\rm frac}$ measures the fraction of seed photons that are being up-scattered by thermal electrons of the corona.
+For Swift J151857.0572147151857.0-572147, LMC X1-1 and MAXI J17442941744-294, null-detection of polarization is observed, whereas all other sources exhibit significant polarization.
+Inset highlights a zoomed-in region of PD–PA space for better clarity.
+We find that the effects of Comptonization are noticeably high during the LHS of Cyg X-1 with covering fraction $cov_{\rm frac}\sim 0.61_{-0.04}^{+0.04}-0.77_{-0.02}^{+0.02}$, indicating Comptonization of up to $\sim 77\%$ of the soft photons in the LHS spectra.
 In addition, a high Comptonized flux contribution of $73-80\%$ is also observed in the spectra (see Table 3).
-Indeed, such a high $cov_{\rm frac}$ results in low photon index ($\Gamma_{\rm th}\lesssim 1.64_{-0.02}^{+0.02}$) as observed during the LHS observations of Cyg X11-1- 1.
-In the HIMS and LHS of Swift J1727.816131727.816131727.8-16131727.8 - 1613, we find $cov_{\rm frac}\sim 0.35_{-0.01}^{+0.01}-0.82_{-0.02}^{+0.02}$ and $\Gamma_{\rm th}\sim 1.71_{-0.02}^{+0.03}-2.11_{-0.04}^{+0.05}$ with up to $\sim 64\%$ Comptonized emission.
-Similarly, LHS of IGR J17091362417091362417091-362417091 - 3624 which is dominated by the Comptonized emission ($\sim 87\%$ flux contribution) yields harder ($\Gamma_{\rm th}\sim 1.62_{-0.01}^{+0.02}$) spectral characteristics.
-On the other hand, the HSS/SIMS spectra of Cyg X11-1- 1 and Swift J1727.816131727.816131727.8-16131727.8 - 1613 results in low Comptonization with $cov_{\rm frac}$ in the range $0.005_{-0.002}^{+0.002}-0.38_{-0.03}^{+0.04}$ and thermal disc emission with $39-98\%$ flux contribution along with a steeper spectral index ($\Gamma_{\rm th}\sim 2.32_{-0.04}^{+0.06}-5.14_{-0.39}^{+0.36}$).
-Moreover, HSS with $76-98\%$ disc contribution are observed for LMC X11-1- 1, 4U $1630-47$, LMC X33-3- 3 and 4U $1957+115$ with negligible Comptonized emission ($cov_{\rm frac}<0.1$) as shown in Table 3.
-We find that GX $339-4$ (Swift J151857.0572147151857.0572147151857.0-572147151857.0 - 572147) exhibits intermediate spectral characteristics with $cov_{\rm frac}\sim 0.34$ ($0.21$) and $\Gamma_{\rm th}\sim 2$ ($2.5$) with $\sim 50\%$ ($78\%$) disc emission during SIMS (HSS).
-The reflection spectroscopy of Cyg X11-1- 1, Swift J1727.816131727.816131727.8-16131727.8 - 1613 and GX $339-4$ with model M3 during the HSS/SIMS results in disc ionization parameter ($\log\xi$) in the range of $3.31_{-0.05}^{+0.05}-4.29_{-0.21}^{+0.14}$ erg cm $\rm s^{-1}$, iron abundance ($A_{\rm fe}$) between $\sim 4.61_{-0.31}^{+0.36}$ and $6.35_{-1.46}^{+2.08}$ $A_{\odot}$, and reflection fraction ($R_{\rm f}$) ranging $\sim 0.59_{-0.11}^{+0.05}$ to $3.16_{-0.33}^{+0.35}$.
+Indeed, such a high $cov_{\rm frac}$ results in low photon index ($\Gamma_{\rm th}\lesssim 1.64_{-0.02}^{+0.02}$) as observed during the LHS observations of Cyg X1-1.
+In the HIMS and LHS of Swift J1727.8-1613, we find $cov_{\rm frac}\sim 0.35_{-0.01}^{+0.01}-0.82_{-0.02}^{+0.02}$ and $\Gamma_{\rm th}\sim 1.71_{-0.02}^{+0.03}-2.11_{-0.04}^{+0.05}$ with up to $\sim 64\%$ Comptonized emission.
+Similarly, LHS of IGR J17091-3624 which is dominated by the Comptonized emission ($\sim 87\%$ flux contribution) yields harder ($\Gamma_{\rm th}\sim 1.62_{-0.01}^{+0.02}$) spectral characteristics.
+On the other hand, the HSS/SIMS spectra of Cyg X-1 and Swift J1727.8-1613 results in low Comptonization with $cov_{\rm frac}$ in the range $0.005_{-0.002}^{+0.002}-0.38_{-0.03}^{+0.04}$ and thermal disc emission with $39-98\%$ flux contribution along with a steeper spectral index ($\Gamma_{\rm th}\sim 2.32_{-0.04}^{+0.06}-5.14_{-0.39}^{+0.36}$).
+Moreover, HSS with $76-98\%$ disc contribution are observed for LMC X-1, 4U 1630-47, LMC X-3 and 4U 1957++115 with negligible Comptonized emission ($cov_{\rm frac}<0.1$) as shown in Table 3.
+We find that GX 339-4 (Swift J151857.0-572147) exhibits intermediate spectral characteristics with $cov_{\rm frac}\sim 0.34$ ($0.21$) and $\Gamma_{\rm th}\sim 2$ ($2.5$) with $\sim 50\%$ ($78\%$) disc emission during SIMS (HSS).
+The reflection spectroscopy of Cyg X1-1, Swift J1727.816131727.8-1613 and GX 3394339-4 with model M3 during the HSS/SIMS results in disc ionization parameter ($\log\xi$) in the range of $3.31_{-0.05}^{+0.05}-4.29_{-0.21}^{+0.14}$ erg cm $\rm s^{-1}$, iron abundance ($A_{\rm fe}$) between $\sim 4.61_{-0.31}^{+0.36}$ and $6.35_{-1.46}^{+2.08}$ $A_{\odot}$, and reflection fraction ($R_{\rm f}$) ranging $\sim 0.59_{-0.11}^{+0.05}$ to $3.16_{-0.33}^{+0.35}$.
 Here, $A_{\odot}$ represents the solar iron abundance.
@@
 In addition, we find that the model component relxilllp (Dauser et al., 2013) also delineates the reflection features assuming the primary source emissivity profile from a lamppost corona.
-Moreover, another reflection flavor relxillCp, in which the spectrum of the coronal emission is computed from the nthComp (Zdziarski et al., 1996) model of spherical geometry, also fits the data for Cyg X11-1- 1 (E6 and E7) and Swift J1727.816131727.816131727.8-16131727.8 - 1613 (E4).
+Moreover, another reflection flavor relxillCp, in which the spectrum of the coronal emission is computed from the nthComp (Zdziarski et al., 1996) model of spherical geometry, also fits the data for Cyg X1-1 (E6 and E7) and Swift J1727.816131727.8-1613 (E4).
 Therefore, it is evident that the reflection spectroscopy also results in degeneracy among several models relying on contemporary coronal geometries.
@@
 This highlights the importance of incorporating X-ray polarimetric observations into the study of BH-XRBs.
-Subsequently, we explore how polarization measurements can offer deeper insights into this issue.
+Subsequently, we explore how polarization measurements offer deeper insights to this issue.
 4.3 Polarization Measurements
@@
 In general, the polarization measurements are performed using two distinct methods.
-The first one relies on the model-independent approach based on the analysis of the polarization cube (Kislat et al., 2015) using the dedicated python package IXPEOBSSIMv31.0.3121212https://ixpeobssim.readthedocs.io/en/latest/ (Baldini et al., 2022).
+The first one relies on the model-independent approach based on the analysis of the polarization cube (Kislat et al., 2015) using the dedicated python package IXPEOBSSIMv31.0.3777https://ixpeobssim.readthedocs.io/en/latest/ (Baldini et al., 2022).
 This provides the polarimetric properties by computing the polarization degree (PD) and polarization angle (PA) from the normalized Stokes parameters without any prior assumption on the emission mechanism.
@@
 Following Kushwaha et al.
-2023b; Majumder et al.
+2023; Majumder et al.
 2024a, b, we consider all events from the three DUs of IXPE in the analysis.
-Subsequently, the normalized Q and U Stokes parameters, the polarization degree $({\rm PD}=\sqrt{(Q/I)^{2}+(U/I)^{2}})$ and polarization angle $({\rm PA}=\frac{1}{2}\tan^{-1}(U/Q))$ are computed in $2-8$ keV energy band for all observations of the respective sources.
+Subsequently, the normalized Q and U Stokes parameters, the polarization degree $({\rm PD}=\sqrt{(Q/I)^{2}+(U/I)^{2}})$ and polarization angle $({\rm PA}=\frac{1}{2}\tan^{-1}(U/Q))$ are computed in 282-8 keV energy band for all observations of the respective sources.
 Following the guidelines of IAU [^8] https://aas.org/posts/news/2015/12/iau-calls-consistency-use-polarization-angle (Contopoulos & Jappel, 1974), we adopt the convention that PA increases counter-clockwise from North to East direction in the sky.
@@
 4, we present $3\sigma$ confidence contours associated with the measurements of PD and PA for the respective sources in different spectral states.
-We find significant degree of polarized emission ($\rm PD>\rm MDP_{\rm 99}\%$) in all sources except LMC X11-1- 1, Swift J151857.0572147151857.0572147151857.0-572147151857.0 - 572147 and MAXI J1744--294.
-More precisely, we find that PD varies within $1.22\pm 0.35\%-21.41\pm 0.41\%$ including all sources in different spectral states within $2-8$ keV energy band.
-We note that Cyg X33-3- 3 shows the maximum degree of polarization, whereas GX $339-4$ exhibits the lowest PD ($>\rm MDP_{99}\%$) among all sources under consideration.
+We find significant degree of polarized emission ($\rm PD>\rm MDP_{\rm 99}\%$) in all sources except LMC X1-1, Swift J151857.0572147151857.0-572147 and MAXI J1744-294.
+More precisely, we find that PD varies within $1.22\pm 0.35\%-21.41\pm 0.41\%$ including all sources in different spectral states within 282-8 keV energy band.
+We note that Cyg X3-3 shows the maximum degree of polarization, whereas GX 339-4 exhibits the lowest PD ($\rm MDP_{99}\%$) among all sources under consideration.
 Further, a significant variation of PD is observed over different spectral states of the sources.
-For example, Cyg X11-1- 1 manifests a maximum PD of $\sim 4.8\%$ in the LHS, which decreases to $\sim 1.4\%$ in the HSS.
-Interestingly, a similar variation of PD is observed over several spectral states during the outburst phase of Swift J1727.816131727.816131727.8-16131727.8 - 1613 in which the polarization degree drops down to $\sim 0.4\%$ (HSS) from $\sim 4.7\%$ (HIMS) and again increases to $\sim 3.2\%$ in the LHS.
-Most astonishingly, 4U $1630-47$ exhibits exceedingly high PD in the HSS ($\sim 8.3\%$) and SPL ($\sim 6.8\%$), respectively.
-Moreover, we find an unprecedentedly high PD of $\sim 9\%$ in the LHS of the recent outburst of IGR J17091362417091362417091-362417091 - 3624, which is the maximum reported for any confirmed BH-XRB observed with IXPE to date (see also Ewing et al.
+For example, Cyg X1-1 manifests a maximum PD of $\sim 4.8\%$ in the LHS, which decreases to $\sim 1.4\%$ in the HSS.
+Interestingly, a similar variation of PD is observed over several spectral states during the outburst phase of Swift J1727.816131727.8-1613 in which the polarization degree drops down to $\sim 0.4\%$ (HSS) from $\sim 4.7\%$ (HIMS) and again increases to $\sim 3.2\%$ in the LHS.
+Most astonishingly, 4U 1630471630-47 exhibits exceedingly high PD in the HSS ($\sim 8.3\%$) and SPL ($\sim 6.8\%$), respectively.
+Moreover, we find an unprecedentedly high PD of $\sim 9\%$ in the LHS of the recent outburst of IGR J17091362417091-3624, which is the maximum reported for any confirmed BH-XRB observed with IXPE to date (see also Ewing et al.
 2025).
-In contrast, 4U $1957+115$ and LMC X33-3- 3 are found to show $\sim 1.9\%$ and $\sim 2.4-3\%$ polarization in the HSS, respectively.
-It may be noted that, for GX $339-4$, significant PD ($\sim 1.2\%$, E1) is obtained in $3-8$ keV energy range (see also Mastroserio et al.
-2025), whereas a null-detection is seen considering the entire IXPE energy band of $2-8$ keV.
-Moreover, we find that the PA of the sources in different spectral states shows marginal variation (see Table 5).
+In contrast, 4U 1957++115 and LMC X-3 are found to show $\sim 1.9\%$ and $\sim 2.4-3\%$ polarization in the HSS, respectively.
+It may be noted that, for GX 339-4, significant PD ($\sim 1.2\%$, E1) is obtained in 3-8 keV energy range (see also Mastroserio et al., 2025), whereas null-detection is seen considering the entire IXPE energy band of 2-8 keV.
+Moreover, we find that PA of the sources in different spectral states shows marginal variation (see Table 5).
 Null-detection of Polarization: It is worth mentioning that we re-confirm the null-detection of polarized emission ($\rm PD<\rm MDP_{\rm 99}\%$) for a few sources under consideration.
-In particular, LMC X11-1- 1 remains the first source that fails to manifest significant polarization with IXPE (see Table 5 and Podgorný et al.
+In particular, LMC X1-1 remains the first source that fails to manifest significant polarization with IXPE (see Table 5 and Podgorný et al.
 2023).
-Further, Swift J151857.0572147151857.0572147151857.0-572147151857.0 - 572147 exhibits a PD of $\sim 0.25\%$ in $2-8$ keV which remains significantly low from the corresponding $\rm MDP_{\rm 99}\sim 0.78\%$, confirming a null-detection.
+Further, Swift J151857.0572147151857.0-572147 exhibits a PD of $\sim 0.25\%$ in 282-8 keV which remains significantly low from the corresponding $\rm MDP_{\rm 99}\sim 0.78\%$, confirming a null-detection.
 This finding contradicts the polarization detection with PD $\sim 1.3\%$ as reported by Mondal et al.
@@
 This discrepancy possibly arises as the present analysis is carried out using the latest version of the software IXPEOBSSIMv31.0.3, fixing the bugs related to the visualization of polarization cubes in the previous releases [^9] https://ixpeobssim.readthedocs.io/en/latest/release_notes.html.
-We emphasize that our results are consistent with the recent polarization measurements of Swift J151857.0572147151857.0572147151857.0-572147151857.0 - 572147 Ling et al.
+We emphasize that our results are consistent with the recent polarization measurements of Swift J151857.0572147151857.0-572147 Ling et al.
 (2024).
-Further, the null-detection of polarization is also observed in a few observations of Swift J1727.816131727.816131727.8-16131727.8 - 1613 (E6, E7), GX $339-4$ (E2) and LMC X33-3- 3 (E3).
-In addition, for the first time, we report the null-detection of X-ray polarization with PD ($\sim 0.7\%$) $<\rm MDP_{\rm 99}(1.3\%)$ in the recently discovered BH-XRB candidate MAXI J174429417442941744-2941744 - 294 during HSS.
+Further, the null-detection of polarization is also observed in a few observations of Swift J1727.816131727.8-1613 (E6, E7), GX 3394339-4 (E2) and LMC X3-3 (E3).
+In addition, for the first time, we report the null-detection of X-ray polarization with PD ($\sim 0.7\%$) $\rm MDP_{\rm 99}(1.3\%)$ in the recently discovered BH-XRB candidate MAXI J17442941744-294 during HSS.
 4.4 Energy-dependent Polarization
+Refer to captionRefer to caption
 Figure 5: Variation of polarization degree (PD) with energy.
 In each panel, source name and spectral state (observed during different epoch) are marked.
-The histograms in colors represent the minimum detectable polarization at $99\%$ confidence ($\rm MDP_{99}\%$).
+The significance of energy variation of PD for the respective cases are mentioned in units of $\sigma$ in the plot legends.
+The histograms represent the minimum detectable polarization at $99\%$ confidence ($\rm MDP_{99}\%$).
 See the text for details.
-We further explore energy-dependent polarization properties of the selected sources that confirm the presence of significant polarized emission in $2-8$ keV energy range of IXPE.
-For each source except 4U $1957+115$, IGR J17091362417091362417091-362417091 - 3624 and GX $339-4$, we divide $2-8$ keV energy range into six linearly spaced energy bins and estimate polarization using the model-independent approach (see also Svoboda et al.
+We further explore energy-dependent polarization properties of the selected sources that confirm the presence of significant polarized emission in 282-8 keV energy range of IXPE.
+For each source except 4U $1957+115$, IGR J17091362417091-3624 and GX 3394339-4, we divide 282-8 keV energy range into six linearly spaced energy bins and estimate polarization using the model-independent approach (see also Svoboda et al.
 2024a; Veledina et al.
 2023).
-For 4U $1957+115$, IGR J17091362417091362417091-362417091 - 3624 and GX $339-4$, only four energy bins are considered for the energy-dependent study to ensure relatively improved photon statistics per energy bin (see also Marra et al.
+For 4U $1957+115$, IGR J17091362417091-3624 and GX 3394339-4, only four energy bins are considered for the energy-dependent study to ensure relatively improved photon statistics per energy bin (see also Marra et al.
 2024).
@@
 (2024b).
-In doing so, we consider PD in $2-8$ keV energy range (see Table 5) as the null hypothesis against which the significance of energy variations in PD is computed for each epoch of a given source.
+In doing so, we consider PD in 282-8 keV energy range (see Table 5) as the null hypothesis against which the significance of energy variations in PD is computed for each epoch of a given source.
 With this, following Majumder et al.
-(2024b), we calculate the probability of rejecting the null hypothesis using $(N-1)$ degrees of freedom, taking into account the variation of PD over $N$ energy bins for the respective cases.
+(2024b), we calculate the probability of rejecting the null hypothesis using (N1)(N-1) degrees of freedom, taking into account the variation of PD over $N$ energy bins for the respective cases.
 This essentially indicates the confidence level at which the PD variation is significant over different energy bins.
 The obtained significance levels of the energy variation of PD are mentioned in Table 5 for the respective cases.
-We observe that this significance varies from $\lesssim 1\sigma$ to $>4\sigma$ across different epochs of the sources in various spectral states.
+We observe that this significance varies from $\lesssim 1\sigma$ to $4\sigma$ across different epochs of the sources in various spectral states.
 In Fig.
@@
 The histograms in colors represent the $\rm MDP_{99}\%$ associated with the measurements in the selected energy bins for the corresponding epochs.
-Note that, for a given source, when multi-epoch observations are available, we present the results for two epochs of different spectral states only, except LMC X--3.
-Interestingly, we observe that Cyg X11-1- 1 and Swift J1727.816131613-1613- 1613 show similar energy variation of PD within $\sim 2-6\%$ in different spectral states (LHS, HSS, HIMS and SIMS).
-Moreover, for Cyg X11-1- 1, the variation in PD with energy remains most significant at a confidence level exceeding $4\sigma$ in HSS, whereas it remains $2.8\sigma$ for Swift J1727.816131727.816131727.8-16131727.8 - 1613 in HIMS, when all observational epochs are considered (see Fig.
+Note that, for a given source, when multi-epoch observations are available, we present the results for two epochs of different spectral states only, except LMC X-3.
+Interestingly, we observe that Cyg X1-1 and Swift J1727.81613-1613 show similar energy variation of PD within $\sim 2-6\%$ in different spectral states (LHS, HSS, HIMS and SIMS).
+Moreover, for Cyg X1-1, the variation in PD with energy remains most significant at a confidence level exceeding $4\sigma$ in HSS, whereas it remains $2.8\sigma$ for Swift J1727.816131727.8-1613 in HIMS, when all observational epochs are considered (see Fig.
 5).
-Note that, PD $<\rm MDP_{99}\%$ at higher energy bins in a few epochs of several sources, indicates a null-detection.
-Interestingly, unlike Cyg X11-1- 1 and Swift J1727.816131727.816131727.8-16131727.8 - 1613, a weak energy dependency of PD is noticed for IGR $17091-3624$ in the LHS, for which PD corresponding to two out of four energy bins remain below the $\rm MDP_{99}\%$ (see Fig.
+Note that, PD $\rm MDP_{99}\%$ at higher energy bins in a few epochs of several sources, indicates a null-detection.
+Interestingly, unlike Cyg X1-1 and Swift J1727.816131727.8-1613, a weak energy dependency of PD is noticed for IGR 17091362417091-3624 in the LHS, for which PD corresponding to two out of four energy bins remain below the $\rm MDP_{99}\%$ (see Fig.
 5).
-We find a significant increase in PD with energy up to $\sim 11\%$ in HSS ($>4\sigma$) and SPL ($>4\sigma$) state of 4U 16304747-47- 47 (see Fig.
+We find a significant increase in PD with energy up to $\sim 11\%$ in HSS ($4\sigma$) and SPL ($4\sigma$) state of 4U 163047-47 (see Fig.
 5).
-Similarly, distinct variations of PD over several energy bins are observed for Cyg X33-3- 3 during LHS ($>4\sigma$) and SIMS ($>4\sigma$).
-LMC X33-3- 3 shows constant polarization over the initial three energy bins with a sudden jump in PD up to $\sim 7.8\%$ at $\sim 5.4$ keV in both epochs (see Fig.
+Similarly, distinct variations of PD over several energy bins are observed for Cyg X3-3 during LHS ($4\sigma$) and SIMS ($4\sigma$).
+LMC X3-3 shows constant polarization over the initial three energy bins with a sudden jump in PD up to $\sim 7.8\%$ at $\sim 5.4$ keV in both epochs (see Fig.
 5).
-Further, we only observe a marginal variation ($1.7\sigma$) of PD in 4U 1957$+115$ up to $\sim 3\%$.
-Interestingly, significant ($3.3\sigma$) energy variation of polarization degree is observed for GX $339-4$.
-Although, PD $<\rm MDP_{99}\%$ is obtained in all four energy bins (see Fig.
+Further, we only observe a marginal variation ($1.7\sigma$) of PD in 4U 1957+115+115 up to $\sim 3\%$.
+Interestingly, significant ($3.3\sigma$) energy variation of polarization degree is observed for GX 3394339-4.
+Although, PD $\rm MDP_{99}\%$ is obtained in all four energy bins (see Fig.
 5).
-Overall, we state that a significant ($>3\sigma$) variation of polarization degree with energy is observed only for Cyg X11-1- 1, 4U 16304747-47- 47 and Cyg X33-3- 3, whereas a moderate ($2\sigma$ to $3\sigma$) variation is seen for Swift J1727.816131613-1613- 1613 and LMC X33-3- 3.
-However, Veledina et al.
-2024a, b reported an energy-independent variation of PD across 12 energy bins for Cyg X--3.
-We mention that the strong ($>4\sigma$) energy dependence of PD observed in our analysis possibly resulted due to relatively small uncertainties obtained in PD estimates over the energy bins and a significant drop in PD around the unpolarized iron line complex at $\sim 6.4$ keV.
-It is worth mentioning that the energy variation of PA for all sources under consideration remains statistically insignificant ($<2\sigma$).
-However, we note that PA drops from $-30^{\circ}$ to $-60^{\circ}$ at $\sim 6$ keV in 4U $1957+115$, which is marginally significant at $1.6\sigma$ confidence level.
+Overall, we state that a significant ($3\sigma$) variation of polarization degree with energy is observed only for Cyg X1-1, 4U 163047-47 and Cyg X3-3, whereas a moderate ($2\sigma$ to $3\sigma$) variation is seen for Swift J1727.81613-1613 and LMC X3-3.
+It is worth mentioning that the energy variation of PA for all sources under consideration remains statistically insignificant ($2\sigma$).
+However, we note that PA drops from $30^{\circ}$ to $60^{\circ}$ at $\sim 6$ keV in 4U $1957+115$, which is marginally significant at $1.6\sigma$ confidence level.
 4.5 Spectro-polarimetric Modeling
-We explore the model-dependent spectro-polarimetric properties of the sources from the simultaneous fitting of the I, Q and U Stokes spectra of IXPE in $2-8$ keV energy range.
-Due to the calibration issues Krawczynski et al.
-(2022); Veledina et al.
-(2024a); Steiner et al.
-(2024), only DU1 data is considered for the spectro-polarimetric study of Cyg X11-1- 1, Cyg X33-3- 3 and Swift J1727.816131613-1613- 1613, whereas data combining all the DUs are used for the remaining sources.
-For simplicity, we first adopt the model combination M4: Tbabs×$\times$×polconst×$\times$(diskbb + powerlaw) for the spectro-polarimetric fitting.
-Here, polconst represents a constant polarization model, where the degree and angle of polarization are treated as the model parameters.
-The best-fit results yield $\chi_{\rm red}^{2}$ ranging from $0.91-1.79$ across all selected sources, with PD varying between $\sim 1.16-13.06\%$ and PA spanning from $-89.13^{\circ}$ to $85.15^{\circ}$.
-Note that the spectro-polarimetric modeling of Swift J1727.816131727.816131727.8-16131727.8 - 1613 does not constrain the hydrogen column density, likely due to low-energy threshold of IXPE ($\sim 2$ keV; see Veledina et al., 2023; Ingram et al., 2024).
+[…]
 Next, we investigate the energy-dependent polarization properties using model-dependent approach for all the sources that demonstrate the signature of X-ray polarization.
-In doing so, we replace the polconst component with the energy-dependent polarization model polpow [^15] https://heasarc.gsfc.nasa.gov/xanadu/xspec/manual/node218.html.
-Empirically, polpow assumes a power-law variation in the polarization parameters of the form PD($E$) $=A_{\rm norm}\times E^{-A_{\rm index}}$ (in fraction) and PA($E$) $=\psi_{\rm norm}\times E^{-\psi_{\rm index}}$ (in ) while modeling the Q and U Stokes spectra.
+In doing so, we replace the polconst component with the energy-dependent polarization model polpow.
+Empirically, polpow assumes a power-law variation in the polarization parameters of the form PD($E$) $A_{\rm norm}\times E^{-A_{\rm index}}$ (in fraction) and PA($E$) $\psi_{\rm norm}\times E^{-\psi_{\rm index}}$ (in ) while modeling the Q and U Stokes spectra.
 We find that the best-fitted $\psi_{\rm index}$ remains consistent with zero within its $1\sigma$ errors and hence, we freeze it to zero during the spectral fitting of all observations.
 Accordingly, we obtain the best spectral fit for all sources with $\chi_{\rm red}^{2}\sim 0.97-1.46$.
-Furthermore, integrating PD($E$) in $2-8$ keV energy range using the best-fitted model parameters ($A_{\rm index}$ and $A_{\rm norm}$), we obtain PD as $3.23-14.33\%$.
-Moreover, with $\psi_{\rm index}$ frozen to zero, the polarization angles are obtained as the best-fitted $\psi_{\rm norm}$ values that vary from $-88.92^{\circ}$ to $85.11^{\circ}$.
-Note that, in a few observations of Cyg X11-1- 1, 4U 1957$+115$ and GX $339-4$ (see Table 6), the data is unable to constrain the parameters of polpow component even at $1\sigma$ level, and hence, we refrain from the modeling with polpow for these observations.
+Furthermore, integrating PD($E$) in 282-8 keV energy range using the best-fitted model parameters ($A_{\rm index}$ and $A_{\rm norm}$), we obtain PD as $3.23-14.33\%$.
+Moreover, with $\psi_{\rm index}$ frozen to zero, the polarization angles are obtained as the best-fitted $\psi_{\rm norm}$ values that vary from $88.92^{\circ}$ to $85.11^{\circ}$.
+Note that, in a few observations of Cyg X1-1, 4U 1957+115+115 and GX 3394339-4 (see Table 6), the data is unable to constrain the parameters of polpow component even at $1\sigma$ level, and hence, we refrain from the modeling with polpow for these observations.
 In Table 6, we present the best-fitted model parameters obtained from the spectro-polarimetric study of all sources.
-We find that the model-dependent polarimetric results are in agreement with the findings of the model-independent study except Cyg X33-3- 3, confirming the robustness of the polarization measurements for the individual sources.
-However, for Cyg X--3, we obtain significantly low PD compared to the results of the model-independent study, likely due to multiple low-energy unpolarized emission lines including strong iron complex around $\sim 6.4$ keV (see Veledina et al.
-2024a), dominating the Stokes spectra.
+We find that the model-dependent polarimetric results are in agreement with the findings of the model-independent study except Cyg X3-3, confirming the robustness of the polarization measurements for the individual sources.
+However, for Cyg X-3, we obtain significantly low PD compared to the results of the model-independent study, likely due to multiple low-energy unpolarized emission lines including strong iron complex around $\sim 6.4$ keV Veledina et al.
+(2024a), dominating Stokes spectra.
 4.6 Spectro-polarimetric Correlation
-Figure 6: Variation of PD ($2-8$ keV), obtained from model-independent analysis, as a function of the disc-to-Comptonized flux ratio ($F_{\rm ratio}$) derived from wide-band spectral modelling.
-The color bar represents the covering fraction ($cov_{\rm frac}$), while each data point is labeled with the corresponding spectral state and observational epoch.
+Figure 6: Variation of PD (282-8 keV), obtained from model-independent analysis, as a function of the disc-to-Comptonized flux ratio ($F_{\rm ratio}$) derived in 11001-100 keV energy range from the wide-band spectral modeling.
+The color bar represents the covering fraction ($cov_{\rm frac}$) obtained from thcomp model, while each data point is labeled with the corresponding spectral state and observational epoch.
 Data points within the shaded region (light blue) are used to compute the Pearson correlation coefficient.
 Downward arrows indicate null-detections of polarization for the respective sources.
-See the text for details.
 We examine the spectro-polarimetric correlation properties for each source based on the results obtained from wide-band spectral modeling using model M1 (see §4.2) and the polarimetric measurements (see §4.3).
-Specifically, we explore the correlation between the observed PD obtained from model-independent study, and best-fitted spectral parameters, namely the ratio of disc to Comptonized spectral flux ($F_{\rm ratio}$) in $1-100$ keV energy range, and the covering fraction ($cov_{\rm frac}$).
+Specifically, we explore the correlation between the observed PD obtained from model-independent study, and best-fitted spectral parameters, namely the ratio of disc to Comptonized spectral flux ($F_{\rm ratio}$) in 11001-100 keV energy range, and the covering fraction ($cov_{\rm frac}$) obtained from the thcomp model.
 In Fig.
@@
 […]
-Furthermore, we find a strong correlation (anti-correlation) between PD and $cov_{\rm frac}$ ($F_{\rm ratio}$) with $\rho\sim 0.9$ ($-0.9$) for Swift J1727.816131727.816131727.8-16131727.8 - 1613.
+Furthermore, we find a strong correlation (anti-correlation) between PD and $cov_{\rm frac}$ ($F_{\rm ratio}$) with $\rho\sim 0.9$ (0.9-0.9) for Swift J1727.816131727.8-1613.
 In particular, a sharp drop in PD is noticed from $\sim 4.7\%$ (HIMS) to $\sim 0.4\%$ (HSS) (see also Svoboda et al.
 2024b) during which $F_{\rm ratio}$ increases from $\sim 0.7$ to $\sim 66$ and $cov_{\rm frac}$ noticeably reduces down to $13\%$ from a maximum value of $82\%$.
-A similar trend is observed for Cyg X11-1- 1 with $\rho=0.9$ ($-0.9$) for the respective positive (negative) correlations between PD and $cov_{\rm frac}$ ($F_{\rm ratio}$).
-As before, a low (high) $F_{\rm ratio}\sim 0.3$ ($11.5$) and high (low) $cov_{\rm frac}\sim 61-77\%$ ($0.5-31\%$) are observed in the LHS (HSS) of Cyg X11-1- 1 with PD $\sim 3.7-4.8\%$ ($1.4-2.1\%$) (see Fig.
+A similar trend is observed for Cyg X1-1 with $\rho=0.9$ (0.9-0.9) for the respective positive (negative) correlations between PD and $cov_{\rm frac}$ ($F_{\rm ratio}$).
+As before, a low (high) $F_{\rm ratio}\sim 0.3$ ($11.5$) and high (low) $cov_{\rm frac}\sim 61-77\%$ ($0.5-31\%$) are observed in the LHS (HSS) of Cyg X1-1 with PD $\sim 3.7-4.8\%$ ($1.4-2.1\%$) (see Fig.
 6).
-Moreover, we also find a similar behavior in the LHS of IGR J17091362417091362417091-362417091 - 3624, for which a very high PD ($\sim 9\%$) is seen with $cov_{\rm frac}\sim 87\%$ and $F_{\rm ratio}\sim 0.2$.
-These findings suggest that a significant contribution in the polarization degree comes from the Comptonized emission for Swift J1727.816131727.816131727.8-16131727.8 - 1613, Cyg X11-1- 1 and IGR J17091362417091362417091-362417091 - 3624 in the HIMS and/or LHS.
-Furthermore, we notice a low polarization degree (PD $\lesssim 2\%$) for all the sources except 4U 16304747-47- 47 and LMC X33-3- 3 in the HSS/SIMS with $F_{\rm ratio}\gtrsim 2$ and $cov_{\rm frac}\lesssim 30\%$.
-Interestingly, 4U 16304747-47- 47 being the source of extreme soft nature ($F_{\rm ratio}\sim 52$, $cov_{\rm frac}\sim 5\%$) manifests a high polarization degree of $\sim 8.3\%$ in the HSS, whereas it reduces to $6.8\%$ in the SPL state with $F_{\rm ratio}\sim 3.6$ and $cov_{\rm frac}\sim 38\%$.
-We observe that 4U $1957+115$ (GX $339-4$) shows PD $\sim 1.2\%$ ($1.9\%$) with $cov_{\rm frac}\sim 2.7\%$ ($34\%$) and $F_{\rm ratio}\sim 18$ ($2.4$) in the HSS (SIMS).
-Moreover, LMC X33-3- 3 exhibits moderate polarization with PD in the range of $2.4-3\%$ for $F_{\rm ratio}\gtrsim 4$ and $cov_{\rm frac}\sim 0.5\%$ during the HSS.
+Moreover, we also find a similar behavior in the LHS of IGR J17091362417091-3624, for which a very high PD ($\sim 9\%$) is seen with $cov_{\rm frac}\sim 87\%$ and $F_{\rm ratio}\sim 0.2$.
+These findings suggest that a significant contribution in the polarization degree comes from the Comptonized emission for Swift J1727.816131727.8-1613, Cyg X1-1 and IGR J17091362417091-3624 in the HIMS and/or LHS.
+Furthermore, we notice a low polarization degree (PD $\lesssim 2\%$) for all the sources except 4U 163047-47 and LMC X3-3 in the HSS/SIMS with $F_{\rm ratio}\gtrsim 2$ and $cov_{\rm frac}\lesssim 30\%$.
+Interestingly, 4U 163047-47 being the source of extreme soft nature ($F_{\rm ratio}\sim 52$, $cov_{\rm frac}\sim 5\%$) manifests a high polarization degree of $\sim 8.3\%$ in the HSS, whereas it reduces to $6.8\%$ in the SPL state with $F_{\rm ratio}\sim 3.6$ and $cov_{\rm frac}\sim 38\%$.
+We observe that 4U $1957+115$ (GX 3394339-4) shows PD $\sim 1.2\%$ ($1.9\%$) with $cov_{\rm frac}\sim 2.7\%$ ($34\%$) and $F_{\rm ratio}\sim 18$ ($2.4$) in the HSS (SIMS).
+Moreover, LMC X3-3 exhibits moderate polarization with PD in the range of $2.4-3\%$ for $F_{\rm ratio}\gtrsim 4$ and $cov_{\rm frac}\sim 0.5\%$ during the HSS.
 5 Discussion
 In this study, we present a comprehensive spectro-polarimetric investigation of eleven BH-XRBs using quasi-simultaneous observations from IXPE, NICER, AstroSat and NuSTAR.
-IXPE data are analyzed to examine the polarization properties in $2-8$ keV energy range, while combined NICER, AstroSat and NuSTAR observations provide insights into the spectro-temporal behavior across the wide energy range ($0.5-100$ keV).
+IXPE data are analyzed to examine the polarization properties in 282-8 keV energy range, while combined NICER, AstroSat and NuSTAR observations provide insights into the spectro-temporal behavior across the wide energy range (0.51000.5-100 keV).
 Subsequently, the spectro-polarimetric findings are used to investigate the accretion-ejection dynamics and geometry surrounding the sources.
 5.1 Spectro-temporal Properties and Spectral States
-Spectro-temporal studies using NICER, NuSTAR and AstroSat observations in $0.5-100$ keV energy range impart distinct variability properties in various spectral states of the sources (see §4.1 and §4.2).
-The obtained results are in line with the spectro-temporal characteristics generally seen in the respective spectral states of BH-XRBs (Belloni et al., 2005; Remillard & McClintock, 2006; Kalemci et al., 2006; McClintock et al., 2009; Nandi et al., 2012; Radhika & Nandi, 2014; Radhika et al., 2016; Sreehari et al., 2019b; Connors et al., 2019; Méndez et al., 2022; Athulya et al., 2022; Aneesha et al., 2024; Shui et al., 2024; Li et al., 2025).
+Spectro-temporal studies using NICER, NuSTAR and AstroSat observations in 0.51000.5-100 keV energy range impart distinct variability properties in various spectral states of the sources (see §4.1 and §4.2).
+The obtained results are in line with the spectro-temporal characteristics generally seen in the respective spectral states of BH-XRBs (Belloni et al., 2005; McClintock et al., 2009; Nandi et al., 2012; Sreehari et al., 2019b; Méndez et al., 2022; Aneesha et al., 2024).
 These findings reinforce the presence of distinct spectral states as observed during the multi-mission campaigns of the sources.
@@
 2).
-In addition, we find the presence of a strong type-C QPO feature at $\sim 1.4$ Hz (E1), which evolves to $\sim 6.7$ Hz (E2-E5), during the HIMS and SIMS of Swift J1727.816131613-1613- 1613 over a wide energy range of $0.5-100$ keV with NICER, NuSTAR and AstroSat observations (see Fig.
+In addition, we find the presence of a strong type-C QPO feature at $\sim 1.4$ Hz (E1), which evolves to $\sim 6.7$ Hz (E2-E5), during the HIMS and SIMS of Swift J1727.81613-1613 over a wide energy range of 0.51000.5-100 keV with NICER, NuSTAR and AstroSat observations (see Fig.
 2).
 The origin of this QPO is attributed to a strong Comptonizing corona surrounding the central source (Chakrabarti & Titarchuk, 1995; Chakrabarti et al., 2008; Nandi et al., 2024); however, based on the correlation between the evolution of QPO frequency and quasi-simultaneous radio jet ejections, Liao et al.
-(2024) suggest a jet-like, vertically elongated corona for Swift J1727.816131727.816131727.8-16131727.8 - 1613.
-A similar type-C QPO at $\sim 0.2$ Hz is detected in LHS of IGR J17091362417091362417091-362417091 - 3624 (see Fig.
+(2024) suggest a jet-like, vertically elongated corona for Swift J1727.816131727.8-1613.
+A similar type-C QPO at $\sim 0.2$ Hz is detected in LHS of IGR J17091362417091-3624 (see Fig.
 2).
-Moreover, we also observe a type-B QPO in the disc-dominated SIMS of GX $339-4$ in the $0.5-80$ keV energy range from NICER and AstroSat data.
-Similar type-B QPOs have been observed during $2024$ and earlier outbursts of GX $339-4$, indicating the presence of a compact corona located very close to the black hole, accompanied by dominant disc emission (Nandi et al., 2012; Aneesha et al., 2024).
-Admittedly, the wide-band energy spectral analysis ($0.7-70$ keV) reveals distinct spectral states of each source.
+Moreover, we also observe a type-B QPO in the disc-dominated SIMS of GX 3394339-4 in the 0.5800.5-80 keV energy range from NICER and AstroSat data.
+Similar type-B QPOs have been observed earlier in GX 3394339-4, indicating the presence of a compact corona located very close to the black hole, accompanied by dominant disc emission (Nandi et al., 2012; Aneesha et al., 2024).
+The wide-band (0.7600.7-60 keV) energy spectral analysis reveals distinct spectral states of each source.
 The observed spectra of all sources are well described by a standard accretion disc component (diskbb) combined with thermal Comptonization of soft photons by a spherical corona (thcomp) (see Table 3 and Fig.
@@
 For example, Svoboda et al.
-2024b found a steeper $\Gamma_{\rm th}\sim 4.9$ and low $cov_{\rm frac}\sim 20\%$ in the HSS of Swift J1727.816131727.816131727.8-16131727.8 - 1613, which agrees well with our findings of $\Gamma_{\rm th}\sim 4.6-5.1$ and $cov_{\rm frac}\sim 17\%$.
-In addition, the LHS results for Cyg X--1 (E1) are consistent with those of Krawczynski et al.
+2024b found a steeper $\Gamma_{\rm th}\sim 4.9$ and low $cov_{\rm frac}\sim 20\%$ in the HSS of Swift J1727.816131727.8-1613, which agrees well with our findings of $\Gamma_{\rm th}\sim 4.6-5.1$ and $cov_{\rm frac}\sim 17\%$.
+In addition, the LHS results for Cyg X-1 (E1) are consistent with those of Krawczynski et al.
 (2022), except for their measurement of a high electron temperature, which is likely because of the differences in spectral energy coverage and/or the choice of model components.
@@
 This evidently indicates the degeneracy in determining the corona geometry (see §4.2).
-Furthermore, strong reflection signatures detected in the soft state observations (HSS and SIMS) of Cyg X11-1- 1, Swift J1727.816131727.816131727.8-16131727.8 - 1613, and GX $339-4$ allow the detailed investigation of alternative coronal geometries using various reflection models (see §4.2).
+Furthermore, strong reflection signatures detected in the soft state observations (HSS and SIMS) of Cyg X1-1, Swift J1727.816131727.8-1613, and GX 3394339-4 allow the detailed investigation of alternative coronal geometries using various reflection models (see §4.2).
 We also observe that the reflected emission, either from a lamppost corona, modeled with relxilllp, or from a spherical corona as described in relxillCp, can be reproduced satisfactorily considering a highly ionized accretion disc ($\log\xi\sim 4$).
@@
 5.2 Model Prescriptions and Limitations of X-ray Polarization in BH-XRBs
-X-ray polarization measurements with IXPE reveal significant polarized emission across several spectral states of the sources in $2-8$ keV energy range, except for LMC X--1 (see §4.3 and Podgorný et al.
-2023), Swift J151857.057214151857.057214151857.0-57214151857.0 - 57214 (see §4.3 and Ling et al.
-2024) and MAXI J174429417442941744-2941744 - 294 (see §4.3).
-We find that Cyg X--1 exhibits PD within $3-4.8\%$ in LHS, which decreases to $1.4-2.5\%$ in HSS.
-A similar noticeable reduction in the PD from $4.7\%$ (HIMS) to $0.4\%$ (HSS) is also observed for the exceptionally bright transient Swift J1727.8--1613.
+X-ray polarization measurements with IXPE reveal significant polarized emission across several spectral states of the sources in 282-8 keV energy range, except for LMC X-1 (see §4.3 and Podgorný et al.
+2023), Swift J151857.057214151857.0-57214 (see §4.3 and Ling et al.
+2024) and MAXI J17442941744-294 (see §4.3).
+We find that Cyg X-1 exhibits PD within $3-4.8\%$ in LHS, which decreases to $1.4-2.5\%$ in HSS.
+A similar noticeable reduction in the PD from $4.7\%$ (HIMS) to $0.4\%$ (HSS) is also observed for the exceptionally bright transient Swift J1727.8-1613.
 The most recent IXPE observation of this source confirms the recovery of PD $\sim 3.2\%$ in the faint LHS.
-We mention that these results are in agreement with the previously reported polarimetric findings of Cyg X--1 (Krawczynski et al., 2022; Jana & Chang, 2024; Steiner et al., 2024) and Swift J1727.8--1613 (Veledina et al., 2023; Ingram et al., 2024; Svoboda et al., 2024b; Podgorný et al., 2024).
-Intriguingly, an exceptionally high PD of $\sim 8.3\%$ is observed in the HSS of 4U 16304747-47- 47, which drops to $\sim 6.8\%$ during the SPL state exhibiting strong ($>4\sigma$) energy dependency (see Table 5 and Fig.
+We mention that these results are in agreement with the previously reported polarimetric findings of Cyg X-1 (Krawczynski et al., 2022; Jana & Chang, 2024; Steiner et al., 2024) and Swift J1727.8-1613 (Veledina et al., 2023; Ingram et al., 2024; Svoboda et al., 2024b; Podgorný et al., 2024).
+Intriguingly, an exceptionally high PD of $\sim 8.3\%$ is observed in the HSS of 4U 163047-47, which drops to $\sim 6.8\%$ during the SPL state exhibiting strong ($4\sigma$) energy dependency (see Table 5 and Fig.
 5; also Kushwaha et al.
-2023b; Rodriguez Cavero et al.
+2023; Rodriguez Cavero et al.
 2023; Rawat et al.
@@
 2024).
-In contrast, low PD values are observed in several sources: LMC X33-3- 3 ($\sim 2.4-3\%$), 4U 1957$+115$ ($\sim 1.9\%$) and GX $339-4$ ($\sim 1.2\%$), being consistent with the previous measurements (Majumder et al., 2024a; Svoboda et al., 2024a; Garg et al., 2024; Marra et al., 2024; Mastroserio et al., 2025).
-In the present work, we report polarization detections analyzing the most recent IXPE observations of Cyg X--3 (PD $\sim 21.4\%$, SIMS), LMC X--3 (PD $\sim 2.4\%$, HSS) and IGR J17091--3624 (PD $\sim 9\%$, LHS), apart from the earlier measurements of these sources.
-In addition, for the first time to the best of our knowledge, we report the null-detection of X-ray polarization in LMC X33-3- 3 (E3) and a newly discovered BH-XRB candidate MAXI J174429417442941744-2941744 - 294.
+In contrast, low PD values are observed in several sources: LMC X3-3 ($\sim 2.4-3\%$), 4U 1957+115+115 ($\sim 1.9\%$) and GX 3394339-4 ($\sim 1.2\%$), being consistent with the previous measurements (Majumder et al., 2024a; Svoboda et al., 2024a; Garg et al., 2024; Marra et al., 2024; Mastroserio et al., 2025).
+In the present work, we report polarization detections analyzing the most recent IXPE observations of Cyg X-3 (PD $\sim 21.4\%$, SIMS), LMC X-3 (PD $\sim 2.4\%$, HSS) and IGR J17091-3624 (PD $\sim 9\%$, LHS), apart from the earlier measurements of these sources.
+In addition, for the first time to the best of our knowledge, we report the null-detection of X-ray polarization in LMC X3-3 (E3) and a newly discovered BH-XRB candidate MAXI J17442941744-294.
+We note that, subsequent to our report, null-detection of polarization for MAXI J17442941744-294 is also confirmed by Marra et al.
+(2025).
 Meanwhile, several models have been proposed to account for the observed polarization in BH-XRBs.
-For instance, remarkably high PD ($\sim 8.3\%$) of 4U 1630--47 is attributed to the emission from a partially ionized thin/slim accretion disc viewed at inclination $i\gtrsim 60^{\circ}$ (Ratheesh et al., 2024).
-Similar explanation is also suggested for LMC X33-3- 3 in the HSS that exhibits PD $\sim 3\%$ at $i\sim 70^{\circ}$ (Majumder et al., 2024a; Svoboda et al., 2024a).
-Further, scattering of photons in strong disc winds of different opening angles ($\alpha_{\rm w}\sim 30^{\circ}-40^{\circ}$) has also been suggested as another possible mechanism to explain high PD values of 4U 1630--47 (Kushwaha et al., 2023b; Rawat et al., 2023b; Nitindala et al., 2025).
+For instance, remarkably high PD ($\sim 8.3\%$) of 4U 1630-47 is attributed to the emission from a partially ionized thin/slim accretion disc viewed at inclination $i\gtrsim 60^{\circ}$ (Ratheesh et al., 2024).
+Similar explanation is also suggested for LMC X3-3 in the HSS that exhibits PD $\sim 3\%$ at $i\sim 70^{\circ}$ (Majumder et al., 2024a; Svoboda et al., 2024a).
+Further, scattering of photons in strong disc winds of different opening angles ($\alpha_{\rm w}\sim 30^{\circ}-40^{\circ}$) has also been suggested as another possible mechanism to explain high PD values of 4U 1630-47 (Kushwaha et al., 2023; Rawat et al., 2023b; Nitindala et al., 2025).
+However, Tomaru et al.
+(2024) demonstrated that electron scattering in an optically thin wind produces polarization in a direction opposite to that obtained from the intrinsic emission from an optically thick disc.
+As a result, the wind component is expected to depolarize the total emission, including the intrinsic disc emission.
+Therefore, the presence of disc-wind from the optically thick disc appears to be insufficient to account for the observed high PD in 4U 1630-47.
 On the other hand, in high-spin black hole systems, strong relativistic effects cause intense gravitational lensing that bends emitted photons back onto the disc.
 This enhances the reflection features and significantly affects the observed polarization (see §4.2, Schnittman & Krolik, 2009).
-Such an effect appears consistent with the polarimetric observations (PD $\sim 1.2-2.5\%$) in the soft state as reported for rapidly spinning systems like Cyg X11-1- 1 ($a_{*}>0.998$) and 4U $1957+115$ ($a_{*}\sim 0.988$) (Marra et al., 2024; Steiner et al., 2024).
-Furthermore, it has been suggested that $\sim 4\%$ (Cyg X11-1- 1, LHS) and $\sim 9\%$ (IGR J17091362417091362417091-362417091 - 3624, LHS) polarization can be obtained from a wedge-shaped corona with mildly relativistic outflowing plasma having inclination in the range $30^{\circ}\lesssim i\lesssim 60^{\circ}$ (Poutanen et al., 2023; Ewing et al., 2025).
+Such an effect appears consistent with the polarimetric observations (PD $\sim 1.2-2.5\%$) in the soft state as reported for rapidly spinning systems like Cyg X1-1 ($a_{*}>0.998$) and 4U $1957+115$ ($a_{*}\sim 0.988$) (Marra et al., 2024; Steiner et al., 2024).
+Furthermore, it has been suggested that $\sim 4\%$ (Cyg X1-1, LHS) and $\sim 9\%$ (IGR J17091362417091-3624, LHS) polarization can be obtained from a wedge-shaped corona with mildly relativistic outflowing plasma model of Beloborodov (1998) having inclination in the range $30^{\circ}\lesssim i\lesssim 60^{\circ}$ (Poutanen et al., 2023; Ewing et al., 2025).
 A similar model of bulk Comptonization involving a mildly relativistic outflow predicts an even higher polarization degree up to $\lesssim 10\%$ (Dexter & Begelman, 2024).
-Notably, observed $\sim 4.7\%$ PD at $i\sim 40^{\circ}$ in HIMS of Swift J1727.816131727.816131727.8-16131727.8 - 1613 also seems to be consistent with the predictions from a wedge-shaped corona with outflowing plasma velocity $\lesssim 0.2c$ (Poutanen et al., 2023).
-However, QPOs are detected in Swift J1727.816131727.816131727.8-16131727.8 - 1613 ($0.5-100$ keV) and IGR J17091362417091362417091-362417091 - 3624 ($3-79$ keV) during observations simultaneous with IXPE, although their physical origin remains unclear within the framework of the outflowing corona geometry.
+Notably, observed $\sim 4.7\%$ PD at $i\sim 40^{\circ}$ in HIMS of Swift J1727.816131727.8-1613 also seems to be consistent with the predictions from a wedge-shaped corona with outflowing plasma velocity $\lesssim 0.2c$ (Poutanen et al., 2023).
+However, QPOs are detected in Swift J1727.816131727.8-1613 (0.51000.5-100 keV) and IGR J17091362417091-3624 (3793-79 keV) during observations simultaneous with IXPE, although their physical origin remains unclear within the framework of the outflowing corona geometry.
 On the other hand, Kumar (2024) predicted PD $\sim 0.5-4\%$ assuming a simple static spherical corona for $30^{\circ}\lesssim i\lesssim 60^{\circ}$.
-This appears broadly consistent with the low PD of $\sim 1.2\%$ observed at $i\sim 30^{\circ}$ in the SIMS of GX $339-4$.
+This appears broadly consistent with the low PD of $\sim 1.2\%$ observed at $i\sim 30^{\circ}$ in the SIMS of GX 3394339-4.
 All the above interpretations of the polarimetric findings point toward a persistent degeneracy in understanding the different accretion-ejection dynamics.
@@
 5.3 Accretion-Ejection Scenarios in BH-XRBs
-Figure 7: Schematic representation of the possible disc-corona-jet configurations for seven BH-XRB sources based on the framework of two-component accretion flow scenario.
-The Keplerian and sub-Keplerian components, as indicated in panel (b), are also present in the other two configurations.
-Colored vertical arrows in the corona regions of panels (a) and (c) indicate the direction of polarization angle for the respective sources.
-In panels (a) and (c), the jet configuration applies only to the sources in distinct spectral states, which are marked by their respective serial numbers.
+Figure 7: Schematic representation of the possible disc–corona–jet configurations for seven BH-XRB sources within the two-component accretion flow framework.
+The Keplerian and sub-Keplerian components shown in panel (b) are also present in the other three configurations.
+Colored vertical arrows in the corona regions of panels (a) and (c) denote the polarization angle for the respective sources.
+In panels (a) and (d), the jet/radio emission configuration applies only to sources in distinct spectral states, indicated by their serial numbers.
 See the text for details.
 Following the discussion of model degeneracies and the need for a unified model framework, we next explore physically motivated accretion-ejection geometries capable of explaining the commonly observed spectro-temporal features, while also offering a promising basis for interpreting the spectro-polarimetric signatures in BH-XRBs.
-Towards this, a plausible disc-corona-jet configuration based on the two-component accretion flow model Chakrabarti & Titarchuk (1995) is illustrated in Fig.
-7 and discussed in detail below.
+Towards this, a plausible disc-corona-jet configuration based on the two-component accretion flow model Chakrabarti & Titarchuk (1995); Chakrabarti (1996) is illustrated in Fig.
+7.
+In this model, the accretion flow around black holes comprise with Keplerian disc flanked between the sub-Keplerian components.
+This sub-Keplerian component generally forms the post-shock corona (PSC, equivalently Comptonizing corona) in the inner region of the disc Chakrabarti & Titarchuk (1995); Molteni et al.
+(1996); Chakrabarti et al.
+(2008); Nandi et al.
+(2012); Das et al.
+(2014).
+Soft photons from the Keplerian disc are up-scattered by hot electrons present in the corona and produce the hard spectral tail seen in BH-XRBs.
+The inner edge of the Keplerian disc smoothly connects to the outer boundary of the coronal region.
 The spectro-polarimetric correlation study (see §4.6 and Fig.
-6) suggests that the polarization signatures (PD $\sim 3-9\%$) in the LHS/HIMS of Cyg X11-1- 1, Swift J1727.816131727.816131727.8-16131727.8 - 1613 and IGR J17091362417091362417091-362417091 - 3624 are primarily governed by the dominant Comptonization process characterized by $F_{\rm ratio}<1$ and $cov_{\rm frac}\sim 35-87\%$.
-Interestingly, the radio jet position angle of Cyg X11-1- 1 and Swift J1727.816131727.816131727.8-16131727.8 - 1613 roughly aligns along the observed PA throughout the respective IXPE campaigns (Miller-Jones et al., 2021; Wood et al., 2024).
+6) suggests that the polarization signatures (PD $\sim 3-9\%$) in the LHS/HIMS of Cyg X1-1, Swift J1727.816131727.8-1613 and IGR J17091362417091-3624 are primarily governed by the dominant Comptonization process characterized by $F_{\rm ratio}<1$ and $cov_{\rm frac}\sim 35-87\%$.
+Interestingly, the radio jet position angle of Cyg X1-1 and Swift J1727.816131727.8-1613 roughly aligns along the observed PA throughout the respective IXPE campaigns (Miller-Jones et al., 2021; Wood et al., 2024).
 This suggests that PA is possibly oriented perpendicular to the disc-corona geometry.
@@
 These pieces of evidence suggest a disc-corona configuration comprising a radial corona located close to the black hole, along with a truncated accretion disc at larger radius during LHS/HIMS (see Fig.
-7a) of Swift J1727.816131727.816131727.8-16131727.8 - 1613 and Cyg X11-1- 1.
-The above configuration potentially explains the QPOs in Swift J1727.816131727.816131727.8-16131727.8 - 1613 (see §4.2 and Fig.
+7a) of Swift J1727.816131727.8-1613 and Cyg X1-1.
+The above configuration potentially explains the QPOs in Swift J1727.816131727.8-1613 (see §4.2 and Fig.
 2c) in terms of the oscillating boundary of the Comptonizing region (see Fig.
@@
 2024).
-It is worth mentioning that the coronal geometry in IGR J17091362417091362417091-362417091 - 3624 remains inconclusive from the polarimetric results due to the unresolved position angle of the radio jet.
-Although the presence of QPO in the LHS of IGR J17091362417091362417091-362417091 - 3624 (see Fig.
+It is worth mentioning that the coronal geometry in IGR J17091362417091-3624 remains inconclusive from the polarimetric results due to the unresolved position angle of the radio jet.
+Although the presence of QPO in LHS of IGR J17091362417091-3624 (see Fig.
 2g) perhaps indicates a similar disc-corona configuration presented in Fig.
 7a.
-However, the absence of QPO in Cyg X11-1- 1 implies that either the coronal boundary fails to satisfy the necessary resonance condition for modulation, or the oscillation is too subtle to be detected as predicted by Majumder et al.
+However, the absence of QPO in Cyg X1-1 implies that either the coronal boundary fails to satisfy the necessary resonance condition for modulation, or oscillation is too subtle to be detected as predicted by Majumder et al.
 (2025a).
-For 4U $1630-47$, an absorption feature around $\sim 7$ keV is observed, suggesting the presence of strong disc winds in the HSS ($F_{\rm ratio}\sim 52$ and $cov_{\rm frac}\sim 5\%$) spectra (see §4.2).
+For 4U 1630471630-47, an absorption feature around $\sim 7$ keV is observed, suggesting the presence of strong disc winds in the HSS ($F_{\rm ratio}\sim 52$ and $cov_{\rm frac}\sim 5\%$) spectra (see §4.2).
 Accordingly, a disc-dominated accretion configuration along with a weak corona, accompanied by strong disc winds with varying opening angles (see also Nitindala et al.
@@
 7b).
-For GX $339-4$, a low PD ($\sim 1.2\%$) is observed within $3-8$ keV energy range during SIMS ($cov_{\rm frac}\sim 34\%$ and $F_{\rm ratio}\sim 2.4$), while null-detection of PD is seen in the HSS (Table 5).
+For GX 3394339-4, a low PD ($\sim 1.2\%$) is observed within 383-8 keV energy range during SIMS ($cov_{\rm frac}\sim 34\%$ and $F_{\rm ratio}\sim 2.4$), while null-detection of PD is seen in the HSS (Table 5).
 This possibly associates the polarization signature of the source with Comptonized emission.
-We notice $\rm PD<\rm MDP_{99}\%$ in the entire IXPE energy range of $2-8$ keV during SIMS, which possibly resulted from the depolarization of the overall radiation by unpolarized and/or oppositely polarized (as compared to corona) disc emission (Ingram et al., 2024) at lower energies ($2-3$ keV).
-The polarization angle of GX $339-4$ aligns with the direction of the discrete radio jet knot (Mastroserio et al., 2025).
-These findings predict the GX $339-4$ configuration comprising a radially extended weak compact corona coexisting with a strong disc ($F_{\rm ratio}\sim 2.4$, $\Gamma_{\rm th}\sim 2$), as illustrated in Fig.
+We notice $\rm PD<\rm MDP_{99}\%$ in the entire IXPE energy range of 282-8 keV during SIMS, which possibly resulted from the depolarization of the overall radiation by unpolarized and/or oppositely polarized (as compared to corona) disc emission (Ingram et al., 2024) at lower energies (232-3 keV).
+The polarization angle of GX 3394339-4 aligns with the direction of the discrete radio jet knot (Mastroserio et al., 2025).
+These findings predict the GX 3394339-4 configuration comprising a radially extended weak compact corona coexisting with a strong disc ($F_{\rm ratio}\sim 2.4$, $\Gamma_{\rm th}\sim 2$), as illustrated in Fig.
 7c.
@@
 2024).
+Notably, the SIMS of Swift J1727.816131727.8-1613 also corresponds to the configuration shown in Fig.
+7c.
 Furthermore, we conjecture that a similar disc-corona configuration (see Fig.
-7c) is also preferred in the disc-dominated HSS/SIMS of other sources (LMC X33-3- 3, 4U $1957+115$, Cyg X11-1- 1 and Swift J1727.816131727.816131727.8-16131727.8 - 1613) with major contributions from thermal disc and/or various general relativistic effects (Majumder et al., 2024a; Svoboda et al., 2024a; Marra et al., 2024; Steiner et al., 2024; Svoboda et al., 2024b) in the observed PD ($1.4-3\%$).
-It is noteworthy that although the presence of a radial corona is favored for 4U $1630-47$, LMC X33-3- 3, and 4U $1957+115$ in Fig.
+7d) is also preferred in the disc-dominated HSS of other sources (LMC X3-3, 4U $1957+115$, Cyg X1-1 and Swift J1727.816131727.8-1613) with major contributions from thermal disc and/or various general relativistic effects (Majumder et al., 2024a; Svoboda et al., 2024a; Marra et al., 2024; Steiner et al., 2024; Svoboda et al., 2024b) in the observed PD ($1.4-3\%$).
+Among these, quenched radio emission with flux densities of $\sim 2.4-11.7$ mJy is observed in the HSS of Cyg X1-1 (see Table 1 and Fig.
+7d), which is found to be consistent with previous reports of radio quenching in this source (Tigelaar et al., 2004).
+It is noteworthy that although the presence of a radial corona is favored for 4U 1630471630-47, LMC X3-3, and 4U $1957+115$ in Fig.
 7, the coronal geometry remains uncertain due to the lack of resolved radio jet orientation for these sources.
-Moreover, high PD (up to $\sim 21\%$) dominated by reflected emission in Cyg X33-3- 3 appears inconsistent with the disc-corona configurations presented in Fig.
+Moreover, high PD (up to $\sim 21\%$) dominated by reflected emission in Cyg X3-3 appears inconsistent with the disc-corona configurations presented in Fig.
 7.
-It has been suggested that a disc-corona geometry featuring an optically thick funnel in the innermost region of the accretion disc, which obscures most of the primary emission, is favored for Cyg X33-3- 3 Veledina et al.
+It has been suggested that a disc-corona geometry featuring an optically thick funnel in the innermost region of the accretion disc, which obscures most of the primary emission, is favored for Cyg X3-3 Veledina et al.
 (2024a, b).
 Finally, we mention that, although the two-component accretion flow configuration aligns with the disc-corona geometry inferred from polarimetric studies for most of the BH-XRBs, a detailed radiative transfer computation is needed within this framework for the predictions on X-ray polarization.
+Therefore, the schematic in Fig.
+7 illustrates a possible disc-corona-jet geometry in BH-XRBs; however, it does not yet provide a quantitative explanation for the polarimetric results.
 5.4 Frontiers of X-ray Polarimetry: A Promising Era Ahead
-Figure 8: (a) Time binned ($50$ s) light curve of Swift J1727.816131727.816131727.8-16131727.8 - 1613 observed with IXPE in the hard-intermediate state (HIMS).
-(b) Best-fitted power density spectra ($0.1-50$ Hz) obtained from the quasi-simultaneous NICER ($0.5-10$ keV), NuSTAR ($3-20$ keV) and AstroSat ($20-100$ keV) observations.
-(c) Best-fitted energy spectra from the quasi-simultaneous NICER and NuSTAR observations in $0.7-40$ keV energy range.
-(d) IXPE image ($2-8$ keV) of Swift J1727.816131727.816131727.8-16131727.8 - 1613.
-The red circular boundary denotes $60$ arcsec source region considered for polarimetric analysis.
+Figure 8: (a) Time binned ($50$ s) light curve of Swift J1727.816131727.8-1613 observed with IXPE in the hard-intermediate state (HIMS).
+(b) Best-fitted power density spectra (0.1500.1-50 Hz) obtained from the quasi-simultaneous NICER (0.5100.5-10 keV), NuSTAR (3203-20 keV) and AstroSat (2010020-100 keV) observations.
+(c) Best-fitted energy spectra from the quasi-simultaneous NICER and NuSTAR observations in 0.7400.7-40 keV energy range.
+(d) Colour composite X-ray image (NuSTAR: reddish and IXPE: greenish) of Swift J1727.816131727.8-1613 with the overlap in yellow colour.
+The magenta cross marks the source location in X-ray and the cyan circle of radius 60 arcsec around it denotes the region considered for polarimetric analysis.
 (e) Variation of the polarization degree (PD) with energy obtained from the model-independent polarimetric analysis.
 The gray histograms represent the $\rm MDP_{\rm 99}\%$ level.
-See the text for details.
-Indeed, the first X-ray polarization measurements were conducted in the 1970s by OSO-8, revealing significant polarization in Cyg X11-1- 1 with PD $\sim 2.4\%$ at $2.6$ keV and $\sim 5.3\%$ at $5.2$ keV (Long et al., 1980).
-Thereafter, higher polarization was observed in hard X-rays for Cyg X11-1- 1 with INTEGRAL as $\lesssim 20\%$ in the $130-230$ keV band (Jourdain et al., 2012) and $\sim 75\%$ above $400$ keV (Rodriguez et al., 2015).
+Indeed, the first X-ray polarization measurements were conducted in the 1970s by OSO-8, revealing significant polarization in Cyg X1-1 with PD $\sim 2.4\%$ at $2.6$ keV and $\sim 5.3\%$ at $5.2$ keV (Long et al., 1980).
+Thereafter, higher polarization was observed in hard X-rays for Cyg X1-1 with INTEGRAL as $\lesssim 20\%$ in the 130230130-230 keV band (Jourdain et al., 2012) and $\sim 75\%$ above $400$ keV (Rodriguez et al., 2015).
 Recently, Chattopadhyay et al.
-(2024) detected $\sim 23\%$ polarization in the HIMS of Cyg X11-1- 1 using AstroSat/CZTI.
+(2024) detected $\sim 23\%$ polarization in the HIMS of Cyg X1-1 using AstroSat/CZTI.
 Despite these advances, low-energy X-ray polarimetry remained relatively underexplored until the launch of IXPE, which has made substantial progress in this domain during its first three and a half years of operation.
-Nevertheless, several fundamental questions remain unanswered, paving the way for future research.
-At present, Swift J1727.816131727.816131727.8-16131727.8 - 1613 is the only BH-XRB transient observed during its outburst with IXPE exhibiting polarization over four spectral states (HIMS, SIMS, HSS and LHS) along with QPO features.
+Nevertheless, several questions remain unanswered, paving the way for future research.
+At present, Swift J1727.816131727.8-1613 is the only BH-XRB transient observed during its outburst with IXPE exhibiting polarization over four spectral states (HIMS, SIMS, HSS and LHS) along with QPO features.
 To establish robust spectro-polarimetric correlations across different spectral states, coordinated IXPE observations along with simultaneous spectro-temporal coverage from other missions are indispensable.
 Further, since the effects of Comptonization become prominent beyond the coverage of IXPE up to $8$ keV, complementary wide-band X-ray spectro-polarimetry and timing studies are crucial to constrain the coronal geometry.
-Thus, the four windows (see Fig.
-8) of X-ray astronomy (imaging, timing, spectroscopy and polarimetry) with wide energy coverage is crucial to unravel the complete understanding of BH-XRBs (see Fig.
+Thus, the four windows of X-ray astronomy (imaging, timing, spectroscopy and polarimetry) with wide energy coverage is crucial to unravel the complete understanding of BH-XRBs (see Fig.
 8).
 In this context, the recently launched XPoSat [^10] https://www.isro.gov.in/XPoSat.html mission and the upcoming XL-Calibur mission (Abarr et al., 2021) are set to play a crucial role in investigating hard X-ray polarimetric properties of BH-XRBs.
-In particular, POLIX ($8-30$ keV) and XSPECT ($0.8-15$ keV) onboard on XPoSat are suitable for mid-energy polarimetry and simultaneous spectro-temporal studies, respectively.
-The balloon-borne XL-Calibur mission, operating in $15-80$ keV energy range, completed a successful flight [^17] https://sscspace.com/nasa-xl-calibur-balloon-launched/ in July 2024, aiming to measure polarization in Cyg X11-1- 1.
-Additionally, the future soft X-ray spectro-polarimetric mission eXTP [^18] https://heasarc.gsfc.nasa.gov/docs/heasarc/missions/extp.html ($2-10$ keV), expected to launch in $2027$, will offer improved sensitivity for low-energy polarization measurements.
-Needless to mention that these dedicated missions will significantly advance our understanding of the accretion dynamics and coronal geometries in BH-XRBs.
+In particular, POLIX (8308-30 keV) and XSPECT (0.8150.8-15 keV) onboard on XPoSat are suitable for mid-energy polarimetry and simultaneous spectro-temporal studies, respectively.
+The balloon-borne XL-Calibur mission, operating in 158015-80 keV energy range, completed a successful flight [^11] https://sscspace.com/nasa-xl-calibur-balloon-launched/ in July 2024, aiming to measure polarization in Cyg X1-1.
+This eventually led to the recent measurement of a polarization degree of $\sim 5\%$ in LHS of Cyg X1-1 in $\sim 19-64$ keV energy range (Awaki et al., 2025), which is comparable to the low energy (282-8 keV) polarimetric results of Cyg X1-1 (PD $\sim 4.8\%$, LHS) with IXPE (see Fig.
+4 and Table 5).
+The consistency of the measured PDs and PAs across the 2642-64 keV range observed with IXPE and XL-Calibur indicates that the same underlying physical processes likely govern the observed polarization in both energy bands (Awaki et al., 2025).
+In addition, the future soft X-ray spectro-polarimetric mission eXTP [^12] https://heasarc.gsfc.nasa.gov/docs/heasarc/missions/extp.html (2102-10 keV), expected to launch in $2027$, will offer improved sensitivity for low-energy polarization measurements.
+Indeed, these dedicated missions will significantly advance our understanding of the accretion dynamics and coronal geometries in BH-XRBs.
 However, simultaneous observations combining wide-band timing and spectro-polarimetric measurement are still lacking, which emphasizes the need for future missions with broader energy ranges and enhanced capabilities.
@@
 The key findings from our study along with their implications are summarized below.
-  • The combined spectro-temporal results in the wide-band energy range ($0.5-100$ keV) reveal the presence of distinct canonical spectral states of BH-XRBs, which are closely connected to the temporal characteristics and emission mechanism of the sources.
+  • The combined spectro-temporal results in the wide-band energy range (0.51000.5-100 keV) reveal the presence of distinct canonical spectral states of BH-XRBs, which are closely connected to the temporal characteristics and emission mechanism of the sources.
 The timing and spectral features result in degeneracy among the different viable disc-corona geometries of the sources.
@@
   • Timing and spectro-polarimetric results, combined with known radio-jet angles, suggest that a two-component disc-corona model may be plausible.
-In this framework, a radially extended strong corona is likely to be present at the inner part of a truncated accretion disc during harder states, while a comparatively weaker corona persists in softer states for Swift J1727.8--1613, Cyg X--1, GX 339--4, and IGR J17091--3624.
-  • For 4U 1630--47, LMC X--3, and 4U 1957$+$115, a thermally dominated accretion disc with a weak corona seems to be preferred.
-However, the coronal geometry remains unclear due to the minimal contribution of Comptonized emission in the IXPE band and the lack of complementary radio observations, though this analysis does support the presence of disc winds in 4U 1630--47.
-We also note that the disc-corona geometry outlined in this work bears limitations to explain the observed features of Cyg X--3, for which an alternative configuration involving an optically thick inner funnel has been proposed (Veledina et al., 2024a, b).
+In this framework, a radially extended strong corona is likely to be present at the inner part of a truncated accretion disc during harder states, while a comparatively weaker corona persists in softer states for Swift J1727.8-1613, Cyg X-1, GX 339-4, and IGR J17091-3624.
+  • For 4U 1630-47, LMC X-3, and 4U 1957++115, a thermally dominated accretion disc with a weak corona seems to be preferred.
+However, the coronal geometry remains unclear due to the minimal contribution of Comptonized emission in the IXPE band and the lack of complementary radio observations, though this analysis does support the presence of disc winds in 4U 1630-47.
+We also note that the disc-corona geometry outlined in this work bears limitations to explain the observed features of Cyg X-3, for which an alternative configuration involving an optically thick inner funnel has been proposed (Veledina et al., 2024a, b).
 7 Acknowledgments
+Authors thank the anonymous reviewer for constructive comments and useful suggestions that helped to improve the quality of the manuscript.
 SM and SD thank the Department of Physics, IIT Guwahati, India for providing the facilities to complete this work.
@@
 Start (ks) (cts/s) Flux (mCrab) Flux (mCrab) (mJy) State
-($2-20$ keV) ($15-50$ keV)
-Cyg X11-1- 1 E1 IXPE 01002901 2022-05-15 59714.64 242 $9.82\pm 0.21$ $401\pm 11$ $928\pm 34$- LHS
-NICER 5100320101 2022-05-15 59714.26 8.2 $2938\pm 25$- -- -- --
-NuSTAR 30702017002 2022-05-18 59717.57 16 $324\pm 15$- -- -- --
-AstroSat A11_080T01_9000005146 2022-05-15 59714.03 $47.3$ $1505\pm 37$- -- -- --
+(2202-20 keV) (155015-50 keV)
+Cyg X1-1 E1 IXPE 01002901 2022-05-15 59714.64 242 $9.82\pm 0.21$ $401\pm 11$ $928\pm 34$ - LHS
+NICER 5100320101 2022-05-15 59714.26 8.2 $2938\pm 25$ - - - -
+NuSTAR 30702017002 2022-05-18 59717.57 16 $324\pm 15$ - - - -
+AstroSat A11_080T01_9000005146 2022-05-15 59714.03 $47.3$ $1505\pm 37$ - - - -
 E2 IXPE 01250101 2022-06-18 59748.86 86 $10.32\pm 0.12$ $420\pm 14$ $885\pm 33$ - LHS
-NICER 5100320108 2022-06-20 59750.60 3 $3001\pm 25$- -- -- --
-NuSTAR 90802013002 2022-06-20 59750.50 13 $338\pm 14$- -- -- --
+NICER 5100320108 2022-06-20 59750.60 3 $3001\pm 25$ - - - -
+NuSTAR 90802013002 2022-06-20 59750.50 13 $338\pm 14$ - - - -
 E3 IXPE 02008201 2023-05-02 60066.96 21 $26.93\pm 0.18$ $888\pm 12$ $546\pm 21$ $9.1\pm 0.3^{\boxdot}$ HSS
@@
 E5 IXPE 02008401 2023-05-24 60088.82 25 $44.26\pm 0.30$ $809\pm 20$ $569\pm 35$ $2.4\pm 0.2^{\boxdot}$ HSS
-AstroSat T05_105T01_9000005662 2023-05-24 60088.79 60.5 $1354\pm 36$- -- -- --
+AstroSat T05_105T01_9000005662 2023-05-24 60088.79 60.5 $1354\pm 36$ - - - -
 E6 IXPE 02008501 2023-06-13 60108.96 29 $39.33\pm 0.28$ $765\pm 27$ $593\pm 45$ $8.9\pm 0.2^{\boxdot}$ HSS
-NuSTAR 80902318004 2023-06-14 60109.02 9.8 $694\pm 22$- -- -- HSS
+NuSTAR 80902318004 2023-06-14 60109.02 9.8 $694\pm 22$ - - - HSS
 E7 IXPE 02008601 2023-06-20 60115.04 34 $48.09\pm 0.24$ $803\pm 24$ $859\pm 46$ $9.4\pm 0.4^{\boxdot}$ HSS
-NICER 6643010104 2023-06-20 60115.97 7.6 $23227\pm 201$- -- -- --
-NuSTAR 80902318006 2023-06-20 60115.82 10.2 $878\pm 26$- -- -- --
+NICER 6643010104 2023-06-20 60115.97 7.6 $23227\pm 201$ - - - -
+NuSTAR 80902318006 2023-06-20 60115.82 10.2 $878\pm 26$ - - - -
 E8 IXPE 03002201 2024-04-12 60412.02 55.8 $7.28\pm 0.12$ $333\pm 9$ $736\pm 31$ - LHS
-NICER 7100320104 2024-04-11 60411.10 0.7 $1508\pm 18$- -- -- --
+NICER 7100320104 2024-04-11 60411.10 0.7 $1508\pm 18$ - - - -
 E9 IXPE 03003101 2024-05-06 60436.37 53.9 $7.14\pm 0.11$ $283\pm 8$ $919\pm 42$ - LHS
@@
 E11 IXPE 03010101 2024-06-14 60475.65 55.8 $5.72\pm 0.09$ $293\pm 11$ $725\pm 27$ - LHS
-NICER 7706010104 2024-06-14 60475.99 6.1 $1698\pm 54$- -- -- --
+NICER 7706010104 2024-06-14 60475.99 6.1 $1698\pm 54$ - - - -
 E12 IXPE 03002599 2024-10-10 60593.22 110 $22.1\pm 2.2$ $600\pm 13$ $795\pm 27$ - HSS
-4U 16304747-47- 47 E1 IXPE 01250401 2022-08-23 59814.95 459 $4.56\pm 0.08$ $247\pm 8$ $10\pm 2$- HSS
-NICER 5501010102 2022-08-23 59814.01 2 $479\pm 10$- -- -- --
-NuSTAR 80802313002 2022-08-25 59816.19 17 $348\pm 14$- -- -- --
+4U 163047-47 E1 IXPE 01250401 2022-08-23 59814.95 459 $4.56\pm 0.08$ $247\pm 8$ $10\pm 2$ - HSS
+NICER 5501010102 2022-08-23 59814.01 2 $479\pm 10$ - - - -
+NuSTAR 80802313002 2022-08-25 59816.19 17 $348\pm 14$ - - - -
 E2 IXPE 02250601 2023-03-10 60013.78 138 $10.16\pm 0.12$ $550\pm 16$ $246\pm 20$ - SPL
-NICER 6557010101 2023-03-10 60013.76 5 $1009\pm 15$- -- -- --
-NuSTAR 80801327002 2023-03-09 60012.36 12 $884\pm 29$- -- -- --
-AstroSat A12_056T02_9000005538 2023-03-10 60013.08 29.4 $1938\pm 43$- -- -- --
-Cyg X33-3- 3 E1 IXPE 02001899 2022-10-14 59866.04 538 $1.33\pm 0.04$ $106\pm 20$ $129\pm 26$ $106\pm 24^{*}$ LHS
-NICER 5142010105 2022-10-14 59866.28 5 $44.79\pm 3.12$- -- -- --
-NuSTAR 90802323002 2022-10-13 59865.62 18 $197\pm 1$- -- -- --
+NICER 6557010101 2023-03-10 60013.76 5 $1009\pm 15$ - - - -
+NuSTAR 80801327002 2023-03-09 60012.36 12 $884\pm 29$ - - - -
+AstroSat A12_056T02_9000005538 2023-03-10 60013.08 29.4 $1938\pm 43$ - - - -
+Cyg X3-3 E1 IXPE 02001899 2022-10-14 59866.04 538 $1.33\pm 0.04$ $106\pm 20$ $129\pm 26$ $106\pm 24^{*}$ LHS
+NICER 5142010105 2022-10-14 59866.28 5 $44.79\pm 3.12$ - - - -
+NuSTAR 90802323002 2022-10-13 59865.62 18 $197\pm 1$ - - - -
 E2 IXPE 02250301 2022-12-25 59938.41 199 $4.55\pm 0.09$ $295\pm 31$ $144\pm 35$ $107\pm 36^{*}$ HSS
-NuSTAR 90801336002 2022-12-25 59938.33 36 $467\pm 2$- -- -- --
+NuSTAR 90801336002 2022-12-25 59938.33 36 $467\pm 2$ - - - -
 E3 IXPE 02009101 2023-11-17 60265.83 291 $1.43\pm 0.05$ $131\pm 8$ $159\pm 13$ - SIMS
-NICER 6692010101 2023-11-17 60265.77 1.9 $125\pm 7$- -- -- --
+NICER 6692010101 2023-11-17 60265.77 1.9 $125\pm 7$ - - - -
 E4 IXPE 03250301 2024-06-02 60463.77 50 $7.45\pm 0.12$ $411\pm 16$ $14\pm 9$ - HSS
-LMC X11-1- 1 E1 IXPE 02001901 2022-10-19 59871.61 563 $0.96\pm 0.03$ $14\pm 9$ $13\pm 3$- HSS
-NICER 5100070101 2022-10-19 59871.79 2.8 $193\pm 6$- -- -- --
-NuSTAR 90801324002 2022-10-24 59876.16 19 $15\pm 2$- -- -- --
-4U 1957$+115$ E1 IXPE 02006601 2023-05-12 60076.10 572 $1.09\pm 0.04$ $19\pm 4$ $2\pm 1$- HSS
-NICER 6100400101 2023-05-12 60076.55 0.6 $265\pm 8$- -- -- --
-NuSTAR 30902042002 2023-05-15 60079.33 19 $29\pm 4$- -- -- --
-LMC X33-3- 3 E1 IXPE 02006599 2023-07-07 60132.77 562 $0.97\pm 0.03$ $16\pm 4$ $18\pm 3$- HSS
-NICER 6101010117 2023-07-08 60133.42 0.1 $358\pm 23$- -- -- --
-NuSTAR 30902041002 2023-07-09 60134.51 28 $15\pm 3$- -- -- --
+LMC X1-1 E1 IXPE 02001901 2022-10-19 59871.61 563 $0.96\pm 0.03$ $14\pm 9$ $13\pm 3$ - HSS
+NICER 5100070101 2022-10-19 59871.79 2.8 $193\pm 6$ - - - -
+NuSTAR 90801324002 2022-10-24 59876.16 19 $15\pm 2$ - - - -
+4U 1957+115+115 E1 IXPE 02006601 2023-05-12 60076.10 572 $1.09\pm 0.04$ $19\pm 4$ $2\pm 1$ - HSS
+NICER 6100400101 2023-05-12 60076.55 0.6 $265\pm 8$ - - - -
+NuSTAR 30902042002 2023-05-15 60079.33 19 $29\pm 4$ - - - -
+LMC X3-3 E1 IXPE 02006599 2023-07-07 60132.77 562 $0.97\pm 0.03$ $16\pm 4$ $18\pm 3$ - HSS
+NICER 6101010117 2023-07-08 60133.42 0.1 $358\pm 23$ - - - -
+NuSTAR 30902041002 2023-07-09 60134.51 28 $15\pm 3$ - - - -
 E2 IXPE 03004899 2024-10-03 60586.51 385 $1.62\pm 0.04$ $40\pm 5$ $14\pm 9$ - HSS
-NICER 7704010101 2024-10-03 60586.63 2.4 $606\pm 11$- -- -- --
-E3 IXPE 03004901 2024-11-21 60635.94 382 $0.42\pm 0.02$ $8\pm 5$- -- HSS
-NICER 7704010201 2024-11-23 60637.10 1.7 $246\pm 7$- -- -- --
+NICER 7704010101 2024-10-03 60586.63 2.4 $606\pm 11$ - - - -
+E3 IXPE 03004901 2024-11-21 60635.94 382 $0.42\pm 0.02$ $8\pm 5$ - - HSS
+NICER 7704010201 2024-11-23 60637.10 1.7 $246\pm 7$ - - - -
+E4 IXPE 03005001 2025-01-11 60686.18 389 $0.86\pm 0.41$ $11\pm 4$ $9\pm 7$ - HSS
+NuSTAR 31002032006 2025-01-11 60686.68 35 $16\pm 5$ - - -
   • Data are not available in public domain. AMI-LA ($15.5$ GHz, Steiner et al. 2024). RATAN ($4.7$ GHz, Veledina et al. 2024a).
@@
 Start (ks) (cts/s) Flux (mCrab) Flux (mCrab) (mJy) State
-($2-20$ keV) ($15-50$ keV)
-Swift J1727.816131613-1613- 1613 E1 IXPE 02250901 2023-09-07 60194.81 19 $42.21\pm 0.28$ $6053\pm 71$ $4898\pm 173$ $120\pm 12^{**}$ HIMS
-NICER 6203980115 2023-09-08 60195.61 0.2 $64955\pm 274$- -- -- --
-NuSTAR 80902333006 2023-09-07 60194.78 0.7 $6536\pm 8$- -- -- --
-AstroSat T05_145T01_9000005836 2023-09-08 60195.07 $32$ $541\pm 23$- -- -- --
+(2202-20 keV) (155015-50 keV)
+Swift J1727.81613-1613 E1 IXPE 02250901 2023-09-07 60194.81 19 $42.21\pm 0.28$ $6053\pm 71$ $4898\pm 173$ $120\pm 12^{**}$ HIMS
+NICER 6203980115 2023-09-08 60195.61 0.2 $64955\pm 274$ - - - -
+NuSTAR 80902333006 2023-09-07 60194.78 0.7 $6536\pm 8$ - - - -
+AstroSat T05_145T01_9000005836 2023-09-08 60195.07 $32$ $541\pm 23$ - - - -
 E2 IXPE 02251001 2023-09-16 60203.70 37 $41.29\pm 0.25$ $5058\pm 31$ $2472\pm 98$ $144\pm 4.32^{*}$ HIMS
-NICER 6203980119 2023-09-13 60200.18 4.9 $65392\pm 122$- -- -- --
-NuSTAR 80902313002 2023-09-16 60203.81 0.6 $5478\pm 26$- -- -- --
-E3 IXPE 02251101 2023-09-27 60214.92 21 $36.85\pm 0.22$ $3395\pm 49$ $300\pm 10$ $88\pm 2.64^{*}$ HIMS
+NICER 6203980119 2023-09-13 60200.18 4.9 $65392\pm 122$ - - - -
+NuSTAR 80902313002 2023-09-16 60203.81 0.6 $5478\pm 26$ - - - -
+E3 IXPE 02251101 2023-09-27 60214.92 21 $6.85\pm 0.22 33$$395\pm 49$ $300\pm 10$ $88\pm 2.64^{*}$ HIMS
 E4 IXPE 02251201 2023-10-04 60221.54 17 $41.31\pm 0.24$ $4689\pm 46$ $206\pm 38$ $11\pm 0.33^{*}$ SIMS
-NICER 6557020401 2023-10-04 60221.18 6.5 $66263\pm 62$- -- -- --
-NuSTAR 80902313008 2023-10-04 60221.54 0.5 $4111\pm 4$- -- -- --
+NICER 6557020401 2023-10-04 60221.18 6.5 $66263\pm 62$ - - - -
+NuSTAR 80902313008 2023-10-04 60221.54 0.5 $4111\pm 4$ - - - -
 E5 IXPE 02251301 2023-10-10 60227.47 18 $34.62\pm 0.20$ $3446\pm 38$ $206\pm 38$ $35\pm 1.05^{*}$ SIMS
-NICER 6203980136 2023-10-09 60226.02 1.9 $57789\pm 149$- -- -- --
-NuSTAR 80902313016 2023-10-10 60227.82 1.1 $3196\pm 4$- -- -- --
+NICER 6203980136 2023-10-09 60226.02 1.9 $57789\pm 149$ - - - -
+NuSTAR 80902313016 2023-10-10 60227.82 1.1 $3196\pm 4$ - - - -
 E6 IXPE 03005701 2024-02-11 60351.39 67 $8.09\pm 0.13$ $80\pm 7$ $21\pm 12$ - HSS
-NICER 7708010101 2024-02-11 60351.39 3.1 $6415\pm 37$- -- -- --
+NICER 7708010101 2024-02-11 60351.39 3.1 $6415\pm 37$ - - - -
 E7 IXPE 03006001 2024-02-20 60360.07 151 $5.59\pm 0.09$ $67\pm 8$ $26\pm 22$ - HSS
-NICER 7708010106 2024-02-19 60359.07 0.6 $5354\pm 34$- -- -- --
+NICER 7708010106 2024-02-19 60359.07 0.6 $5354\pm 34$ - - - -
 E8 IXPE 03005801 2024-04-03 60403.66 202 $7.87\pm 0.11$ $76\pm 11$ $109\pm 14$ - LHS
-NICER 7708010109 2024-04-03 60403.29 2.5 $675\pm 12$- -- -- --
-GX $339-4$ E1 IXPE 03005101 2024-02-14 60354.93 95 $17.44\pm 0.21$- $79\pm 25$- SIMS
-NICER 7702010112 2024-02-14 60354.98 3.8 $5263\pm 34$- -- -- --
-NuSTAR 91002306002 2024-02-14 60354.70 16 $308\pm 13$- -- -- --
-AstroSat A05_166T01_9000006070 2024-02-14 60354.31 $50.7$ $629\pm 26$- -- -- --
+NICER 7708010109 2024-04-03 60403.29 2.5 $675\pm 12$ - - - -
+GX 3394339-4 E1 IXPE 03005101 2024-02-14 60354.93 95 $17.44\pm 0.21$ - $79\pm 25$ - SIMS
+NICER 7702010112 2024-02-14 60354.98 3.8 $5263\pm 34$ - - - -
+NuSTAR 91002306002 2024-02-14 60354.70 16 $308\pm 13$ - - - -
+AstroSat A05_166T01_9000006070 2024-02-14 60354.31 $50.7$ $629\pm 26$ - - - -
 E2 IXPE 03005301 2024-03-08 60377.68 98 $7.37\pm 0.11$ $103\pm 6$ $12\pm 8$ - HSS
-AstroSat A13_028T01_9000006122 2024-03-10 60379.44 -- -- -- -- -- --
-Swift J151857.0572147151857.0572147151857.0-572147151857.0 - 572147 E1 IXPE 03250201 2024-03-18 60387.15 96 $12.18\pm 0.12$ $422\pm 8$ $152\pm 17$- HSS
-NICER 7204220111 2024-03-18 60387.55 4.2 $1468\pm 18$- -- -- --
-NuSTAR 91001311004 2024-03-18 60387.64 9.2 $444\pm 17$- -- -- --
-IGR J17091362417091362417091-362417091 - 3624 E1 IXPE 04250201 2025-03-07 60741.30 163 $0.27\pm 0.02$ $755\pm 11$ $82\pm 27$- LHS
-NuSTAR 81002342008 2025-03-07 60741.58 21 $27\pm 15$- -- -- --
-MAXI J174429417442941744-2941744 - 294 E1 IXPE 04250301 2025-04-05 60770.09 149 $0.17\pm 0.11$ $45\pm 28$- -- HSS
+AstroSat A13_028T01_9000006122 2024-03-10 60379.44 - - - - - -
+Swift J151857.0572147151857.0-572147 E1 IXPE 03250201 2024-03-18 60387.15 96 $12.18\pm 0.12$ $422\pm 8$ $152\pm 17$ - HSS
+NICER 7204220111 2024-03-18 60387.55 4.2 $1468\pm 18$ - - - -
+NuSTAR 91001311004 2024-03-18 60387.64 9.2 $444\pm 17$ - - - -
+IGR J17091362417091-3624 E1 IXPE 04250201 2025-03-07 60741.30 163 $0.27\pm 0.02$ $755\pm 11$ $82\pm 27$ - LHS
+NuSTAR 81002342008 2025-03-07 60741.58 21 $27\pm 15$ - - - -
+MAXI J17442941744-294 E1 IXPE 04250301 2025-04-05 60770.09 149 $0.17\pm 0.11$ $45\pm 28$ - - HSS
+GRS $1915+105$ E1 IXPE 04003501 2025-05-10 60805.69 142 - $6\pm 3$ - - -
   • Data are not available in public domain. RATAN ($4.7$ GHz, Veledina et al. 2023; Ingram et al. 2024); ∗∗E-MERLIN ($5.1$ GHz, Williams-Baldwin et al. 2023).
@@
 Source Epoch $n_{\rm H}$ $kT_{\rm in}$ $kT_{\rm e}$ $\Gamma_{\rm th}$ $cov_{\rm frac}$ $\chi^{2}/d.o.f$ $F_{\rm disc}^{\boxtimes}$ $F_{\rm Comp}^{\boxtimes}$ $F_{\rm bol}^{\boxtimes}$ $L_{\rm bol}^{\boxtimes}$ Spectral Energy $rms_{\rm tot}^{\dagger}$
-[…]
-Cyg X11-1- 1 E1 $0.44_{-0.02}^{+0.02}$ $0.41_{-0.03}^{+0.03}$ $30.35_{-2.22}^{+2.87}$ $1.61_{-0.01}^{+0.02}$ $0.77_{-0.02}^{+0.02}$ $1517/1468$ $0.68$ $2.45$ $3.07$ $1.4$ LHS $0.7-60$ $32.5$
-($d=2.2$ kpc) E2 $0.34_{-0.02}^{-0.02}$ $0.38_{-0.01}^{+0.01}$ $22.57_{-1.13}^{+1.50}$ $1.61_{-0.01}^{+0.01}$ $0.73_{-0.03}^{+0.04}$ $1479/1403$ $0.74$ $2.53$ $3.29$ $1.5$ LHS $0.7-60$ $31$
-E6 $0.51^{*}$ $0.52_{-0.02}^{+0.02}$ $10^{*}$ $1.01^{*}$ $0.005_{-0.002}^{+0.002}$ $2122/1913$ $1.75$ $0.22$ $4.43$ $2$ HSS $3-60$-
-E7 $0.52_{-0.01}^{+0.01}$ $0.43_{-0.01}^{+0.02}$ $10^{*}$ $3.21_{-0.14}^{+0.11}$ $0.31_{-0.02}^{+0.03}$ $2185/2167$ $3.78$ $0.33$ $6.18$ $2.8$ HSS $0.7-60$ $9.1$
-E8 $0.43_{-0.03}^{+0.03}$ $0.27_{-0.03}^{+0.03}$ $20^{*}$ $1.64_{-0.02}^{+0.02}$ $0.65_{-0.07}^{+0.06}$ $115/133$ $0.37$ $1.15$ $1.53$ $0.7$ LHS $0.7-10$ $41.5$
-E11 $0.39_{-0.04}^{+0.02}$ $0.32_{-0.04}^{+0.03}$ $20^{*}$ $1.62_{-0.03}^{+0.03}$ $0.61_{-0.04}^{+0.04}$ $152/131$ $0.39$ $1.18$ $1.58$ $0.7$ LHS $0.7-10$ $52.4$
-4U 16304747-47- 47 E1 $6.69_{-0.30}^{+0.24}$ $1.36_{-0.01}^{+0.01}$ $20^{*}$ $3.34_{-0.11}^{+0.12}$ $0.052_{-0.007}^{+0.008}$ $1134/928$ $1.56$ $0.03$ $1.59$ $15.1$ HSS $0.7-40$ $6.7$
-($d=10$ kpc) E2 $4.55_{-0.05}^{+0.07}$ $1.46_{-0.02}^{+0.02}$ $20^{*}$ $2.47_{-0.04}^{+0.05}$ $0.38_{-0.01}^{+0.02}$ $2366/1980$ $2.76$ $0.77$ $3.65$ $34.7$ SPL $0.7-60$ $4.8$
-Cyg X33-3- 3 E1 $4.04_{-0.11}^{+0.26}$ $1.21_{-0.02}^{+0.02}$ $6.14_{-0.34}^{+0.28}$ $1.55_{-0.04}^{+0.03}$ $0.83_{-0.01}^{+0.02}$ $2921/1842$ $0.21$ $0.41$ $0.64$ $5.7$ LHS $3-50$-
-($d=9.7$ kpc) E2 $5.54_{-0.13}^{+0.07}$ $1.01_{-0.02}^{+0.01}$ $7.95_{-0.14}^{+0.22}$ $2.21_{-0.01}^{+0.01}$ $0.63_{-0.01}^{+0.01}$ $2709/1903$ $0.94$ $0.62$ $1.58$ $14.1$ HSS $3-50$-
-LMC X11-1- 1 E1 $0.50_{-0.01}^{+0.01}$ $0.98_{-0.01}^{+0.01}$ $10^{*}$ $2.58_{-0.09}^{+0.10}$ $0.091_{-0.010}^{+0.012}$ $744/613$ $0.065$ $0.004$ $0.074$ $16$ HSS $0.8-30$ $10.1$
+($10^{22}$ ${\rm cm}^{-2}$) (keV) (keV) ($\tiny{10^{-8}}$ erg ${\rm cm}^{-2}$ ${\rm s}^{-1}$) ($\%L_{\rm Edd}$) State range (keV) ($\%$)
+Cyg X1-1 E1 $0.44_{-0.02}^{+0.02}$ $0.41_{-0.03}^{+0.03}$ $30.35_{-2.22}^{+2.87}$ $1.61_{-0.01}^{+0.02}$ $0.77_{-0.02}^{+0.02}$ $1517/1468$ $0.68$ $2.45$ $3.07$ $1.4$ LHS 0.7600.7-60 $32.5$
+($d=2.2$ kpc) E2 $0.34_{-0.02}^{-0.02}$ $0.38_{-0.01}^{+0.01}$ $22.57_{-1.13}^{+1.50}$ $1.61_{-0.01}^{+0.01}$ $0.73_{-0.03}^{+0.04}$ $1479/1403$ $0.74$ $2.53$ $3.29$ $1.5$ LHS 0.7600.7-60 $31$
+E6 $0.51^{*}$ $0.52_{-0.02}^{+0.02}$ $10^{*}$ $1.01^{*}$ $0.005_{-0.002}^{+0.002}$ $2122/1913$ $1.75$ $0.22$ $4.43$ 22 HSS 3603-60 -
+E7 $0.52_{-0.01}^{+0.01}$ $0.43_{-0.01}^{+0.02}$ $10^{*}$ $3.21_{-0.14}^{+0.11}$ $0.31_{-0.02}^{+0.03}$ $2185/2167$ $3.78$ $0.33$ $6.18$ $2.8$ HSS 0.7600.7-60 $9.1$
+E8 $0.43_{-0.03}^{+0.03}$ $7_{-0.03}^{+0.03} 2020^{*}$ $1.64_{-0.02}^{+0.02}$ $0.65_{-0.07}^{+0.06}$ $115/133$ $0.37$ $1.15$ $1.53$ $0.7$ LHS 0.7100.7-10 $41.5$
+E11 $0.39_{-0.04}^{+0.02}$ $0.32_{-0.04}^{+0.03}$ $20^{*}$ $1.62_{-0.03}^{+0.03}$ $0.61_{-0.04}^{+0.04}$ $152/131$ $0.39$ $1.18$ $1.58$ $0.7$ LHS 0.7100.7-10 $52.4$
+4U 163047-47 E1 $6.69_{-0.30}^{+0.24}$ $1.36_{-0.01}^{+0.01}$ $20^{*}$ $3.34_{-0.11}^{+0.12}$ $0.052_{-0.007}^{+0.008}$ $1134/928$ $1.56$ $0.03$ $1.59$ $15.1$ HSS 0.7400.7-40 $6.7$
+($d=10$ kpc) E2 $4.55_{-0.05}^{+0.07}$ $1.46_{-0.02}^{+0.02}$ $20^{*}$ $2.47_{-0.04}^{+0.05}$ $0.38_{-0.01}^{+0.02}$ $2366/1980$ $2.76$ $0.77$ $3.65$ $34.7$ SPL 0.7600.7-60 $4.8$
+Cyg X3-3 E1 $4.04_{-0.11}^{+0.26}$ $1.21_{-0.02}^{+0.02}$ $6.14_{-0.34}^{+0.28}$ $1.55_{-0.04}^{+0.03}$ $0.83_{-0.01}^{+0.02}$ $2921/1842$ $0.21$ $0.41$ $0.64$ $5.7$ LHS 3503-50 -
+($d=9.7$ kpc) E2 $5.54_{-0.13}^{+0.07}$ $1.01_{-0.02}^{+0.01}$ $7.95_{-0.14}^{+0.22}$ $2.21_{-0.01}^{+0.01}$ $0.63_{-0.01}^{+0.01}$ $2709/1903$ $0.94$ $0.62$ $1.58$ $14.1$ HSS 3503-50 -
+LMC X1-1 E1 $0.50_{-0.01}^{+0.01}$ $0.98_{-0.01}^{+0.01}$ $10^{*}$ $2.58_{-0.09}^{+0.10}$ $0.091_{-0.010}^{+0.012}$ $744/613$ $0.065$ $0.004$ $0.074$ $16$ HSS 0.8300.8-30 $10.1$
 ($d=48.1$ kpc)
-4U 1957$+115$ E1 $0.08_{-0.01}^{+0.01}$ $1.40_{-0.01}^{+0.01}$ $10^{*}$ $1.86_{-0.10}^{+0.12}$ $0.027_{-0.004}^{+0.004}$ $728/701$ $0.092$ $0.005$ $0.096$ $0.9$ HSS $0.7-40$ $5$
+4U 1957+115+115 E1 $0.08_{-0.01}^{+0.01}$ $1.40_{-0.01}^{+0.01}$ $10^{*}$ $1.86_{-0.10}^{+0.12}$ $0.027_{-0.004}^{+0.004}$ $728/701$ $0.092$ $0.005$ $0.096$ $0.9$ HSS 0.7400.7-40 55
 ($d=10$ kpc)
-LMC X33-3- 3 E1 $0.04^{*}$ $1.10_{-0.01}^{+0.01}$ $10^{*}$ $1.78_{-0.35}^{+0.45}$ $0.005_{-0.002}^{+0.003}$ $567/543$ $0.061$ $0.001$ $0.062$ $13.4$ HSS $0.7-20$-
-($d=48.1$ kpc) E2 $0.04^{*}$ $1.09_{-0.01}^{+0.01}$ $10^{*}$ $2.51_{-0.07}^{+0.08}$ $0.3^{*}$ $153/128$ $0.11$ $0.03$ $0.14$ $30.8$ HSS $1-10$-
-E3 $0.04^{*}$ $0.6_{-0.03}^{+0.04}$ $1.04_{-0.05}^{+0.08}$ $1.72_{-0.08}^{+0.11}$ $0.4^{*}$ $118/117$ $0.03$ $0.006$ $0.035$ $7.7$ HSS $1-10$-
-Swift J1727.816131613-1613- 1613 E1 $0.23_{-0.01}^{+0.01}$ $0.29_{-0.02}^{+0.01}$ $8.44_{-0.11}^{+0.16}$ $1.98_{-0.02}^{+0.04}$ $0.74_{-0.01}^{+0.01}$ $1181/931$ $10.07$ $15.01$ $25.24$ $17.5$ HIMS $0.7-40$ $12$
-($d=2.7$ kpc) E2 $0.20_{-0.01}^{+0.01}$ $0.43_{-0.02}^{+0.02}$ $9.34_{-0.23}^{+0.22}$ $2.11_{-0.04}^{+0.05}$ $0.82_{-0.02}^{+0.02}$ $912/847$ $11.35$ $13.58$ $25.11$ $17.4$ HIMS $0.7-40$-
-E4 $0.25_{-0.01}^{+0.01}$ $0.85_{-0.02}^{+0.03}$ $10^{*}$ $2.74_{-0.04}^{+0.04}$ $0.38_{-0.03}^{+0.04}$ $1373/1447$ $15.89$ $3.63$ $25.04$ $17.3$ SIMS $0.7-40$ $6.1$
-E5 $0.29_{-0.01}^{+0.02}$ $0.88_{-0.02}^{+0.02}$ $10^{*}$ $2.32_{-0.04}^{+0.06}$ $0.19_{-0.02}^{+0.02}$ $1572/1601$ $10.12$ $2.09$ $16.78$ $11.6$ SIMS $0.7-60$ $6.4$
-E6 $0.17_{-0.01}^{+0.01}$ $0.48_{-0.01}^{+0.01}$ $10^{*}$ $5.14_{-0.39}^{+0.36}$ $0.17_{-0.05}^{+0.03}$ $127/132$ $0.91$ $0.02$ $0.93$ $0.6$ HSS $0.7-10$ $1.4$
-E7 $0.18_{-0.01}^{+0.01}$ $0.45_{-0.01}^{+0.01}$ $10^{*}$ $4.57_{-0.60}^{+0.65}$ $0.13_{-0.05}^{+0.09}$ $102/111$ $0.66$ $0.01$ $0.67$ $0.5$ HSS $0.7-10$ $1.3$
-E8 $0.12^{*}$ $0.24_{-0.01}^{+0.01}$ $20^{*}$ $1.71_{-0.02}^{+0.03}$ $0.35_{-0.01}^{+0.01}$ $126/130$ $0.08$ $0.12$ $0.19$ $0.1$ LHS $0.7-10$ $30.2$
-GX $339-4$ E1 $0.51_{-0.01}^{+0.01}$ $0.60_{-0.02}^{+0.02}$ $0.97_{-0.01}^{+0.02}$ $2.02_{-0.52}^{+0.44}$ $0.34_{-0.05}^{+0.08}$ $1603/1538$ $0.84$ $0.36$ $1.71$ $11.5$ SIMS $0.7-50$ $14.2$
+LMC X3-3 E1 $0.04^{*}$ $1.10_{-0.01}^{+0.01}$ $10^{*}$ $1.78_{-0.35}^{+0.45}$ $0.005_{-0.002}^{+0.003}$ $567/543$ $0.061$ $0.001$ $0.062$ $13.4$ HSS 0.7200.7-20 -
+($d=48.1$ kpc) E2 $0.04^{*}$ $9_{-0.01}^{+0.01} 1010^{*}$ $2.51_{-0.07}^{+0.08}$ $0.3^{*}$ $153/128$ $0.11$ $0.03$ $0.14$ $30.8$ HSS 1101-10 -
+E3 $0.04^{*}$ $0.6_{-0.03}^{+0.04}$ $1.04_{-0.05}^{+0.08}$ $1.72_{-0.08}^{+0.11}$ $0.4^{*}$ $118/117$ $0.03$ $0.006$ $0.035$ $7.7$ HSS 1101-10 -
+E4 $0.04^{*}$ $7_{-0.01}^{+0.01} 1010^{*}$ $2.05_{-0.04}^{+0.05}$ $0.005_{-0.003}^{+0.003}$ $771/706$ $0.06$ $0.006$ $0.07$ $15.4$ HSS 3303-30 -
+Swift J1727.81613-1613 E1 $0.23_{-0.01}^{+0.01}$ $0.29_{-0.02}^{+0.01}$ $8.44_{-0.11}^{+0.16}$ $1.98_{-0.02}^{+0.04}$ $0.74_{-0.01}^{+0.01}$ $1181/931$ $10.07$ $15.01$ $25.24$ $17.5$ HIMS 0.7400.7-40 $12$
+($d=2.7$ kpc) E2 $0.20_{-0.01}^{+0.01}$ $0.43_{-0.02}^{+0.02}$ $9.34_{-0.23}^{+0.22}$ $2.11_{-0.04}^{+0.05}$ $0.82_{-0.02}^{+0.02}$ $912/847$ $11.35$ $13.58$ $25.11$ $17.4$ HIMS 0.7400.7-40 -
+E4 $0.25_{-0.01}^{+0.01}$ $0.85_{-0.02}^{+0.03}$ $10^{*}$ $2.74_{-0.04}^{+0.04}$ $0.38_{-0.03}^{+0.04}$ $1373/1447$ $15.89$ $3.63$ $25.04$ $17.3$ SIMS 0.7400.7-40 $6.1$
+E5 $0.29_{-0.01}^{+0.02}$ $0.88_{-0.02}^{+0.02}$ $10^{*}$ $2.32_{-0.04}^{+0.06}$ $0.19_{-0.02}^{+0.02}$ $1572/1601$ $10.12$ $2.09$ $16.78$ $11.6$ SIMS 0.7600.7-60 $6.4$
+E6 $0.17_{-0.01}^{+0.01}$ $0.48_{-0.01}^{+0.01}$ $10^{*}$ $5.14_{-0.39}^{+0.36}$ $0.17_{-0.05}^{+0.03}$ $127/132$ $0.91$ $0.02$ $0.93$ $0.6$ HSS 0.7100.7-10 $1.4$
+E7 $0.18_{-0.01}^{+0.01}$ $0.45_{-0.01}^{+0.01}$ $10^{*}$ $4.57_{-0.60}^{+0.65}$ $0.13_{-0.05}^{+0.09}$ $102/111$ $0.66$ $0.01$ $0.67$ $0.5$ HSS 0.7100.7-10 $1.3$
+E8 $0.12^{*}$ $4_{-0.01}^{+0.01} 2020^{*}$ $1.71_{-0.02}^{+0.03}$ $0.35_{-0.01}^{+0.01}$ $126/130$ $0.08$ $0.12$ $0.19$ $0.1$ LHS 0.7100.7-10 $30.2$
+GX 3394339-4 E1 $0.51_{-0.01}^{+0.01}$ $0.60_{-0.02}^{+0.02}$ $0.97_{-0.01}^{+0.02}$ $2.02_{-0.52}^{+0.44}$ $0.34_{-0.05}^{+0.08}$ $1603/1538$ $0.84$ $0.36$ $1.71$ $11.5$ SIMS 0.7500.7-50 $14.2$
 ($d=8.4$ kpc)
-Swift J151857.0572147151857.0572147151857.0-572147151857.0 - 572147 E1 $3.83_{-0.01}^{+0.02}$ $0.94_{-0.01}^{+0.01}$ $20^{*}$ $2.50_{-0.01}^{+0.01}$ $0.21_{-0.01}^{+0.01}$ $1572/1367$ $2.11$ $0.34$ $2.71$ $8.7$ HSS $0.7-50$ $12.6$
+Swift J151857.0572147151857.0-572147 E1 $3.83_{-0.01}^{+0.02}$ $0.94_{-0.01}^{+0.01}$ $20^{*}$ $2.50_{-0.01}^{+0.01}$ $0.21_{-0.01}^{+0.01}$ $1572/1367$ $2.11$ $0.34$ $2.71$ $8.7$ HSS 0.7500.7-50 $12.6$
 ($d=5.8$ kpc)
-IGR J17091362417091362417091-362417091 - 3624 E1 $1.1^{*}$ $0.88_{-0.38}^{+0.27}$ $23.54_{-2.91}^{+4.01}$ $1.62_{-0.01}^{+0.02}$ $0.87_{-0.03}^{+0.03}$ $1465/1486$ $0.02$ $0.13$ $0.15$ $1.7$ LHS $3-60$ $24.3$
+IGR J17091362417091-3624 E1 $1.1^{*}$ $0.88_{-0.38}^{+0.27}$ $23.54_{-2.91}^{+4.01}$ $1.62_{-0.01}^{+0.02}$ $0.87_{-0.03}^{+0.03}$ $1465/1486$ $0.02$ $0.13$ $0.15$ $1.7$ LHS 3603-60 $24.3$
 ($d=11$ kpc)
   • $L_{\rm Edd}=1.26\times 10^{39}$ erg $\rm s^{-1}$ for $10M_{\odot}$ BH.
-Calculated in $1-100$ keV energy range.
-Computed in $0.1-50$ Hz using NICER observations in $0.5-10$ keV energy range.
+Calculated in 11001-100 keV energy range.
+Computed in 0.1500.1-50 Hz using NICER observations in 0.5100.5-10 keV energy range.
 Frozen values.
@@
 ($a_{*}$) () ($A_{\odot}$) ($\log$ erg $\rm cm$} ${\rm s}^{-1}$) ($\tiny{10^{-8}}$ erg ${\rm cm}^{-2}$ ${\rm s}^{-1}$) State
-aCyg X11-1- 1 E6 $0.98^{*}$ $27^{*}$ $2.32_{-0.06}^{+0.04}$ $6.35_{-1.46}^{+2.08}$ $3.86_{-0.08}^{+0.09}$ $0.67_{-0.04}^{+0.04}$ $2.33$ HSS
+aCyg X1-1 E6 $0.98^{*}$ $27^{*}$ $2.32_{-0.06}^{+0.04}$ $6.35_{-1.46}^{+2.08}$ $3.86_{-0.08}^{+0.09}$ $0.67_{-0.04}^{+0.04}$ $2.33$ HSS
 E7 $0.98^{*}$ $27^{*}$ $2.04_{-0.01}^{+0.01}$ $4.61_{-0.31}^{+0.36}$ $3.31_{-0.05}^{+0.05}$ $1.11_{-0.03}^{+0.03}$ $2.14$ HSS
-bGX $339-4$ E1 $0.85^{*}$ $50^{*}$ $2.27_{-0.03}^{+0.03}$ $0.5^{\boxtimes}$ $3.92_{-0.05}^{+0.12}$ $3.16_{-0.33}^{+0.35}$ $0.54$ SIMS
-cSwift J1727.816131727.816131727.8-16131727.8 - 1613 E4 $0.98^{*}$ $40^{*}$ $2.19_{-0.02}^{+0.03}$ $5.02^{\boxtimes}$ $4.11_{-0.13}^{+0.12}$ $0.68_{-0.14}^{+0.17}$ $5.44$ SIMS
+bGX 3394339-4 E1 $0.85^{*}$ $50^{*}$ $2.27_{-0.03}^{+0.03}$ $0.5^{\boxtimes}$ $3.92_{-0.05}^{+0.12}$ $3.16_{-0.33}^{+0.35}$ $0.54$ SIMS
+cSwift J1727.816131727.8-1613 E4 $0.98^{*}$ $40^{*}$ $2.19_{-0.02}^{+0.03}$ $5.02^{\boxtimes}$ $4.11_{-0.13}^{+0.12}$ $0.68_{-0.14}^{+0.17}$ $5.44$ SIMS
 E5 $0.98^{*}$ $40^{*}$ $2.48_{-0.03}^{+0.02}$ $5.1^{*}$ $4.29_{-0.21}^{+0.14}$ $0.59_{-0.11}^{+0.05}$ $4.38$ SIMS
@@
 Fixed to the best-fitted values.
-Calculated in $1-100$ keV energy range.
+Calculated in 11001-100 keV energy range.
 $A_{\odot}$ is the solar iron abundance.
@@
 (2024a).
-Table 5: Model-independent polarimetric results obtained using the PCUBE algorithm in the $2-8$ keV energy band from IXPE observations of eleven BH-XRBs.
+Table 5: Model-independent polarimetric results obtained using the PCUBE algorithm in the 282-8 keV energy band from IXPE observations of eleven BH-XRBs.
 PD and PA represent the polarization degree and angle, respectively.
-Q///I and U///I are the normalized Stokes parameters.
+Q//I and U//I are the normalized Stokes parameters.
 MDP denotes the minimum detectable polarization, and SIGNIF is the detection significance in units of $\sigma$.
@@
 See text for details.
-Source Epoch Obs. ID PD PA Q///I U///I $\rm MDP_{99}$ SIGNIF Spectral Remarks PDE
+Source Epoch Obs. ID PD PA Q//I U//I $\rm MDP_{99}$ SIGNIF Spectral Remarks PDE
 (%) () (%) (%) (%) ($\sigma$) State Sig. ($\sigma$)
-Cyg X11-1- 1 E1 01002901 $3.70\pm 0.19$ $-20.75\pm 1.50$ $2.77\pm 0.19$ $-2.45\pm 0.19$ $0.59$ $19$ LHS Detection $2.8$
-E2 01250101 $3.79\pm 0.32$ $-25.54\pm 2.39$ $2.38\pm 0.32$ $-2.95\pm 0.32$ $0.96$ $12$ LHS Detection $<1$
-E3 02008201 $2.50\pm 0.40$ $-19.23\pm 4.63$ $1.95\pm 0.40$ $-1.55\pm 0.40$ $1.22$ $5.7$ HSS Detection $<1$
-E4 02008301 $2.46\pm 0.30$ $-22.89\pm 3.45$ $1.71\pm 0.30$ $-1.76\pm 0.30$ $0.90$ $8$ HSS Detection $3$
-E5 02008401 $2.01\pm 0.29$ $-25.82\pm 4.11$ $1.25\pm 0.29$ $-1.58\pm 0.29$ $0.88$ $6.6$ HSS Detection $>4$
-E6 02008501 $1.44\pm 0.28$ $-25.38\pm 5.66$ $0.91\pm 0.28$ $-1.12\pm 0.28$ $0.86$ $4.5$ HSS Detection $1.2$
-E7 02008601 $2.07\pm 0.23$ $-36.37\pm 3.19$ $0.61\pm 0.23$ $-1.98\pm 0.23$ $0.70$ $9$ HSS Detection $1.1$
-E8 03002201 $3.75\pm 0.47$ $-24.22\pm 3.61$ $2.49\pm 0.47$ $-2.80\pm 0.47$ $1.43$ $7.6$ LHS Detection $1.8$
-E9 03003101 $3.04\pm 0.48$ $-18.35\pm 4.54$ $2.44\pm 0.48$ $-1.82\pm 0.48$ $1.46$ $5.9$ LHS Detection $<1$
-E10 03010001 $4.65\pm 0.46$ $-27.77\pm 2.85$ $2.63\pm 0.46$ $-3.83\pm 0.46$ $1.4$ $10.1$ LHS Detection $<1$
-E11 03010101 $4.76\pm 0.53$ $-32.64\pm 3.18$ $1.99\pm 0.53$ $-4.32\pm 0.53$ $1.60$ $9$ LHS Detection $1.5$
-E12 03002599 $2.79\pm 0.19$ $-21.97\pm 1.97$ $2.01\pm 0.19$ $-1.94\pm 0.19$ $0.58$ $14.5$ HSS Detection $>4$
-4U 16304747-47- 47 E1 01250401 $8.34\pm 0.17$ $17.80\pm 0.60$ $6.79\pm 0.17$ $4.86\pm 0.17$ $0.53$ $48$ HSS Detection $>4$
-E2 02250601 $6.77\pm 0.21$ $21.36\pm 0.90$ $4.97\pm 0.21$ $4.59\pm 0.21$ $0.65$ $31.7$ SPL Detection $>4$
-Cyg X33-3- 3 E1 02001899 $20.60\pm 0.31$ $-89.79\pm 0.43$ $-20.60\pm 0.31$ $-0.15\pm 0.31$ $0.94$ $66.3$ LHS Detection $>4$
-E2 02250301 $10.58\pm 0.28$ $-87.39\pm 0.75$ $-10.54\pm 0.28$ $-0.96\pm 0.28$ $0.85$ $38$ HSS Detection $>4$
-E3 02009101 $21.41\pm 0.41$ $-88.09\pm 0.55$ $-21.36\pm 0.41$ $-1.43\pm 0.41$ $1.24$ $52.7$ SIMS Detection $>4$
-E4 03250301 $11.98\pm 0.43$ $-85.22\pm 1.03$ $-11.82\pm 0.43$ $-1.99\pm 0.43$ $1.31$ $27.8$ HSS Detection $>4$
-LMC X11-1- 1 E1 02001901 $1.04\pm 0.40$ $53.97\pm 11.09$ $-0.32\pm 0.40$ $0.99\pm 0.40$ $1.23$ $1.8$ HSS Null-detection --
-4U 1957$+115$ E1 02006601 $1.85\pm 0.37$ $-42.09\pm 5.75$ $0.19\pm 0.37$ $-1.84\pm 0.37$ $1.13$ $4.5$ HSS Detection $1.7$
-LMC X33-3- 3 E1 02006599 $3.04\pm 0.40$ $-44.24\pm 3.77$ $0.08\pm 0.40$ $-3.04\pm 0.40$ $1.21$ $7.2$ HSS Detection $1.8$
-E2 03004899 $2.39\pm 0.37$ $-38.43\pm 4.45$ $0.54\pm 0.37$ $-2.33\pm 0.37$ $1.13$ $6$ HSS Detection $2.3$
-E3 03004901 $2.19\pm 0.77$ $-39.42\pm 10.01$ $0.42\pm 0.77$ $-2.15\pm 0.77$ $2.32$ $2.4$ HSS Null-detection --
-Swift J1727.816131613-1613- 1613 E1 02250901 $4.72\pm 0.28$ $2.50\pm 1.69$ $4.71\pm 0.28$ $0.41\pm 0.28$ $0.84$ $17$ HIMS Detection $<1$
+Cyg X1-1 E1 01002901 $3.70\pm 0.19$ $20.75\pm 1.50$ $2.77\pm 0.19$ $2.45\pm 0.19$ $0.59$ 1919 LHS Detection $2.8$
+E2 01250101 $3.79\pm 0.32$ $25.54\pm 2.39$ $2.38\pm 0.32$ $2.95\pm 0.32$ $0.96$ $12$ LHS Detection <1<1
+E3 02008201 $2.50\pm 0.40$ $19.23\pm 4.63$ $1.95\pm 0.40$ $1.55\pm 0.40$ $1.22$ $5.7$ HSS Detection <1<1
+E4 02008301 $46\pm 0.30 22.$$2.89\pm 3.45$ $1.71\pm 0.30$ $1.76\pm 0.30$ $0.90$ 88 HSS Detection 33
+E5 02008401 $2.01\pm 0.29$ $25.82\pm 4.11$ $1.25\pm 0.29$ $1.58\pm 0.29$ $0.88$ $6.6$ HSS Detection >4>4
+E6 02008501 $1.44\pm 0.28$ $25.38\pm 5.66$ $0.91\pm 0.28$ $1.12\pm 0.28$ $0.86$ $4.5$ HSS Detection $1.2$
+E7 02008601 $2.07\pm 0.23$ $36.37\pm 3.19$ $0.61\pm 0.23$ $1.98\pm 0.23$ $0.70$ 99 HSS Detection $1.1$
+E8 03002201 $3.75\pm 0.47$ $24.22\pm 3.61$ $2.49\pm 0.47$ $2.80\pm 0.47$ $1.43$ $7.6$ LHS Detection $1.8$
+E9 03003101 $3.04\pm 0.48$ $18.35\pm 4.54$ $2.44\pm 0.48$ $1.82\pm 0.48$ $1.46$ $5.9$ LHS Detection <1<1
+E10 03010001 $4.65\pm 0.46$ $27.77\pm 2.85$ $2.63\pm 0.46$ $3.83\pm 0.46$ $1.4$ $10.1$ LHS Detection <1<1
+E11 03010101 $4.76\pm 0.53$ $32.64\pm 3.18$ $1.99\pm 0.53$ $4.32\pm 0.53$ $1.60$ 99 LHS Detection $1.5$
+E12 03002599 $2.79\pm 0.19$ $21.97\pm 1.97$ $2.01\pm 0.19$ $1.94\pm 0.19$ $0.58$ $14.5$ HSS Detection >4>4
+4U 163047-47 E1 01250401 $8.34\pm 0.17$ $17.80\pm 0.60$ $6.79\pm 0.17$ $4.86\pm 0.17$ $0.53$ $48$ HSS Detection >4>4
+E2 02250601 $6.77\pm 0.21$ $21.36\pm 0.90$ $4.97\pm 0.21$ $4.59\pm 0.21$ $0.65$ $31.7$ SPL Detection >4>4
+Cyg X3-3 E1 02001899 $20.60\pm 0.31$ $89.79\pm 0.43$ $20.60\pm 0.31$ $0.15\pm 0.31$ $0.94$ $66.3$ LHS Detection >4>4
+E2 02250301 $10.58\pm 0.28$ $87.39\pm 0.75$ $10.54\pm 0.28$ $0.96\pm 0.28$ $0.85$ $38$ HSS Detection >4>4
+E3 02009101 $21.41\pm 0.41$ $88.09\pm 0.55$ $21.36\pm 0.41$ $1.43\pm 0.41$ $1.24$ $52.7$ SIMS Detection >4>4
+E4 03250301 $11.98\pm 0.43$ $85.22\pm 1.03$ $11.82\pm 0.43$ $1.99\pm 0.43$ $1.31$ $27.8$ HSS Detection >4>4
+LMC X1-1 E1 02001901 $1.04\pm 0.40$ $53.97\pm 11.09$ $0.32\pm 0.40$ $0.99\pm 0.40$ $1.23$ $1.8$ HSS Null-detection -
+4U 1957+115+115 E1 02006601 $1.85\pm 0.37$ $42.09\pm 5.75$ $0.19\pm 0.37$ $84\pm 0.37 1.1$31.13 $4.5$ HSS Detection $1.7$
+LMC X3-3 E1 02006599 $3.04\pm 0.40$ $44.24\pm 3.77$ $0.08\pm 0.40$ $3.04\pm 0.40$ $1.21$ $7.2$ HSS Detection $1.8$
+E2 03004899 $2.39\pm 0.37$ $38.43\pm 4.45$ $0.54\pm 0.37$ $2.33\pm 0.37$ $1.13$ 66 HSS Detection $2.3$
+E3 03004901 $2.19\pm 0.77$ $39.42\pm 10.01$ $0.42\pm 0.77$ $2.15\pm 0.77$ $2.32$ $2.4$ HSS Null-detection -
+E4 03005001 $2.84\pm 0.52$ $47.32\pm 5.24$ $0.23\pm 0.52$ $2.83\pm 0.52$ $1.57$ 55 HSS Detection <1<1
+Swift J1727.81613-1613 E1 02250901 $4.72\pm 0.28$ $2.50\pm 1.69$ $4.71\pm 0.28$ $0.41\pm 0.28$ $0.84$ $17$ HIMS Detection <1<1
 E2 02251001 $4.48\pm 0.20$ $2.28\pm 1.31$ $4.46\pm 0.20$ $0.36\pm 0.20$ $0.62$ $22$ HIMS Detection $2.8$
 E3 02251101 $4.39\pm 0.29$ $0.95\pm 1.87$ $4.38\pm 0.29$ $0.15\pm 0.29$ $0.87$ $15.3$ HIMS Detection $2.6$
-E4 02251201 $3.81\pm 0.29$ $-0.49\pm 2.22$ $3.81\pm 0.29$ $-0.07\pm 0.29$ $0.89$ $12.8$ SIMS Detection $1.1$
-E5 02251301 $3.35\pm 0.32$ $-2.01\pm 2.75$ $3.34\pm 0.32$ $-0.23\pm 0.32$ $0.98$ $10.4$ SIMS Detection $2.5$
-E6 03005701 $1.18\pm 0.44$ $-2.74\pm 10.55$ $1.18\pm 0.44$ $-0.11\pm 0.44$ $1.32$ $2$ HSS Null-detection --
-E7 03006001 $0.40\pm 0.36$ $6.05\pm 25.83$ $0.39\pm 0.36$ $0.08\pm 0.36$ $1.08$ $<1$ HSS Null-detection --
+E4 02251201 $3.81\pm 0.29$ $0.49\pm 2.22$ $3.81\pm 0.29$ $0.07\pm 0.29$ $0.89$ $12.8$ SIMS Detection $1.1$
+E5 02251301 $3.35\pm 0.32$ $2.01\pm 2.75$ $3.34\pm 0.32$ $0.23\pm 0.32$ $0.98$ $10.4$ SIMS Detection $2.5$
+E6 03005701 $1.18\pm 0.44$ $2.74\pm 10.55$ $1.18\pm 0.44$ $0.11\pm 0.44$ $1.32$ 22 HSS Null-detection -
+E7 03006001 $0.40\pm 0.36$ $6.05\pm 25.83$ $39\pm 0.36 0.0$$08\pm 0.36$ $1.08$ <1<1 HSS Null-detection -
 E8 03005801 $3.22\pm 0.60$ $4.18\pm 5.37$ $3.19\pm 0.60$ $0.47\pm 0.60$ $1.83$ $4.8$ LHS Detection $1.6$
 […]
-E2 03005301 $0.47\pm 0.36$ $-25.15\pm 22.07$ $0.30\pm 0.36$ $-0.36\pm 0.36$ $1.09$ $0.8$ HSS Null-detection --
-Swift J151857.0572147151857.0572147151857.0-572147151857.0 - 572147 E1 03250201 $0.25\pm 0.26$ $-22.35\pm 29.98$ $0.17\pm 0.26$ $-0.17\pm 0.26$ $0.78$ $0.5$ HSS Null-detection --
-IGR J17091362417091362417091-362417091 - 3624 E1111 04250201 $9\pm 1.41$ $83.71\pm 4.48$ $-8.78\pm 1.41$ $1.96\pm 1.41$ $4.27$ $6.1$ LHS Detection $1.2$
-MAXI J174429417442941744-2941744 - 294 E1111 04250301 $0.71\pm 0.42$ $-15.37\pm 17.21$ $-0.60\pm 0.42$ $-0.36\pm 0.42$ $1.27$ $1.2$ HSS Null-Detection --
-  • Measurements in $3-8$ keV.
-Null-detection in $2-8$ keV.
-Computed considering $6$ linear energy bins for all sources except GX $339-4$, 4U 1957$+115$ and IGR J17091362417091362417091-362417091 - 3624, where $4$ linear bins are used.
-Table 6: Results from the spectro-polarimetric modeling of IXPE Stokes spectra across different observation epochs of the sources in $2-8$ keV energy band.
+E2 03005301 $0.47\pm 0.36$ $25.15\pm 22.07$ $0.30\pm 0.36$ $0.36\pm 0.36$ $1.09$ $0.8$ HSS Null-detection -
+Swift J151857.0572147151857.0-572147 E1 03250201 $0.25\pm 0.26$ $22.35\pm 29.98$ $0.17\pm 0.26$ $0.17\pm 0.26$ $0.78$ $0.5$ HSS Null-detection -
+IGR J17091362417091-3624 E11 04250201 $9\pm 1.41$ $83.71\pm 4.48$ $8.78\pm 1.41$ $1.96\pm 1.41$ $4.27$ $6.1$ LHS Detection $1.2$
+MAXI J17442941744-294 E11 04250301 $0.71\pm 0.42$ $15.37\pm 17.21$ $0.60\pm 0.42$ $0.36\pm 0.42$ $1.27$ $1.2$ HSS Null-Detection -
+  • Measurements in 383-8 keV.
+Null-detection in 282-8 keV.
+Computed considering $6$ linear energy bins for all sources except GX 3394339-4, 4U 1957+115+115 and IGR J17091362417091-3624, where $4$ linear bins are used.
+Table 6: Results from the spectro-polarimetric modeling of IXPE Stokes spectra across different observation epochs of the sources in 282-8 keV energy band.
 The parameters, with their standard definitions, are obtained from the constant polarization model (polconst) and the energy-dependent polarization model (polpow).
@@
 Source Epoch $A_{\rm norm}$ $A_{\rm index}$ $\psi_{\rm norm}$ $\chi^{2}_{\rm red}$ $\rm PD_{polpow}$ $\rm PA_{polpow}$ $\rm PD_{polconst}$ $\rm PA_{polconst}$ $\chi^{2}_{\rm red}$
-[…]
-Cyg X11-1- 1a E1 $0.020_{-0.006}^{+0.008}$ $-0.43\pm 0.25$ $-19.83\pm 2.22$ $1.01$ $3.93\pm 0.95$ $-19.83\pm 2.22$ $3.40\pm 0.26$ $-19.73\pm 2.24$ $0.91$
-E2 -- -- -- -- -- -- $3.41\pm 0.44$ $-25.13\pm 3.73$ $1.09$
-E3 -- -- -- -- -- -- $2.24\pm 0.61$ $-11.87\pm 7.95$ $1.15$
-E4 $0.008_{-0.004}^{+0.005}$ $-0.90\pm 0.56$ $-17.74\pm 5.45$ $1.13$ $3.39\pm 0.40$ $-17.74\pm 5.45$ $2.29\pm 0.45$ $-18.33\pm 5.68$ $1.13$
-E5 $0.003_{-0.002}^{+0.003}$ $-1.57\pm 0.57$ $-8.63\pm 6.16$ $1.10$ $3.95\pm 0.31$ $-8.63\pm 6.16$ $1.91\pm 0.47$ $-9.48\pm 7.18$ $1.10$
-E6 $0.003\pm 0.002$ $-1.52\pm 0.68$ $-16.26\pm 5.74$ $1.12$ $3.63\pm 0.47$ $-16.26\pm 5.74$ $2.06\pm 0.45$ $-16.64\pm 6.27$ $1.13$
-E7 $0.009\pm 0.005$ $-0.80\pm 0.47$ $-25.03\pm 4.56$ $1.11$ $3.23\pm 0.99$ $-25.03\pm 4.56$ $2.19\pm 0.37$ $-25.77\pm 4.69$ $1.11$
-E8b -- -- -- -- -- -- $3.04\pm 0.66$ $-27.09\pm 6.24$ $1.07$
-E9b -- -- -- -- -- -- $3.35\pm 0.67$ $-25.73\pm 5.79$ $1.25$
-E10b -- -- -- -- -- -- $5.83\pm 0.64$ $-32.72\pm 3.15$ $0.96$
-E11b -- -- -- -- -- -- $4\pm 0.7$ $-34.98\pm 5.27$ $1.05$
-E12 $0.009_{-0.004}^{+0.005}$ $-0.9_{-0.4}^{+0.4}$ $-24.1\pm 3.26$ $1.03$ $3.38\pm 0.48$ $-24.1\pm 3.26$ $2.52\pm 0.48$ $-23.03\pm 3.34$ $1.05$
-4U 16304747-47- 47 E1c $0.028\pm 0.003$ $-0.75\pm 0.07$ $17.90\pm 0.52$ $1.28$ $9.25\pm 0.02$ $17.90\pm 0.52$ $7.89\pm 0.14$ $17.94\pm 0.53$ $1.36$
-E2 $0.022\pm 0.004$ $-0.74\pm 0.11$ $21.45\pm 0.77$ $1.13$ $8.02\pm 0.02$ $21.45\pm 0.77$ $6.27\pm 0.17$ $21.30\pm 0.78$ $1.17$
-Cyg X33-3- 3a E1c $0.028\pm 0.004$ $-0.73\pm 0.09$ $-88.92\pm 0.93$ $1.46$ $8.89\pm 0.11$ $-88.92\pm 0.93$ $7.09\pm 0.24$ $-89.13\pm 0.95$ $1.79$
-E2c $0.018\pm 0.002$ $-0.58\pm 0.11$ $-88.48\pm 0.97$ $1.31$ $4.43\pm 0.16$ $-88.48\pm 0.97$ $3.58\pm 0.21$ $-88.58\pm 1.01$ $1.36$
-E3c $0.024\pm 0.005$ $-0.87\pm 0.12$ $-86.98\pm 3.45$ $1.32$ $9.75\pm 0.21$ $-86.98\pm 3.45$ $7.33\pm 0.31$ $-87.84\pm 1.19$ $1.63$
-E4c $0.11\pm 0.03$ $-0.17\pm 0.11$ $-86.35\pm 1.34$ $1.31$ $14.33\pm 0.44$ $-86.35\pm 1.34$ $13.06\pm 1.01$ $-86.29\pm 2.19$ $1.31$
-4U 1957$+115$ E1 -- -- -- -- -- -- $1.83\pm 0.33$ $-43.73\pm 5.19$ $1.09$
-LMC X33-3- 3 E1 $0.011_{-0.006}^{+0.007}$ $-0.82\pm 0.5$ $-44.29\pm 3.59$ $1.12$ $4.08\pm 0.26$ $-44.29\pm 3.59$ $2.94\pm 0.37$ $-44.94\pm 3.63$ $1.12$
-E2 $0.003_{-0.001}^{+0.002}$ $-1.67\pm 0.51$ $-40.47\pm 4.02$ $1.07$ $4.71\pm 0.58$ $-40.47\pm 4.02$ $2.24\pm 0.56$ $-38.21\pm 7.17$ $1.08$
-Swift J1727.816131613-1613- 1613a E1 $0.022_{-0.009}^{+0.015}$ $-0.42\pm 0.36$ $2.21\pm 2.74$ $1.08$ $4.26\pm 0.21$ $2.21\pm 2.74$ $3.94\pm 0.38$ $2.02\pm 2.75$ $1.08$
-E2 $0.018_{-0.006}^{+0.008}$ $-0.55\pm 0.26$ $1.53\pm 2.08$ $1.38$ $4.29\pm 0.85$ $1.53\pm 2.08$ $3.85\pm 0.28$ $1.41\pm 2.10$ $1.39$
-E3 $0.021_{-0.010}^{+0.018}$ $-0.37\pm 0.23$ $-3.55\pm 3.24$ $1.04$ $3.75\pm 0.88$ $-3.55\pm 3.24$ $3.53\pm 0.39$ $-3.63\pm 3.25$ $1.04$
-E4 $0.007_{-0.004}^{+0.005}$ $-1.08\pm 0.54$ $-1.21\pm 3.76$ $1.05$ $4.00\pm 0.36$ $-1.21\pm 3.76$ $3.11\pm 0.41$ $1.06$
-E5 $0.004_{-0.002}^{+0.003}$ $-1.49\pm 0.54$ $-7.21\pm 4.11$ $0.97$ $4.59\pm 0.41$ $-7.21\pm 4.11$ $2.94\pm 0.45$ $-8.38\pm 4.44$ $0.98$
-E8 $0.006_{-0.003}^{+0.005}$ $-1.28\pm 0.47$ $0.76\pm 4.21$ $0.97$ $4.81\pm 0.17$ $0.76\pm 4.21$ $3.22\pm 0.51$ $2.68\pm 4.54$ $0.97$
-GX $339-4$d E1 -- -- -- -- -- -- $1.16\pm 0.31$ $-72.52\pm 7.78$ $1.03$
-IGR J17091362417091362417091-362417091 - 3624b E1 $0.025_{-0.013}^{+0.016}$ $-0.89\pm 0.43$ $85.11\pm 3.88$ $1.04$ $10.41\pm 0.51$ $85.11\pm 3.88$ $8.24\pm 1.16$ $85.15\pm 4.03$ $1.05$
-  • aOnly IXPE spectra of DU1 is fitted. Not constrained. bdiskbb is not required. cpowerlaw is not required. dSpectral fitting performed within $3-8$ keV range.
+() (polpow) ($\%$) () ($\%$) () (polconst)
+Cyg X1-1a E1 $0.020_{-0.006}^{+0.008}$ $0.43\pm 0.25$ $19.83\pm 2.22$ $1.01$ $3.93\pm 0.95$ $19.83\pm 2.22$ $3.40\pm 0.26$ $19.73\pm 2.24$ $0.91$
+E2 - - - - - - $3.41\pm 0.44$ $25.13\pm 3.73$ $1.09$
+E3 - - - - - - $2.24\pm 0.61$ $1.87\pm 7.95 1.1$51.15
+E4 $0.008_{-0.004}^{+0.005}$ $0.90\pm 0.56$ $7.74\pm 5.45 1.1$31.13 $3.39\pm 0.40$ $17.74\pm 5.45$ $2.29\pm 0.45$ $8.33\pm 5.68 1.1$31.13
+E5 $0.003_{-0.002}^{+0.003}$ $1.57\pm 0.57$ $8.63\pm 6.16$ $1.10$ $3.95\pm 0.31$ $8.63\pm 6.16$ $1.91\pm 0.47$ $9.48\pm 7.18$ $1.10$
+E6 $0.003\pm 0.002$ $1.52\pm 0.68$ $6.26\pm 5.74 1.1$21.12 $3.63\pm 0.47$ $16.26\pm 5.74$ $2.06\pm 0.45$ $6.64\pm 6.27 1.1$31.13
+E7 $0.009\pm 0.005$ $0.80\pm 0.47$ $25.03\pm 4.56$ $1.11$ $3.23\pm 0.99$ $25.03\pm 4.56$ $2.19\pm 0.37$ $25.77\pm 4.69$ $1.11$
+E8b - - - - - - $3.04\pm 0.66$ $27.09\pm 6.24$ $1.07$
+E9b - - - - - - $3.35\pm 0.67$ $25.73\pm 5.79$ $1.25$
+E10b - - - - - - $5.83\pm 0.64$ $32.72\pm 3.15$ $0.96$
+E11b - - - - - - $4\pm 0.7$ $34.98\pm 5.27$ $1.05$
+E12 $0.009_{-0.004}^{+0.005}$ $0.9_{-0.4}^{+0.4}$ $24.1\pm 3.26$ $1.03$ $3.38\pm 0.48$ $24.1\pm 3.26$ $2.52\pm 0.48$ $23.03\pm 3.34$ $1.05$
+4U 163047-47 E1c $0.028\pm 0.003$ $0.75\pm 0.07$ $17.90\pm 0.52$ $1.28$ $9.25\pm 0.02$ $17.90\pm 0.52$ $7.89\pm 0.14$ $17.94\pm 0.53$ $1.36$
+E2 $0.022\pm 0.004$ $0.74\pm 0.11$ $21.45\pm 0.77$ $1.13$ $8.02\pm 0.02$ $21.45\pm 0.77$ $6.27\pm 0.17$ $21.30\pm 0.78$ $1.17$
+Cyg X3-3a E1c $0.028\pm 0.004$ $0.73\pm 0.09$ $88.92\pm 0.93$ $1.46$ $89\pm 0.11 88.$$8.92\pm 0.93$ $7.09\pm 0.24$ $89.13\pm 0.95$ $1.79$
+E2c $0.018\pm 0.002$ $0.58\pm 0.11$ $88.48\pm 0.97$ $1.31$ $4.43\pm 0.16$ $88.48\pm 0.97$ $3.58\pm 0.21$ $88.58\pm 1.01$ $1.36$
+E3c $0.024\pm 0.005$ $0.87\pm 0.12$ $86.98\pm 3.45$ $1.32$ $9.75\pm 0.21$ $86.98\pm 3.45$ $7.33\pm 0.31$ $87.84\pm 1.19$ $1.63$
+E4c $0.11\pm 0.03$ $0.17\pm 0.11$ $86.35\pm 1.34$ $1.31$ $14.33\pm 0.44$ $86.35\pm 1.34$ $13.06\pm 1.01$ $86.29\pm 2.19$ $1.31$
+4U 1957+115+115 E1 - - - - - - $1.83\pm 0.33$ $43.73\pm 5.19$ $1.09$
+LMC X3-3 E1 $0.011_{-0.006}^{+0.007}$ $0.82\pm 0.5$ $44.29\pm 3.59$ $1.12$ $08\pm 0.26 44.$$4.29\pm 3.59$ $2.94\pm 0.37$ $44.94\pm 3.63$ $1.12$
+E2 $0.003_{-0.001}^{+0.002}$ $1.67\pm 0.51$ $40.47\pm 4.02$ $1.07$ $4.71\pm 0.58$ $40.47\pm 4.02$ $2.24\pm 0.56$ $38.21\pm 7.17$ $1.08$
+E4 - - - - - - $2.85\pm 0.83$ $44.18\pm 8.08$ $1.03$
+Swift J1727.81613-1613a E1 $0.022_{-0.009}^{+0.015}$ $0.42\pm 0.36$ $2.21\pm 2.74$ $1.08$ $4.26\pm 0.21$ $2.21\pm 2.74$ $3.94\pm 0.38$ $2.02\pm 2.75$ $1.08$
+E2 $0.018_{-0.006}^{+0.008}$ $0.55\pm 0.26$ $1.53\pm 2.08$ $1.38$ $4.29\pm 0.85$ $1.53\pm 2.08$ $3.85\pm 0.28$ $1.41\pm 2.10$ $1.39$
+E3 $0.021_{-0.010}^{+0.018}$ $0.37\pm 0.23$ $3.55\pm 3.24$ $1.04$ $3.75\pm 0.88$ $3.55\pm 3.24$ $3.53\pm 0.39$ $3.63\pm 3.25$ $1.04$
+E4 $0.007_{-0.004}^{+0.005}$ $1.08\pm 0.54$ $1.21\pm 3.76$ $1.05$ $4.00\pm 0.36$ $1.21\pm 3.76$ $3.11\pm 0.41$ * $1.06$
+E5 $0.004_{-0.002}^{+0.003}$ $1.49\pm 0.54$ $7.21\pm 4.11$ $0.97$ $4.59\pm 0.41$ $7.21\pm 4.11$ $2.94\pm 0.45$ $8.38\pm 4.44$ $0.98$
+E8 $0.006_{-0.003}^{+0.005}$ $1.28\pm 0.47$ $0.76\pm 4.21$ $0.97$ $4.81\pm 0.17$ $0.76\pm 4.21$ $3.22\pm 0.51$ $2.68\pm 4.54$ $0.97$
+GX 3394339-4d E1 - - - - - - $1.16\pm 0.31$ $72.52\pm 7.78$ $1.03$
+IGR J17091362417091-3624b E1 $0.025_{-0.013}^{+0.016}$ $0.89\pm 0.43$ $85.11\pm 3.88$ $1.04$ $10.41\pm 0.51$ $85.11\pm 3.88$ $8.24\pm 1.16$ $85.15\pm 4.03$ $1.05$
+  • aOnly IXPE spectra of DU1 is fitted. Not constrained. bdiskbb is not required. cpowerlaw is not required. dSpectral fitting performed within 383-8 keV range.